### v6

Abstract
This paper studies the convergence of a spatial semidiscretization of a three-dimensional stochastic Allen-Cahn equation with multiplicative noise. For non-smooth initial data, the regularity of the mild solution is investigated, and an error estimate is derived within the spatial $L^{2}$-norm setting. In the case of smooth initial data, two error estimates are established within the framework of general spatial $L^{q}$-norms.
Keywords: stochastic Allen-Cahn equation, spatial semidiscretization, discrete stochastic maximal $L^{p}$-regularity
1 Introduction
Let $\mathcal{O}$ be a bounded convex polyhedral domain in $\mathbb{R}^{3}$, and let $\Delta$ be the Laplace operator with homogeneous Dirichlet boundary condition. We consider the following stochastic Allen-Cahn equation:
$\mathrm{d}y(t)=(\Delta y+y-y^{3})(t)\,\mathrm{d}t+F(y(t))\,\mathrm{d}W_{H}(t),\quad 0<t\leqslant T,$ (1.1a)
$y(0)=v,$ (1.1b)
where the initial value $v$ is given and $W_{H}$ is an infinite-dimensional Wiener process.
The stochastic Allen-Cahn equations are widely used to model phase transitions in materials science; see e.g., [3, 9, 13, 23, 26]. They are also an important class of nonlinear stochastic partial differential equations from a mathematical perspective. Recently, researchers in the field of numerical analysis have shown considerable interest in this topic; see [1, 2, 4, 5, 8, 12, 19, 20, 22, 23, 24, 31] and the references therein. We briefly summarize some closely related works of numerical analysis of three-dimensional stochastic Allen-Cahn equations as follows. Feng et al. [8] analyzed finite element approximations of a stochastic Allen-Cahn equation with gradient-type multiplicative noise. Majee and Prohl [22] derived strong convergence rates for a stochastic Allen-Cahn equation driven by a multiplicative finite-dimensional Wiener process. Bréhier et al. [4] analyzed an explicit temporal splitting numerical scheme for a stochastic Allen-Cahn equation driven by an additive infinite-dimensional Wiener process. Jentzen and Pusnik [12] proposed a numerical method for approximating a class of semilinear stochastic evolution equations with non-globally Lipschitz continuous nonlinearities. Qi and Wang [24] derived strong approximation errors for a stochastic Allen-Cahn equation driven by an additive infinite-dimensional Wiener process. Liu and Qiao [20] established a general theory of the numerical analysis for the stochastic Allen-Cahn equations driven by a multiplicative infinite-dimensional Wiener process. Qi et al. [25] proposed an efficient approximation method to the stochastic Allen-Cahn equation with random diffusion coefficient field and multiplicative noise.
The main motivation of this work includes two components:
  • To the best of our knowledge, the existing works on the numerical analysis of the three-dimensional stochastic Allen-Cahn equation have not adequately addressed scenarios involving non-smooth initial data. This scenario is of practical significance, as initial phases may encompass randomly distributed particles. The inclusion of non-smooth initial data introduces heightened complexity to numerical analysis when compared to cases with smooth initial data. Firstly, the theoretical analysis of the three-dimensional stochastic Allen-Cahn equation with multiplicative noise and non-smooth initial data is relatively scarce (see, e.g., [17, 18]). Secondly, the singularity introduced by non-smooth initial data significantly complicates the process of numerical analysis.
  • In the existing literature, the numerical analysis of the three-dimensional stochastic Allen-Cahn equation with multiplicative noise primarily focuses on the spatial $L^{2}$-norm. However, error estimates employing the more general spatial $L^{q}$-norms are relatively scarce.
Accordingly, the main contributions of this work are as follows:
  • For the model problem Eq. 1.1 with non-smooth initial data $v\in L^{\infty}$, we derive some regularity estimates within the framework of the general spatial $L^{q}$-norms, which are pivotal for the numerical analysis. Specifically, we affirm that
    $y\in L^{p}(\Omega;C([0,T];L^{q})),\quad\forall p,q\in(2,\infty).$
    Our analysis uses a suite of primary tools, including: (i) stochastic integration and Itô’s formula in UMD spaces [7]; (ii) the stochastic maximal $L^{p}$-regularity estimates developed in [29, 30]; (iii) the theory of stochastic partial differential equations with Lyapunov conditions from [17, 18]; (iv) the maximal $L^{p}$-regularity theory for deterministic parabolic equations [32].
  • To handle the singularity caused by the non-smooth initial data, by decomposing the error into a singular part and a relatively non-singular part, we derive the error estimate
    $\lVert{(y-y_{h})(t)}\rVert_{L^{p}(\Omega;L^{2})}\leqslant c\big{(}h^{\alpha/p}+h^{2/p}t^{-1/p}\big{)},\quad\forall t\in(0,T],$
    where $p\in(2,\infty)$ and $\alpha\in(1,2)$ can be arbitrarily chosen. This result is a stochastic counterpart of the well-known error estimate in [28, Theoerem 3.2].
  • Given smooth initial data $v\in W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$, we mainly use the standard theory of finite element methods (see, e.g., [6, 28]), the discrete maximal $L^{p}$-regularity [15], and the discrete stochastic maximal $L^{p}$-regularity [16] to derive two rigorous error estimates within the framework of general spatial $L^{q}$-norms. Specifically, for arbitrary parameters $p\in(2,\infty)$ and $q\in[2,\infty)$, and for any chosen $\epsilon>0$, we establish the following error estimates:
    $\lVert{y-y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2},$
    $\lVert{y-y_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2-\epsilon}.$
To achieve a second-order spatial convergence rate, the previous numerical analysis in the literature typically enforce stringent regularity assumptions on the diffusion term. However, in this study, we impose only some moderate regularity conditions on $F$ (see Section 2). Furthermore, this work introduces a novel approach to the numerical analysis of finite element method-based spatial semidiscretizations for a broad class of nonlinear stochastic parabolic equations.
The rest of this paper is organized as follows. Section 2 introduces some conventions and definitions used in this paper, and the spatial semidiscretization analyzed in this research. Section 3 presents the main results of this paper, including the regularity of Eq. 1.1 with non-smooth initial data and the convergence of the spatial semidiscretization under both non-smooth and smooth initial conditions. Section 4 provides a rigorous proof of the aforementioned results. Lastly, Section 5 summarizes the key findings of this work.
The rest of this paper is organized as follows. Section 2 establishes the foundational notation, definitions, and preliminaries that underpin the analysis within this study, as well as detailing the spatial semidiscretization that is the focus of our investigation. Section 3 is dedicated to the principal findings of this work, which encompass the treatment of the regularity of the solution to (1.1) with nonsmooth initial data, along with the convergence analysis of the spatial semidiscretization scheme under both nonsmooth and smooth initial conditions. Section 4 furnishes a detailed and exacting proof of the results previously outlined. Finally, Section 5 offers a concise synthesis of the pivotal insights and contributions that this research brings to the field.
2 Preliminaries
Notational Conventions. Throughout this paper, we will adhere to the following notational conventions. For Banach spaces $E_{1}$ and $E_{2}$, we denote the space of all bounded linear operators from $E_{1}$ to $E_{2}$ as $\mathcal{L}(E_{1},E_{2})$, and it is abbreviated to $\mathcal{L}(E_{1})$ when $E_{1}=E_{2}$. The identity map is denoted by $I$. For any $\theta\in(0,1)$ and $p\in(1,\infty)$, we use $(\cdot,\cdot)_{\theta,p}$ to represent the real interpolation space constructed by the $K$-method, as described in [21, Chapter 1]. The conjugate exponent of $p$ is denoted by $p^{\prime}$. For $q\in[1,\infty]$, we briefly write $L^{q}(\mathcal{O})$ as $L^{q}$. Let $W_{0}^{1,\infty}(\mathcal{O})$ and $W^{2,\infty}(\mathcal{O})$ denote two standard Sobolev spaces (see, e.g., [14]). Let $(\Omega,\mathcal{F},\mathbb{P})$ represent a complete probability space equipped with a filtration $\mathbb{F}:=(\mathcal{F}_{t})_{t\geqslant 0}$, on which a sequence of independent $\mathbb{F}$-adapted Brownian motions $(\beta_{n})_{n\in\mathbb{N}}$ is defined. The expectation operator associated with this probability space is denoted by $\mathbb{E}$. For $p\in(1,\infty)$ and a Banach space $X$, we use $L^{p}_{\mathbb{F}}(\Omega\times(0,T);X)$ to denote the space of all $\mathbb{F}$-adapted processes in $L^{p}(\Omega\times(0,T);X)$. For any Banach space $X$, the notation $\gamma(H,X)$ denotes the space of all $\gamma$-radonifying operators from $H$ to $X$ (see [11, Chapter 9]). Finally, for each $t\in\mathbb{R}_{+}$, define $W_{H}(t)\in\mathcal{L}(H,L^{2}(\Omega))$ by
$W_{H}(t)h:=\sum_{n\in\mathbb{N}}\beta_{n}(t)(h,h_{n})_{H},\quad\forall h\in H,$
where $H$ is a separable Hilbert space with inner product $(\cdot,\cdot)_{H}$ and an orthonormal basis $(h_{n})_{n\in\mathbb{N}}$.
Definition of Operator $F$. Let $(f_{n})_{n\in\mathbb{N}}$ be a sequence of continuous functions mapping from $\overline{\mathcal{O}}\times\mathbb{R}$ to $\mathbb{R}$, satisfying the following conditions for every $n\in\mathbb{N}$:
$f_{n}(x,0)=0,$ $\forall x\in\partial\mathcal{O},$ (2.1)
$\lvert{f_{n}(x,y)}\rvert\leqslant C_{F}(1+\lvert{y}\rvert),$ $\forall x\in\mathcal{O},\,\forall y\in\mathbb{R},$ (2.2)
$\lvert{f_{n}(x,y_{1})-f_{n}(x,y_{2})}\rvert\leqslant C_{F}\lvert{y_{1}-y_{2}}\rvert,$ $\forall x\in\mathcal{O},\,\forall y_{1},y_{2}\in\mathbb{R},$ (2.3)
$\lvert{f_{n}(x_{1},y)-f_{n}(x_{2},y)}\rvert\leqslant C_{F}(1+\lvert{y}\rvert)\lvert{x_{1}-x_{2}}\rvert,$ $\forall x_{1},x_{2}\in\mathcal{O},\,\forall y\in\mathbb{R},$ (2.4)
where $C_{F}$ is a positive constant independent of $n$, $x$, $y$, $x_{1}$, $x_{2}$, $y_{1}$, and $y_{2}$. Hereafter, with a slight abuse of notation, we will also employ $f_{n}$ to denote the Nemytskii operator defined by
$f_{n}(v):=f_{n}(\cdot,v(\cdot)),\quad\forall v\in L^{1}.$
Let $\{\lambda_{n}\}_{n\in\mathbb{N}}$ be a sequence of non-negative numbers satisfying
$\sum_{n\in\mathbb{N}}\lambda_{n}<\infty.$ (2.5)
Given $v\in L^{2}$, we define $F(v)\in\gamma(H,L^{2})$ via
$F(v):=\sum_{n=0}^{\infty}\sqrt{\lambda}_{n}h_{n}\otimes f_{n}(v).$ (2.6)
Consequently, the model equation Eq. 1.1 can be explicitly stated as follows:
$\begin{cases}\mathrm{d}y(t)=(\Delta y+y-y^{3})(t)\,\mathrm{d}t+\sum_{n=0}^{\infty}\sqrt{\lambda_{n}}f_{n}(y(t))\,\mathrm{d}\beta_{n}(t),\quad 0<t\leqslant T,\\ y(0)=v.\end{cases}$
Spatial Semidiscretization. Let $\mathcal{K}_{h}$ represent a conforming and quasi-uniform triangulation of the domain $\mathcal{O}$, consisting of three-dimensional simplexes. We define the finite element space $X_{h}$ as
$X_{h}:=\left\{v_{h}\in C(\overline{\mathcal{O}})\,\middle|\,v_{h}|_{K}\in P_{r}(K)\text{ for each }K\in\mathcal{K}_{h}\text{ and }v_{h}\text{ vanishes on }\partial\mathcal{O}\right\},$
where $C(\overline{\mathcal{O}})$ is the space of continuous functions over the closure of $\mathcal{O}$, $r\geqslant 1$ is an integer denoting the polynomial degree, and $P_{r}(K)$ is the space of polynomials defined on simplex $K$ up to degree $r$. Let $h$ denote the maximum diameter of the elements in $\mathcal{K}_{h}$. Let $P_{h}$ be the $L^{2}$-orthogonal projection operator onto $X_{h}$, and we define the discrete Laplace operator $\Delta_{h}:X_{h}\to X_{h}$ by
$\int_{\mathcal{O}}(\Delta_{h}u_{h})\cdot v_{h}\,\mathrm{d}\mu=-\int_{\mathcal{O}}\nabla u_{h}\cdot\nabla v_{h}\,\mathrm{d}\mu,\quad\text{for all }u_{h},v_{h}\in X_{h},$
where $\mu$ represents the three-dimensional Lebesgue measure on $\mathcal{O}$. The spatial semidiscretization considered in this work is given by
$\mathrm{d}y_{h}(t)=\left(\Delta_{h}y_{h}+y_{h}-P_{h}(y_{h}^{3})\right)(t)\mathrm{d}t+P_{h}F(y_{h}(t))\,\mathrm{d}W_{H}(t),\quad t>0,$ (2.7a)
$y_{h}(0)=P_{h}v,$ (2.7b)
where $v$ is the given initial value.
Mild Solutions. Let $p\in(1,\infty)$ and $q\in[2,\infty)$. We define the convolution operator $S_{0}$ acting on any function $g\in L^{p}(0,T;L^{q})$ as follows:
$(S_{0}g)(t):=\int_{0}^{t}e^{(t-s)\Delta}g(s)\,\mathrm{d}s,\quad 0\leqslant t\leqslant T.$ (2.8)
Similarly, for any $g\in L_{\mathbb{F}}^{p}(\Omega;L^{2}(0,T;\gamma(H,L^{q})))$, we introduce the stochastic convolution operator $S_{1}$ defined by
$(S_{1}g)(t):=\int_{0}^{t}e^{(t-s)\Delta}g(s)\,\mathrm{d}W_{H}(s),\quad 0\leqslant t\leqslant T.$ (2.9)
An $\mathbb{F}$-adapted process $y$ is called a mild solution to Eq. 1.1 if $y\in L_{\mathbb{F}}^{6}(\Omega\times(0,T);L^{6})$ and satisfies
$y=G+S_{0}(y-y^{3})+S_{1}F(y),$ (2.10)
where $G$ is defined as
$G(t):=e^{t\Delta}v,\quad t\in[0,T].$ (2.11)
In the context of the spatial semidiscretization, for any $g_{h}\in L^{p}(0,T;X_{h})$, we define
$(S_{0,h}g_{h})(t):=\int_{0}^{t}e^{(t-s)\Delta_{h}}g_{h}(t)\,\mathrm{d}t,\quad t\in[0,T].$
For any $g_{h}\in L_{\mathbb{F}}^{p}(\Omega;L^{2}(0,T;\gamma(H,X_{h})))$, the discrete stochastic convolution operator $S_{1,h}$ is given by
$(S_{1,h}g_{h})(t):=\int_{0}^{t}e^{(t-s)\Delta_{h}}g_{h}(t)\,\mathrm{d}W_{H}(t),\quad t\in[0,T].$
An $\mathbb{F}$-adapted $X_{h}$-valued stochastic process $y_{h}$ is called a mild solution to the spatial semidiscretization Eq. 2.7 if $y_{h}\in L_{\mathbb{F}}^{6}(\Omega\times(0,T);X_{h})$ and it satisfies that
$y_{h}=G_{h}+S_{0,h}(y_{h}-P_{h}y_{h}^{3})+S_{1,h}P_{h}F(y_{h}),$
where $G_{h}$ is given by
$G_{h}(t):=e^{t\Delta_{h}}P_{h}v,\quad t\in[0,T].$ (2.12)
3 Main results
In the remainder of this paper, the symbol $c$ represents a generic positive constant that depends solely on $T$, $\mathcal{O}$, $\sum_{n\in\mathbb{N}}\lambda_{h}$, the regularity parameters of the triangulation $\mathcal{K}_{h}$, the indices of the pertinent Lebesgue, Sobolev, and interpolation spaces, the initial value $v$ in Eqs. 1.1 and 2.7, and the exponents related to the spatial mesh size $h$. The value of $c$ may differ from line to line. In addition, for any $\alpha\in[0,\infty)$ and $q\in(1,\infty)$, let $\dot{H}^{\alpha,q}$ denote the domain of $(-\Delta)^{\alpha/2}$ in $L^{q}$, endowed with the standard norm defined as follows:
$\lVert{v}\rVert_{\dot{H}^{\alpha,q}}:=\lVert{(-\Delta)^{\alpha/2}v}\rVert_{L^{q}},\quad\forall v\in\dot{H}^{\alpha,q}.$
The main results of this paper are summarized in the following three theorems.
Theorem 3.1.
Assume that $v\in L^{\infty}$. Then equation Eq. 1.1 admits a unique mild solution $y$, which satisfies that
$y\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}),\qquad\forall p,q\in(2,\infty).$ (3.1)
Furthermore, for any $p\in(2,\infty)$, $q\in[2,\infty)$, and $\epsilon>0$, the solution $y$ possesses the following regularity properties:
$y\in L^{p}\big{(}\Omega;C([0,T];L^{q})\big{)},$ (3.2)
$y\in L^{p}\big{(}\Omega\times(0,T);\dot{H}^{2/p-\epsilon,q}\big{)},$ (3.3)
$y-G\in L^{p}\big{(}\Omega;C([0,T];\dot{H}^{1-\epsilon,q})\big{)},$ (3.4)
$y-G\in L^{p}\big{(}\Omega\times(0,T);\dot{H}^{1+2/p-\epsilon,q}\big{)},$ (3.5)
where $G$ is defined by Eq. 2.11.
Theorem 3.2.
Assume that $v\in L^{\infty}$. Let $y$ and $y_{h}$ be the mild solutions of Eqs. 1.1 and 2.7, respectively. Then, for any $\alpha\in(1,2)$ and $p\in(2,\infty)$, the following error estimate holds for all $t\in(0,T]$:
$\lVert{(y-y_{h})(t)}\rVert_{L^{p}(\Omega;L^{2})}\leqslant c\left(h^{\alpha/p}+h^{2/p}t^{-1/p}\right).$ (3.6)
Theorem 3.3.
Assume that $v\in W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$. Let $y$ and $y_{h}$ be the mild solutions of Eqs. 1.1 and 2.7, respectively. Then, for any $p\in(2,\infty)$, $q\in[2,\infty)$, and $\epsilon>0$, the following error estimates hold:
$\lVert{y-y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2},$ (3.7)
$\lVert{y-y_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2-\epsilon}.$ (3.8)
Remark 3.1.
For Theorem 3.1, we have the following remarks.
  1. 1.
    Liu [17] proposed a theoretical framework for a broad spectrum of nonlinear stochastic partial differential equations. Applying this framework to our model yields the result that
    $y\in L^{p}(\Omega;C([0,T];\dot{H}^{1,2})),\quad\forall p\in(2,\infty),$
    provided that $v\in\dot{H}^{1,2}$. However, the theory in [17] does not cover the case of non-smooth initial data. In the context of the stochastic Allen-Cahn equation with additive noise, Bréhier et al. [4, Lemma 3.1] provide some a priori estimates with the general spatial $L^{q}$ norms.
  2. 2.
    Assume that the condition Eq. 2.1 is removed. By an approximation argument we conclude that Eq. 3.1 remains valid. Moreover, for any $p\in(2,\infty)$ and $q\in[2,\infty)$, the following assertions hold:
    $y-G\in L^{p}(\Omega\times(0,T);\dot{H}^{1,q}),$
    $y-G\in L^{p}(\Omega;C([0,T];(L^{q},\dot{H}^{2,q})_{1/2-1/p,p})).$
Remark 3.2.
For Theorems 3.2 and 3.3, we have the following remarks.
  1. 1.
    In the numerical analysis of the three-dimensional stochastic Allen-Cahn equation with multiplicative infinite-dimensional Wiener process, the existing literature typically imposes stringent regularity conditions on the diffusion coefficient. Our chosen formulation of the function $F$ adheres to the criteria specified in Assumption 2.2 of Liu and Qiao’s work [20]; however, it fails to satisfy the more stringent requirements imposed by their Assumption 2.3. Examining the numerical approximations of Eq. 1.1 under less restrictive conditions on $F$, particularly the relaxation of condition (2.1), holds significant theoretical and practical implications.
  2. 2.
    In Section 4.3, we demonstrate that the norm of $y_{h}$ in $L^{p}(\Omega\times(0,T);L^{q})$ is uniformly bounded with respect to the spatial mesh size $h$ for all $p,q\in[2,\infty)$, provided that $v$ is an element of $W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$ (see Eq. 4.51). It can be rigorously shown that this result remains valid even if $v$ is an element of $W_{0}^{\epsilon,\infty}(\mathcal{O})$, where $\epsilon>0$ can be any arbitrarily small positive number. This conclusion is crucial for the numerical analysis of the fully discrete approximations of Eq. 1.1. However, the question of whether this conclusion extends to the case where $v$ is an element of $L^{\infty}$ remains unresolved.
4 Proofs
The objective of this section is to affirm the validity of Theorems 3.1, 3.2, and 3.3. The structure of this section is as follows:
  • Subsection 4.1 introduces key lemmas that serve as foundational results.
  • Subsection 4.2 focuses on the proof of Theorem 3.1, which establishes the regularity of the mild solution of Eq. 1.1.
  • Subsection 4.3 is dedicated to the proofs of Theorems 3.2 and 3.3, which concern the error estimates of the spatial semidiscretization.
4.1 Preliminary lemmas
Lemma 4.1.
Let $0\leqslant\beta<\gamma\leqslant 1$ and $q\in(1,\infty)$. Then the following continuous embeddings hold:
  • (i)
    $\dot{H}^{\gamma,q}$ is continuously embedded in $(L^{q},\dot{H}^{1,q})_{\beta,q}$;
  • (ii)
    $(L^{q},\dot{H}^{1,q})_{\gamma,q}$ is continuously embedded in $\dot{H}^{\beta,q}$.
Lemma 4.2.
Let $0\leqslant\alpha\leqslant\beta<\infty$ and $q\in(1,\infty)$. Then for all $t\in(0,T]$ we have
$\lVert{e^{t\Delta}}\rVert_{\mathcal{L}(\dot{H}^{\alpha,q},\dot{H}^{\beta,q})}\leqslant ct^{(\alpha-\beta)/2}.$
Lemma 4.3.
Let $p,q\in(1,\infty)$. Then for any $g\in L^{p}(0,T;L^{q})$, the following maximal $L^{p}$-regularity estimate holds:
$\Big{\lVert}{\frac{\mathrm{d}}{\mathrm{d}t}S_{0}g}\Big{\rVert}_{L^{p}(0,T;L^{q})}+\lVert{S_{0}g}\rVert_{L^{p}(0,T;\dot{H}^{2,q})}+\lVert{S_{0}g}\rVert_{C([0,T];(L^{q},\dot{H}^{2,q})_{1-1/p,p})}$
[…]
Lemma 4.4.
Let $p\in(2,\infty)$ and $q\in[2,\infty)$. Then for any $g\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,L^{q}))$, the following stochastic maximal $L^{p}$-regularity estimate holds:
$\lVert{S_{1}g}\rVert_{L^{p}(\Omega;C([0,T];(L^{q},\dot{H}^{2,q})_{1/2-1/p,p}))}+\lVert{S_{1}g}\rVert_{L^{p}(\Omega\times(0,T);\dot{H}^{1,q})}$
[…]
For any $\alpha\in\mathbb{R}$ and $q\in(1,\infty)$, let $\dot{H}_{h}^{\alpha,q}$ denote the space $X_{h}$ equipped with the norm
$\lVert{v_{h}}\rVert_{\dot{H}_{h}^{\alpha,q}}:=\lVert{(-\Delta_{h})^{\alpha/2}v_{h}}\rVert_{L^{q}},\quad\forall v_{h}\in X_{h}.$
Lemma 4.5.
Let $p,q\in(1,\infty)$. Then for any $g_{h}\in L^{p}(0,T;X_{h})$, the following discrete maximal $L^{p}$-regularity estimate holds:
$\Big{\lVert}{\frac{\mathrm{d}}{\mathrm{d}t}S_{0,h}g_{h}}\Big{\rVert}_{L^{p}(0,T;\dot{H}_{h}^{0,q})}+\lVert{S_{0,h}g_{h}}\rVert_{L^{p}(0,T;\dot{H}_{h}^{2,q})}\\ \leqslant c\lVert{g_{h}}\rVert_{L^{p}(0,T;\dot{H}_{h}^{0,q})}.$
Lemma 4.6.
Let $p\in(2,\infty)$ and $q\in[2,\infty)$. Then for any $g_{h}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,X_{h}))$, the following discrete stochastic maximal $L^{p}$-regularity estimate holds:
$\lVert{S_{1,h}g_{h}}\rVert_{L^{p}(\Omega;C([0,T];(\dot{H}_{h}^{0,q},\dot{H}_{h}^{2,q})_{1/2-1/p,p}))}+\lVert{S_{1,h}g_{h}}\rVert_{L^{p}(\Omega\times(0,T);\dot{H}_{h}^{1,q})}$ (4.1)
[…]
Lemma 4.7.
The following assertions are valid:
  1. (i)
    For all $q\in[1,\infty]$, it holds that
    $\lVert{P_{h}}\rVert_{\mathcal{L}(L^{q})}\text{ is uniformly bounded with respect to $h$}.$ (4.2)
  2. (ii)
    For all $\alpha\in[0,2]$ and $q\in(1,\infty)$, $P_{h}$ has the following approximation property:
    $\lVert{I-P_{h}}\rVert_{\mathcal{L}(\dot{H}^{\alpha,q},L^{q})}\leqslant ch^{\alpha}.$ (4.3)
  3. (iii)
    For all $0\leqslant\alpha<\beta\leqslant 2$ and $q\in(1,\infty)$, the following inequality holds:
    $\lVert{\Delta_{h}P_{h}-P_{h}\Delta}\rVert_{\mathcal{L}(\dot{H}^{\beta,q},\dot{H}_{h}^{\alpha-2,q})}\leqslant ch^{\beta-\alpha}.$ (4.4)
  4. (iv)
    For all $\alpha\in[0,2]$, $q\in(1,\infty)$, and $t>0$, the following inequalities are satisfied:
    $\lVert{e^{t\Delta_{h}}}\rVert_{\mathcal{L}(\dot{H}_{h}^{0,q},\dot{H}_{h}^{\alpha,q})}\leqslant ct^{-\alpha/2},$ (4.5)
    $\lVert{e^{t\Delta_{h}}P_{h}-P_{h}e^{t\Delta}}\rVert_{\mathcal{L}(\dot{H}^{\alpha,q},L^{q})}+\lVert{(I-P_{h})e^{t\Delta}}\rVert_{\mathcal{L}(\dot{H}^{\alpha,q},L^{q})}$
    $\leqslant c\min\Big{\{}h^{\alpha},\frac{h^{2}}{t^{1-\alpha/2}}\Big{\}}.$ (4.6)
Remark 4.1.
Lemma 4.1 is established by Proposition 1.4 from [21] and Theorem C.4.1 in [10]. The inequality asserted in Lemma 4.2 is a direct consequence of Theorem 2.12 and Equation (2.128) as presented in [33]. The proof of Lemma 4.3 is provided in [32] and Corollary 1.14 from [21]. The proof of Lemma 4.4 is detailed in [30]. The proof of Lemma 4.5 is accessible in [15], and the proof of Lemma 4.6 can be found in [16]. The assertions in Lemma 4.7 are widely recognized and can be found in standard references such as [6, 28].
Lemma 4.8.
For any $q\in(1,\infty)$, the following inequalities hold:
$\lVert{F(v)}\rVert_{\gamma(H,L^{q})}\leqslant c(1+\lVert{v}\rVert_{L^{q}}),$ $\forall v\in L^{q};$ (4.7)
$\lVert{F(v)}\rVert_{\gamma(H,\dot{H}^{1,q})}\leqslant c(1+\lVert{v}\rVert_{\dot{H}^{1,q}}),$ $\forall v\in\dot{H}^{1,q};$ (4.8)
$\lVert{F(v)}\rVert_{\gamma(H,\dot{H}^{s_{1},q})}\leqslant c(1+\lVert{v}\rVert_{\dot{H}^{s_{2},q}}),$ $\forall v\in\dot{H}^{s_{2},q},\,\forall 0\leqslant s_{1}<s_{2}<1;$ (4.9)
$\lVert{F(u)-F(v)}\rVert_{\gamma(H,L^{q})}\leqslant c\lVert{u-v}\rVert_{L^{q}},$ $\forall u,v\in L^{q}.$ (4.10)
Proof.
We only prove Eq. 4.9, the proof of Eqs. 4.7, 4.8 and 4.10 being easier. Let $\Omega_{\gamma}$ denote a probability space on which a sequence $(\gamma_{n})_{n\in\mathbb{N}}$ of independent standard Gaussian random variables is defined, and let $\mathbb{E}_{\gamma}$ denote the expectation operator with respect to $\Omega_{\gamma}$. The rest of the proof is divided into two steps.
Step 1. We proceed to demonstrate that for any $\theta\in(0,1)$ and $N\in\mathbb{N}_{>0}$, the following inequality holds:
$\lVert{A_{N}v}\rVert_{L^{2}(\Omega_{\gamma};(L^{q},\dot{H}^{1,q})_{\theta,q})}\leqslant c\big{(}1+\lVert{v}\rVert_{(L^{q},\dot{H}^{1,q})_{\theta,q}}\big{)},\quad\forall v\in(L^{q},\dot{H}^{1,q})_{\theta,q},$ (4.11)
where
$A_{N}v:=\sum_{n=0}^{N}\gamma_{n}\sqrt{\lambda_{n}}f_{n}(v),\quad\forall v\in L^{q}.$
For any $v\in\dot{H}^{1,q}$, we infer that
$\lVert{A_{N}v}\rVert_{L^{q}(\Omega_{\gamma};\dot{H}^{1,q})}$ $\leqslant c\Big{(}\mathbb{E}_{\gamma}\int_{\mathcal{O}}\Big{\lvert}{\sum_{n=0}^{N}\gamma_{n}\sqrt{\lambda_{n}}\nabla f_{n}(v)}\Big{\rvert}^{q}\,\mathrm{d}\mu\Big{)}^{1/q}$
$\stackrel{{\scriptstyle\mathrm{(i)}}}{{\leqslant}}c\Big{(}\int_{\mathcal{O}}\Big{(}\mathbb{E}_{\gamma}\Big{\lvert}{\sum_{n=0}^{N}\gamma_{n}\sqrt{\lambda_{n}}\nabla f_{n}(v)}\Big{\rvert}^{2}\Big{)}^{q/2}\,\mathrm{d}\mu\Big{)}^{1/q}$
$=c\Big{(}\int_{\mathcal{O}}\Big{(}\sum_{n=0}^{N}\lambda_{n}\lvert{\nabla f_{n}(v)}\rvert^{2}\Big{)}^{q/2}\,\mathrm{d}\mu\Big{)}^{1/q}$
$\stackrel{{\scriptstyle\mathrm{(ii)}}}{{\leqslant}}c\Big{(}\int_{\mathcal{O}}\Big{(}\sum_{n=0}^{N}\lambda_{n}(1+\lvert{v}\rvert+\lvert{\nabla v}\rvert)^{2}\Big{)}^{q/2}\,\mathrm{d}\mu\Big{)}^{1/q},$
where step $(\mathrm{i})$ invokes Theorem 6.2.6 from [11], and step $(\mathrm{ii})$ relies on Lipschitz conditions Eqs. 2.3 and 2.4. Using the summability condition Eq. 2.5 and the standard inequality
$\left(\int_{\mathcal{O}}|v|^{q}+|\nabla v|^{q}\,\mathrm{d}\mu\right)^{1/q}\leqslant c\lVert{v}\rVert_{\dot{H}^{1,q}},$
we further deduce that
$\lVert{A_{N}v}\rVert_{L^{q}(\Omega_{\gamma};\dot{H}^{1,q})}\leqslant c(1+\lVert{v}\rVert_{\dot{H}^{1,q}}).$
Analogously, employing Eq. 2.2 yields
$\lVert{A_{N}v}\rVert_{L^{q}(\Omega_{\gamma};L^{q})}\leqslant c(1+\lVert{v}\rVert_{L^{q}}),\quad\forall v\in L^{q},$
and utilizing Eq. 2.3 results in
$\lVert{A_{N}u-A_{N}v}\rVert_{L^{q}(\Omega_{\gamma};L^{q})}\leqslant c\lVert{u-v}\rVert_{L^{q}},\quad\forall u,v\in L^{q}.$
Thus, following the methodology outlined in Lemma 28.1 of [27], one obtains
$\lVert{A_{N}v}\rVert_{(L^{q}(\Omega_{\gamma};L^{q}),L^{q}(\Omega_{\gamma};\dot{H}^{1,q}))_{\theta,q}}\leqslant c(1+\lVert{v}\rVert_{(L^{q},\dot{H}^{1,q})_{\theta,q}}),\quad\forall v\in(L^{q},\dot{H}^{1,q})_{\theta,q}.$
This result, when combined with the continuous embedding property
$\left(L^{q}(\Omega_{\gamma};L^{q}),L^{q}(\Omega_{\gamma};\dot{H}^{1,q})\right)_{\theta,q}\hookrightarrow L^{q}\left(\Omega_{\gamma};(L^{q},\dot{H}^{1,q})_{\theta,q}\right),$
yields
$\lVert{A_{N}v}\rVert_{L^{q}(\Omega_{\gamma};(L^{q},\dot{H}^{1,q})_{\theta,q})}\leqslant c\big{(}1+\lVert{v}\rVert_{(L^{q},\dot{H}^{1,q})_{\theta,q}}\big{)},\quad\forall v\in(L^{q},\dot{H}^{1,q})_{\theta,q}.$
Applying Theorem 6.2.6 from [11] once again, we arrive at the assertion in Eq. 4.11.
Step 2. For any $\theta\in(0,1)$ and $v\in(L^{q},\dot{H}^{1,q})_{\theta,q}$, we consider the equality
$\lVert{F(v)}\rVert_{\gamma(H,(L^{q},\dot{H}^{1,q})_{\theta,q})}=\lVert{\lim_{N\to\infty}A_{N}v}\rVert_{L^{2}(\Omega_{\gamma};(L^{q},\dot{H}^{1,q})_{\theta,q})},$
which stems directly from the definition of $F$ as stated in (2.6) and is justified by Theorem 9.1.19 in [11]. By invoking the inequality (4.11), we apply Fatou’s Lemma to deduce that
$\lVert{F(v)}\rVert_{\gamma(H,(L^{q},\dot{H}^{1,q})_{\theta,q})}\leqslant c(1+\lVert{v}\rVert_{(L^{q},\dot{H}^{1,q})_{\theta,q}}).$
Consequently, the desired inequality (4.9) is established through the application of Lemma 4.1. This concludes the proof. ∎
Remark 4.2.
As established by Theorem 4.36 in [21], the interpolation space $(L^{2},\dot{H}^{1,q})_{s,2}$ coincides with $\dot{H}^{s,2}$, sharing equivalent norms across the interval $s\in(0,1)$. Consequently, we refine the estimate Eq. 4.9 to the following form:
$\|F(v)\|_{\gamma(H,\dot{H}^{s,2})}\leqslant c(1+\|v\|_{\dot{H}^{s,2}}),\quad\forall v\in\dot{H}^{s,2},\,\forall s\in(0,1).$
4.2 Proof of Theorem 3.1
It is a well-established fact that the space $L^{2}$ admits an orthonormal basis $\{\phi_{n}\}_{n\in\mathbb{N}}$, comprised of eigenvectors of the negative Laplace operator $-\Delta$. Furthermore, this basis $\{\phi_{n}\}_{n\in\mathbb{N}}$ forms a complete orthogonal system within the space $\dot{H}^{s,2}$ for all $s\in(0,\infty)$.
Lemma 4.9.
For each $n\in\mathbb{N}$ and $v_{n}\in\mathrm{span}\{\phi_{0},\ldots,\phi_{n}\}$, the stochastic Allen-Cahn equation
$\mathrm{d}y_{n}(t)=\big{(}\Delta y_{n}+y_{n}-y_{n}^{3}\big{)}(t)\,\mathrm{d}t+F(y_{n}(t))\,\mathrm{d}W_{H}(t),\quad 0<t\leqslant T,$ (4.12a)
$y_{n}(0)=v_{n}$ (4.12b)
has a unique mild solution. Moreover, for all $p,q\in(2,\infty)$, the following inequality holds:
$\lVert{y_{n}}\rVert_{L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q})}\leqslant c(1+\lVert{v_{n}}\rVert_{L^{q}}).$ (4.13)
Proof.
We split the proof into the following two steps.
Step 1. Let $G_{n}(t):=e^{t\Delta}v_{n}$, $t\in[0,T]$. Since $v_{n}\in\dot{H}^{2,q}$ for each $q\in(2,\infty)$, by Lemma 4.2 with $\alpha=\beta=2$, we have
$G_{n}\in L^{p}(0,T;\dot{H}^{2,q}),\quad\forall p,q\in(2,\infty).$ (4.14)
In view of Eq. 4.8 and the fact that $v_{n}\in\dot{H}^{1,2}$, we use [17, Theorem 1.1] to deduce that Eq. 4.12 has a unique mild solution $y_{n}$ and
$y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{1,2}),\quad\forall p\in(2,\infty).$ (4.15)
The continuous embedding of $\dot{H}^{1,2}$ in $L^{6}$ then leads to
$y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{6}),\quad\forall p\in(2,\infty),$
which implies $y_{n}-y_{n}^{3}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{2})$ for each $p\in(2,\infty)$. Hence, using Lemma 4.3 gives
$S_{0}(y_{n}-y_{n}^{3})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,2}),\quad\forall p\in(2,\infty).$ (4.16)
Since Eqs. 4.8 and 4.15 imply that $F(y_{n})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,\dot{H}^{1,2}))$ for each $p\in(2,\infty)$, by Lemma 4.4 we obtain
$S_{1}F(y_{n})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,2}),\quad\forall p\in(2,\infty).$ (4.17)
Considering the mild solution representation
$y_{n}=G_{n}+S_{0}(y_{n}-y_{n}^{3})+S_{1}F(y_{n}),$ (4.18)
and combining Eqs. 4.14, 4.16 and 4.17, we conclude that
$y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,2}),\quad\forall p\in(2,\infty).$
From the continuous embeddings of $\dot{H}^{2,2}$ into $\dot{H}^{1,6}$ and $\dot{H}^{2,2}$ into $L^{q}$ for all $q\in(2,\infty)$, it follows that
$y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}),\quad\forall p,q\in(2,\infty),$
$y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{1,6}),\quad\forall p\in(2,\infty).$
Hence, Eqs. 4.16 and 4.17 can be improved to
$S_{0}(y_{n}-y_{n}^{3})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,q}),\quad\forall p,q\in(2,\infty),$ (4.19)
$S_{1}F(y_{n})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,6}),\quad\forall p\in(2,\infty).$ (4.20)
Combining Eqs. 4.14, 4.18, 4.19 and 4.20 yields
$y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,6}),\quad\forall p\in(2,\infty),$
thereby implying, through the continuous embedding of $\dot{H}^{2,6}$ into $\dot{H}^{1,q}$ for all $q\in(2,\infty)$, that
$y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{1,q}),\quad\forall p,q\in(2,\infty).$
This regularity result allows us to enhance Eq. 4.17 to
$S_{1}F(y_{n})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,q}),\quad\forall p,q\in(2,\infty),$
which, combined with Eqs. 4.14, 4.18 and 4.19, implies
$y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,q}),\quad\forall p,q\in(2,\infty).$ (4.21)
Step 2. Let $p,q\in(2,\infty)$ be arbitrary but fixed. Considering the regularity assertion Eq. 4.21 and the fact
$y_{n}(t)=v_{n}+\int_{0}^{t}(\Delta y_{n}+y_{n}-y_{n}^{3})(s)\,\mathrm{d}s+\int_{0}^{t}F(y_{n}(s))\,\mathrm{d}W_{H}(s),\quad t\in[0,T],$
we use Itô’s formula in UMD Banach spaces (see [7, Theorem 2.4]) to conclude that, for any $t\in(0,T]$,
$\mathbb{E}\Big{(}\lVert{y_{n}(t)}\rVert_{L^{q}}^{p}-\lVert{v_{n}}\rVert_{L^{q}}^{p}\Big{)}$
[…]
$\quad{}+\frac{p(q-1)}{2}\,\mathbb{E}\int_{0}^{t}\lVert{y_{n}(s)}\rVert_{L^{q}}^{p-q}\,\int_{\mathcal{O}}\lvert{y_{n}(s)}\rvert^{q-2}\sum_{k=0}^{\infty}\lambda_{k}\lvert{f_{k}(y_{n}(s))}\rvert^{2}\,\mathrm{d}\mu\,\mathrm{d}s$
$\quad{}+\frac{p(p-q)}{2}\,\mathbb{E}\int_{0}^{t}\lVert{y_{n}(s)}\rVert_{L^{q}}^{p-2q}\sum_{k=0}^{\infty}\lambda_{k}\Big{[}\int_{\mathcal{O}}\lvert{y_{n}(s)}\rvert^{q-2}y_{n}(s)f_{k}(y_{n}(s))\,\mathrm{d}\mu\Big{]}^{2}\,\mathrm{d}s$
[…]
Considering the inequality
$\int_{\mathcal{O}}\lvert{y_{n}(s)}\rvert^{q-2}y_{n}(s)(\Delta y_{n}-y_{n}^{3})(s)\,\mathrm{d}\mu$
[…]
we infer the bound for $I_{1}$:
$I_{1}\leqslant p\,\mathbb{E}\int_{0}^{t}\lVert{y_{n}(s)}\rVert_{L^{q}}^{p}\,\mathrm{d}s.$
Regarding $I_{2}+I_{3}$, by using Hölder’s inequality in conjunction with Young’s inequality, and employing the summability condition Eq. 2.5 alongside the growth condition Eq. 2.2, a straightforward calculation leads to the following bound for $I_{2}+I_{3}$:
$I_{2}+I_{3}\leqslant c\,\mathbb{E}\int_{0}^{t}1+\lVert{y_{n}(s)}\rVert_{L^{q}}^{p}\,\mathrm{d}s.$
Combining the above bounds for $I_{1}$ and $I_{2}+I_{3}$, we arrive at
$\mathbb{E}\lVert{y_{n}(t)}\rVert_{L^{q}}^{p}\leqslant\lVert{v_{n}}\rVert_{L^{q}}^{p}+c\,\mathbb{E}\int_{0}^{t}1+\lVert{y_{n}(s)}\rVert_{L^{q}}^{p}\,\mathrm{d}s,\quad\forall t\in[0,T].$
Applying Gronwall’s lemma to the above inequality yields
$\lVert{y_{n}(t)}\rVert_{L^{p}(\Omega;L^{q})}\leqslant c\big{(}1+\lVert{v_{n}}\rVert_{L^{q}}\big{)},\quad\forall t\in[0,T],$
which implies the desired inequality Eq. 4.13. This completes the proof. ∎
Proof of Theorem 3.1. We split the proof into the following two steps.
Step 1. Let us prove that Eq. 1.1 admits a unique mild solution $y$ and Eq. 3.1 holds. Let $p,q\in(2,\infty)$ be arbitrary but fixed. For each $n\in\mathbb{N}$, let $y_{n}$ be the mild solution of Eq. 4.12 with $v_{n}$ defined by
$v_{n}:=\mathop{\mathrm{argmin}}_{w\in\mathrm{span}\{\phi_{0},\ldots,\phi_{n}\}}\lVert{v-w}\rVert_{L^{3q}}.$
For any $m,n\in\mathbb{N}$, put $e_{m,n}:=y_{m}-y_{n}$. By definition, we infer that, for any $t\in[0,T]$,
$\mathrm{d}e_{m,n}(t)=(\Delta e_{m,n}+e_{m,n}+y_{n}^{3}-y_{m}^{3})(t)\,\mathrm{d}t+\big{(}F(y_{m}(t))-F(y_{n}(t))\big{)}\,\mathrm{d}W_{H}(t).$
Observing that the integral over $\mathcal{O}$ of $\lvert{e_{m,n}(t)}\rvert^{q-2}e_{m,n}(t)(y_{n}^{3}-y_{m}^{3})(t)$ is non-positive for all $t\in[0,T]$, and following the proof of Eq. 4.13, we can deduce that
$\lVert{e_{m,n}}\rVert_{L_{\mathbb{F}}^{3p}(\Omega\times(0,T);L^{3q})}\leqslant c\lVert{v_{m}-v_{n}}\rVert_{L^{3q}}.$
On the other hand, from the fact that $\dot{H}^{2,2}$ is dense in $L^{3q}$ and that $\{\phi_{n}\}_{n\in\mathbb{N}}$ is a complete orthogonal system in $\dot{H}^{2,2}$, we conclude that $\{v_{n}\}_{n\in\mathbb{N}}$ is a Cauchy sequence in $L^{3q}$, converging to $v$. Consequently, there exists a unique $y\in L_{\mathbb{F}}^{3p}(\Omega\times(0,T);L^{3q})$ such that
$\lim_{n\to\infty}y_{n}=y\quad\text{ in }L_{\mathbb{F}}^{3p}(\Omega\times(0,T);L^{3q}).$ (4.22)
For any $n\in\mathbb{N}$, application of Hölder’s inequality yields
$\lVert{y_{n}^{3}-y^{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}$
[…]
Therefore, using the convergence established in Eq. 4.22, we deduce that
$\lim_{n\to\infty}y_{n}^{3}=y^{3}\quad\text{ in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}),$
consequently leading to
$\lim_{n\to\infty}y_{n}-y_{n}^{3}=y-y^{3}\quad\text{ in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}).$
Using Lemma 4.3, this directly implies the following convergence for $S_{0}(y_{n}-y_{n}^{3})$:
$\lim_{n\to\infty}S_{0}(y_{n}-y_{n}^{3})=S_{0}(y-y^{3})\quad\text{ in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}).$
Furthermore, invoking Eqs. 4.10 and 4.22 yields
$\lim_{n\to\infty}F(y_{n})=F(y)\quad\text{ in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,L^{q})),$
and subsequently by Lemma 4.4 we infer the following convergence for $S_{1}F(y_{n})$:
$\lim_{n\to\infty}S_{1}F(y_{n})=S_{1}F(y)\quad\text{ in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}).$
Additionally, applying Lemma 4.2 with $\alpha=\beta=0$ and noting the convergence $\lim_{n\in\infty}\lVert{v_{n}-v}\rVert_{L^{q}}=0$, we establish the following convergence for $G_{n}$:
$\lim_{n\to\infty}G_{n}=G\quad\text{ in }L^{p}(0,T;L^{q}),$
where $G$ is defined by Eq. 2.11, and $G_{n}(t):=e^{t\Delta}v_{n}$ for each $t\in[0,T]$. Finally, combining the derived convergences of $y_{n}$, $G_{n}$, $S_{0}(y_{n}-y_{n}^{3})$, and $S_{1}F(y_{n})$, and recalling that $y_{n}=G_{n}+S_{0}(y_{n}-y_{n}^{3})+S_{1}F(y_{n})$, by passing to the limit $n\to\infty$ we obtain
$y=G+S_{0}(y-y^{3})+S_{1}F(y)\quad\text{in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}).$
This confirms that $y$ is a mild solution of Eq. 1.1, and it is evident that Eq. 3.1 holds. The uniqueness of the mild solution can be proved similarly as in [17, Theorem 1.1].
Step 2. Fix any $p\in(2,\infty)$ and $q\in[2,\infty)$. By virtue of the regularity assertion in Eq. 3.1 and the growth estimate Eq. 4.7 for $F$, we deduce that $y-y^{3}$ belongs to $L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q})$ and $F(y)$ is an element of $L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,L^{q}))$. Hence, the application of Lemma 4.3, in conjunction with Lemma 4.1, gives that
$S_{0}(y-y^{3})\in L^{p}\big{(}\Omega;L^{p}(0,T;\dot{H}^{2,q})\cap C([0,T];\dot{H}^{2-2/p-\epsilon,q})\big{)},\quad\forall\epsilon>0,$ (4.23)
while Lemma 4.4 implies $S_{1}F(y)\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{1,q})$. Additionally, Lemma 4.2 ensures that $G\in L^{p}(0,T;\dot{H}^{2/p-\epsilon,q})$ for all $\epsilon>0$. Consequently, by the mild solution representation Eq. 2.10, we arrive at
$y\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2/p-\epsilon,q}),\quad\forall\epsilon>0,$
which confirms the regularity assertion Eq. 3.3. In conjunction with Eq. 4.9, we further infer that
$F(y)\in L_{\mathbb{F}}^{p}\big{(}\Omega\times(0,T);\gamma(H,\dot{H}^{2/p-\epsilon,q})\big{)},\quad\forall\epsilon>0,$
leading to the enhanced conclusion via Lemmas 4.1 and 4.4 that
$S_{1}F(y)\in L^{p}\big{(}\Omega;L^{p}(0,T;\dot{H}^{1+2/p-\epsilon,q})\cap C([0,T];\dot{H}^{1-\epsilon,q})\big{)},\quad\forall\epsilon>0.$ (4.24)
Finally, combining Eqs. 2.10, 4.23 and 4.24 establishes the regularity assertions in Eqs. 3.4 and 3.5. Moreover, the regularity assertion in Eq. 3.2 follows directly from Eq. 3.4 and the observation that $G\in C([0,T];L^{q})$. This completes the proof of Theorem 3.1.
$\blacksquare$
4.3 Proof of Theorems 3.2 and 3.3
We first summarize some standard results as follows.
Lemma 4.10.
Let $p,q\in(1,\infty)$. For any $g_{h}\in L^{p}(0,T;X_{h})$, the following inequality holds for all $\epsilon>0$:
$\max_{t\in[0,T]}\|S_{0,h}g_{h}(t)\|_{\dot{H}_{h}^{2-\frac{2}{p}-\epsilon,q}}\leqslant c\|g_{h}\|_{L^{p}(0,T;L^{q})}.$
Proof.
For any $t\in[0,T]$, recognizing that $S_{0,h}g_{h}(t)=\int_{0}^{t}e^{(t-s)\Delta_{h}}g_{h}(s)\,\mathrm{d}s$, we proceed as follows:
$\|S_{0,h}g_{h}(t)\|_{\dot{H}_{h}^{2-\frac{2}{p}-\epsilon,q}}$ $=\left\|\int_{0}^{t}e^{(t-s)\Delta_{h}}g_{h}(s)\,\mathrm{d}s\right\|_{\dot{H}_{h}^{2-2/p-\epsilon,q}}$
$\leqslant\int_{0}^{t}\left\|e^{(t-s)\Delta_{h}}g_{h}(s)\right\|_{\dot{H}_{h}^{2-2/p-\epsilon,q}}\,\mathrm{d}s$
$\stackrel{{\scriptstyle\text{(i)}}}{{\leqslant}}c\int_{0}^{t}(t-s)^{-1+\frac{1}{p}+\frac{\epsilon}{2}}\|g_{h}(s)\|_{L^{q}}\,\mathrm{d}s$
$\stackrel{{\scriptstyle\text{(ii)}}}{{\leqslant}}c\Big{(}\int_{0}^{t}(t-s)^{(-1+\frac{1}{p}+\frac{\epsilon}{2})\frac{p}{p-1}}\,\mathrm{d}s\Big{)}^{1-\frac{1}{p}}\|g_{h}\|_{L^{p}(0,T;L^{q})},$
where step (i) invokes Eq. 4.5 with $\alpha=2-2/p-\epsilon$, and step (ii) uses Hölder’s inequality. The uniform boundedness of the integral
$\int_{0}^{t}(t-s)^{\left(-1+\frac{1}{p}+\frac{\epsilon}{2}\right)\frac{p}{p-1}}\,\mathrm{d}s$
with respect to $t\in[0,T]$ then establishes the lemma. ∎
Lemma 4.11.
Assume that $v\in L^{\infty}$. Let $y$ and $y_{h}$ be the mild solutions of Eqs. 1.1 and 2.7, respectively, and let $G$ and $G_{h}$ be defined by Eqs. 2.11 and 2.12, respectively. For simplicity, let $z:=y-G$ and $\xi_{h}:=S_{0,h}(\Delta_{h}P_{h}-P_{h}\Delta)z$. Then, for any $p\in(2,\infty)$, $q\in[2,\infty)$, and $\varepsilon>0$, the following inequalities hold:
$\lVert{G-G_{h}}\rVert_{L^{p}(0,T;L^{q})}+\lVert{P_{h}G-G_{h}}\rVert_{L^{p}(0,T;L^{q})}\leqslant ch^{\frac{2}{p}},$ (4.25)
$\lVert{z-P_{h}z}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{1+\frac{2}{p}-\varepsilon},$ (4.26)
$\lVert{\xi_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{1+\frac{2}{p}-\varepsilon},$ (4.27)
$\lVert{\xi_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{1-\varepsilon},$ (4.28)
$\mathbb{E}\int_{0}^{T}\lVert{(y-P_{h}y)(t)}\rVert_{L^{4}}^{p}\lVert{y(t)}\rVert_{L^{8}}^{2p}\,\mathrm{d}t\leqslant ch^{2-\varepsilon}.$ (4.29)
Furthermore, for any $\beta\in\Big{(}1,\frac{3p}{3p-4}\Big{)}$, the following inequalities hold:
$\mathbb{E}\int_{0}^{T}\lVert{(P_{h}G-G_{h})(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\big{(}\lVert{y(t)}\rVert_{L^{2\beta}}^{p}+\lVert{y_{h}(t)}\rVert_{L^{2\beta}}^{p}\big{)}\,\mathrm{d}t\leqslant ch^{2-\frac{3}{2}\big{(}1-\frac{1}{\beta}\big{)}p},$ (4.30)
$\mathbb{E}\int_{0}^{T}\lVert{\xi_{h}(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\big{(}\lVert{y(t)}\rVert_{L^{2\beta}}^{p}+\lVert{y_{h}(t)}\rVert_{L^{2\beta}}^{p}\big{)}\,\mathrm{d}t\leqslant ch^{p+2-\frac{3}{2}\big{(}1-\frac{1}{\beta}\big{)}p-\varepsilon}.$ (4.31)
Proof.
The inequality Eq. 4.25 is evident by Eq. 4.6, and the inequality Eq. 4.26 follows by Eq. 4.3 and the regularity assertion Eq. 3.5. The inequality Eq. 4.27 is derived as follows:
[…]
The inequality Eq. 4.28 is derived in a similarly manner:
$\lVert{\xi_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}$
[…]
The inequality Eq. 4.29 is straightforward to verify by applying Hölder’s inequality, in conjunction with the approximation property of $P_{h}$ from Eq. 4.3 and the regularity assertions in Theorem 3.1.
Now let us prove Eq. 4.30. By invoking Eq. 4.2 and replicating the reasoning behind [18, Equation (2.2)], we establish that
$\mathbb{E}\int_{0}^{T}\lVert{y_{h}(t)}\rVert_{L^{2}}^{\frac{4\beta}{3\beta-3}-2}\lVert{y_{h}(t)}\rVert_{\dot{H}^{1,2}}^{2}\,\mathrm{d}t$
is uniformly bounded with respect to $h$. Using the continuous embedding of $\dot{H}^{1,2}$ into $L^{6}$, this uniform bound extends to
$\mathbb{E}\int_{0}^{T}\lVert{y_{h}(t)}\rVert_{L^{2}}^{\frac{4\beta}{3\beta-3}-2}\lVert{y_{h}(t)}\rVert_{L^{6}}^{2}\,\mathrm{d}t.$
Employing Hölder’s inequality, we deduce for all $t\in[0,T]$ that
$\lVert{y_{h}(t)}\rVert_{L^{2\beta}}^{\frac{4\beta}{3\beta-3}}\leq\lVert{y_{h}(t)}\rVert_{L^{2}}^{\frac{4\beta}{3\beta-3}-2}\lVert{y_{h}(t)}\rVert_{L^{6}}^{2},$
which in turn confirms the uniform bound on
$\lVert{y_{h}}\rVert_{L^{\frac{4\beta}{3\beta-3}}(\Omega\times(0,T);L^{2\beta})}.$
Subsequently, applying Hölder’s inequality once more yields
$\mathbb{E}\int_{0}^{T}\lVert{(P_{h}G-G_{h})(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\lVert{y_{h}(t)}\rVert_{2\beta}^{p}\,\mathrm{d}t$
$\leqslant\lVert{P_{h}G-G_{h}}\rVert_{L^{\frac{4p\beta}{4\beta-(3\beta-3)p}}(\Omega\times(0,T);L^{2\beta^{\prime}})}^{p}\lVert{y_{h}}\rVert_{L^{\frac{4\beta}{3\beta-3}}(\Omega\times(0,T);L^{2\beta})}^{p}$
$\leqslant c\lVert{P_{h}G-G_{h}}\rVert_{L^{\frac{4p\beta}{4\beta-(3\beta-3)p}}(\Omega\times(0,T);L^{2\beta^{\prime}})}^{p}.$
Invoking the inequality Eq. 4.25 with $p$ replaced by $4p\beta(4\beta-(3\beta-3)p)$, we find
$\mathbb{E}\int_{0}^{T}\lVert{(P_{h}G-G_{h})(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\lVert{y_{h}(t)}\rVert_{2\beta}^{p}\,\mathrm{d}t\leqslant ch^{2-\frac{3}{2}(1-\frac{1}{\beta})p}.$
Given that Theorem 3.1 ensures $y$ resides within $L^{\frac{4\beta}{3\beta-3}}(\Omega\times(0,T);L^{2\beta})$, we can mirror the preceding argument to deduce that
$\mathbb{E}\int_{0}^{T}\lVert{(P_{h}G-G_{h})(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\lVert{y(t)}\rVert_{2\beta}^{p}\,\mathrm{d}t\leqslant ch^{2-\frac{3}{2}(1-1/\beta)p}.$
The inequality Eq. 4.30 is thus substantiated by combining the above derived inequalities.
Finally, by invoking Eq. 4.27 and adhering to the same logical progression as demonstrated in the proof of Eq. 4.30, we deduce Eq. 4.31. This concludes the proof. ∎
Proof of Theorem 3.2. Fix any $\alpha\in(1,2)$ and $p\in(2,\infty)$. Let
$z(t)$ $:=y(t)-G(t),\quad 0\leqslant t\leqslant T,$ (4.32)
$z_{h}(t)$ $:=y_{h}(t)-G_{h}(t),\quad 0\leqslant t\leqslant T,$ (4.33)
where $G$ and $G_{h}$ are defined by Eqs. 2.11 and 2.12, respectively. From Eq. 1.1 and the fact $G^{\prime}(t)=\Delta G(t)$ for $t\in[0,T]$, we deduce that
$\mathrm{d}z(t)=\big{(}\Delta z+z+G-y^{3}\big{)}(t)\,\mathrm{d}t+F(y(t))\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T.$
Consequently,
$\mathrm{d}P_{h}z(t)=P_{h}(\Delta z+z+G-y^{3})(t)\,\mathrm{d}t+P_{h}F(y(t))\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T.$
Similarly, from Eq. 2.7 and the fact $G_{h}^{\prime}(t)=\Delta_{h}G_{h}(t)$ for $t\in[0,T]$, it follows that
$\mathrm{d}z_{h}(t)=(\Delta_{h}z_{h}\!+\!z_{h}\!+\!G_{h}\!-\!P_{h}y_{h}^{3})(t)\,\mathrm{d}t+P_{h}F\big{(}y_{h}(t)\big{)}\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T.$
Defining $e_{h}:=z_{h}-P_{h}z$, we then obtain
$\mathrm{d}e_{h}(t)$ $=\Big{(}\Delta_{h}e_{h}+\Delta_{h}P_{h}z-P_{h}\Delta z+e_{h}+G_{h}-P_{h}G+P_{h}\big{(}y^{3}-y_{h}^{3}\big{)}\Big{)}(t)\,\mathrm{d}t+{}$ (4.34)
$\quad P_{h}\Big{(}F\big{(}y_{h}(t))-F\big{(}y(t)\big{)}\Big{)}\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T.$
Defining $\eta_{h}:=e_{h}-\xi_{h}$, where $\xi_{h}:=S_{0,h}(\Delta_{h}P_{h}z-P_{h}\Delta z)$, and considering the differential equation
$\Big{(}\frac{\mathrm{d}}{\mathrm{d}t}-\Delta_{h}\Big{)}\xi_{h}(t)=(\Delta_{h}P_{h}z-P_{h}\Delta z)(t),\quad t\geqslant 0,$
we use Eq. 4.34 to conclude that
$\mathrm{d}\eta_{h}(t)$ $=\Big{(}\Delta_{h}\eta_{h}+\xi_{h}+\eta_{h}+G_{h}-P_{h}G+P_{h}\big{(}y^{3}-y_{h}^{3}\big{)}\Big{)}(t)\,\mathrm{d}t+{}$ (4.35)
$\quad P_{h}\Big{(}F\big{(}y_{h}(t))-F\big{(}y(t)\big{)}\Big{)}\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T.$
By applying Itô’s formula and considering the initial condition $\eta_{h}(0)=0$ and the definition of $F$ given by Eq. 2.6, we deduce that for any $0\leqslant t\leqslant T$,
$\mathbb{E}\lVert{\eta_{h}(t)}\rVert_{L^{2}}^{p}$ $=p\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot\Big{[}\Delta_{h}\eta_{h}(s)+\xi_{h}(s)+\eta_{h}(s)$
$\qquad\qquad\qquad\qquad\qquad{}+(G_{h}-P_{h}G)(s)+\big{(}y^{3}-y_{h}^{3}\big{)}(s)\Big{]}\,\mathrm{d}\mu\,\mathrm{d}s$
$\qquad{}+\frac{p}{2}\,\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\sum_{k=0}^{\infty}\lambda_{k}\lVert{P_{h}\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}}\rVert_{L^{2}}^{2}\,\mathrm{d}s$
$\qquad{}+\frac{p(p-2)}{2}\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-4}\sum_{k=0}^{\infty}\lambda_{k}\Big{[}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}\,\mathrm{d}\mu\Big{]}^{2}\,\mathrm{d}s$
$=I_{1}+I_{2}+I_{3}+I_{4}+I_{5}+I_{6},$
where
$I_{1}$ $:=p\,\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot\Delta_{h}\eta_{h}(s)\,\mathrm{d}\mu\,\mathrm{d}s,$
$I_{2}$ $:=p\,\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot(\xi_{h}+\eta_{h}+G_{h}-P_{h}G)(s)\,\mathrm{d}\mu\,\mathrm{d}s,$
$I_{3}$ $:=p\,\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{[}y^{3}-(P_{h}y)^{3}\big{]}(s)\,\mathrm{d}\mu\,\mathrm{d}s,$
$I_{4}$ $:=p\,\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{[}(P_{h}y)^{3}-y_{h}^{3}\big{]}(s)\,\mathrm{d}\mu\,\mathrm{d}s,$
$I_{5}$ $:=\frac{p}{2}\,\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\sum_{k=0}^{\infty}\lambda_{k}\lVert{P_{h}\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}}\rVert_{L^{2}}^{2}\,\mathrm{d}s,$
$I_{6}$ $:=\frac{p(p-2)}{2}\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-4}\sum_{k=0}^{\infty}\lambda_{k}\Big{[}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}\,\mathrm{d}\mu\Big{]}^{2}\,\mathrm{d}s.$
Next, let us estimate $I_{1}$, $I_{2}$, $I_{3}$, and $I_{4}$. For the term $I_{1}$, by the definition of $\Delta_{h}$, we establish that
$I_{1}=-p\mathbb{E}\int_{0}^{t}\|{\eta_{h}(s)}\|_{L^{2}}^{p-2}\|{\eta_{h}(s)}\|_{\dot{H}^{1,2}}^{2}\,\mathrm{d}s.$ (4.36)
Regarding $I_{2}$, we can deduce that
$I_{2}$ $\leqslant\,c\mathbb{E}\int_{0}^{t}\|{\eta_{h}(s)}\|_{L^{2}}^{p-2}\big{(}\|{\xi_{h}(s)}\|_{L^{2}}^{2}+\|{\eta_{h}(s)}\|_{L^{2}}^{2}+\|{(G_{h}-P_{h}G)(s)}\|_{L^{2}}^{2}\big{)}\,\mathrm{d}s$
$\stackrel{{\scriptstyle\mathrm{(i)}}}{{\leqslant}}c\,\mathbb{E}\int_{0}^{t}\|{\xi_{h}(s)}\|_{L^{2}}^{p}+\|{\eta_{h}(s)}\|_{L^{2}}^{p}+\|{(G_{h}-P_{h}G)(s)}\|_{L^{2}}^{p}\,\mathrm{d}s$
=$\stackrel{{\scriptstyle\mathrm{(ii)}}}{{=}}c\int_{0}^{t}\|{\xi_{h}(s)}\|_{L^{p}(\Omega;L^{2})}^{p}+\|{\eta_{h}(s)}\|_{L^{p}(\Omega;L^{2})}^{p}+\|{(G_{h}-P_{h}G)(s)}\|_{L^{2}}^{p}\,\mathrm{d}s$
$\stackrel{{\scriptstyle\mathrm{(iii)}}}{{\leqslant}}ch^{2}+c\int_{0}^{t}\|{\eta_{h}(s)}\|_{L^{p}(\Omega;L^{2})}^{p}\,\mathrm{d}s,$ (4.37)
where (i) follows from Young’s inequality, (ii) uses the determinism of $G$ and $G_{h}$, and (iii) is justified by the inequalities (4.25) and (4.27). For $I_{3}$, by employing Hölder’s inequality, Young’s inequality, and the stability property of $P_{h}$ as stated in Eq. 4.2, a straightforward calculation yields
$I_{3}$ $\leqslant c\,\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-1}\lVert{(y-P_{h}y)(s)}\rVert_{L^{4}}\Big{(}\lVert{y(s)}\rVert_{L^{8}}^{2}+\lVert{P_{h}y(s)}\rVert_{L^{8}}^{2}\Big{)}\,\mathrm{d}s$
$\leqslant c\,\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-1}\lVert{(y-P_{h}y)(s)}\rVert_{L^{4}}\lVert{y(s)}\rVert_{L^{8}}^{2}\,\mathrm{d}s$
$\leqslant c\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p}+\lVert{(y-P_{h}y)(s)}\rVert_{L^{4}}^{p}\lVert{y(s)}\rVert_{L^{8}}^{2p}\,\mathrm{d}s.$
By utilizing Eq. 4.29, we arrive at the following bound:
$I_{3}\leqslant ch^{2-\epsilon}+c\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{p}(\Omega;L^{2})}^{p}\,\mathrm{d}s,\quad\forall\epsilon>0.$ (4.38)
To estimate $I_{4}$, we proceed as follows. Using the definitions in (4.32) and (4.33), along with the relation $z_{h}-P_{h}z=\xi_{h}+\eta_{h}$, we deduce that $P_{h}y-y_{h}=-\xi_{h}-\eta_{h}+P_{h}G-G_{h}$. For any $s\in[0,t]$, we can then write:
$\int_{\mathcal{O}}\eta_{h}(s)\left[(P_{h}y)^{3}-y_{h}^{3}\right](s)\,\mathrm{d}\mu$
[…]
Hence, applying Hölder’s inequality, Young’s inequality, and the stability property of $P_{h}$ from (4.2), we obtain the following bound for $I_{4}$:
$I_{4}$ $\leqslant c\mathbb{E}\int_{0}^{t}\bigg{[}\|\eta_{h}(s)\|_{L^{2}}^{p}+\left(\|\xi_{h}(s)\|^{p}_{L^{2\beta^{\prime}}}+\|(P_{h}G-G_{h})(s)\|^{p}_{L^{2\beta^{\prime}}}\right)$
×$\qquad\qquad\qquad{}\times\Big{(}\|y(s)\|^{p}_{L^{2\beta}}+\|y_{h}(s)\|^{p}_{L^{2\beta}}\Big{)}\bigg{]}\,\mathrm{d}s,\quad\forall\beta\in\big{(}1,\frac{3p}{3p-4}\big{)}.$
Using (4.30), we further refine this bound to:
$I_{4}\leqslant c\int_{0}^{t}\|\eta_{h}(s)\|^{p}_{L^{p}(\Omega;L^{2})}\,\mathrm{d}s+ch^{2-\frac{3}{2}p\left(1-\frac{1}{\beta}\right)},\quad\forall\beta\in\left(1,\frac{3p}{3p-4}\right).$
By selecting $\beta$ sufficiently small within $\big{(}1,\frac{3p}{3p-4}\big{)}$ for any $\epsilon>0$, we arrive at the final bound for $I_{4}$:
$I_{4}\leqslant ch^{2-\epsilon}+c\int_{0}^{t}\|\eta_{h}(s)\|^{p}_{L^{p}(\Omega;L^{2})}\,\mathrm{d}s.$ (4.39)
Now, let us proceed to estimate $I_{5}+I_{6}$. Using Hölder’s inequality and the stability property of $P_{h}$ from Eq. 4.2, we obtain the following bound:
$I_{5}+I_{6}\leqslant c\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\sum_{k=0}^{\infty}\lambda_{k}\lVert{f_{k}(y_{h}(s))-f_{k}(y(s))}\rVert_{L^{2}}^{2}\,\mathrm{d}s.$
Invoking the Lipschitz continuity condition Eq. 2.3 and the summability condition Eq. 2.5, we further deduce that
$I_{5}+I_{6}\leqslant c\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\lVert{(y_{h}-y)(s)}\rVert_{L^{2}}^{2}\,\mathrm{d}s.$
Subsequently, by Young’s inequality, we obtain
$I_{5}+I_{6}\leqslant c\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p}+\lVert{(y_{h}-y)(s)}\rVert_{L^{2}}^{p}\,\mathrm{d}s.$
Recalling the definitions in Eqs. 4.32 and 4.33 and considering the relation $z_{h}-P_{h}z=\xi_{h}+\eta_{h}$, we infer that
$(y_{h}-y)(s)=\xi_{h}(s)+\eta_{h}(s)-(z-P_{h}z)(s)-(G-G_{h})(s),\quad\forall s\in[0,T],$ (4.40)
which, in conjunction with Eq. 4.25, Eq. 4.26, and Eq. 4.27, yields
$\mathbb{E}\int_{0}^{t}\lVert{(y_{h}-y)(s)}\rVert_{L^{2}}^{p}\,\mathrm{d}s\leqslant ch^{2}+\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{p}(\Omega;L^{2})}^{p}\,\mathrm{d}s.$
Consequently, we obtain the following bound for $I_{5}+I_{6}$:
$I_{5}+I_{6}\leqslant ch^{2}+c\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{p}(\Omega;L^{2})}^{p}\,\mathrm{d}s.$ (4.41)
Finally, by integrating the estimates for $I_{1}$, $I_{2}$, $I_{3}$, $I_{4}$, and $I_{5}+I_{6}$, as given by Eqs. 4.36, 4.37, 4.38, 4.39 and 4.41, we obtain the following inequality:
$\|\eta_{h}(t)\|_{L^{p}(\Omega;L^{2})}^{p}+\mathbb{E}\int_{0}^{t}\|\eta_{h}(s)\|_{L^{2}}^{p-2}\|\eta_{h}(s)\|_{\dot{H}_{h}^{1,2}}^{2}\,\mathrm{d}s$
[…]
Applying Gronwall’s lemma to the above inequality, we infer that
$\max_{t\in[0,T]}\|\eta_{h}(t)\|_{L^{p}(\Omega;L^{2})}+\left[\mathbb{E}\int_{0}^{T}\|\eta_{h}(t)\|_{L^{2}}^{p-2}\|\eta_{h}(t)\|_{\dot{H}_{h}^{1,2}}^{2}\,\mathrm{d}t\right]^{1/p}\leqslant ch^{\alpha/p}.$
Moreover, for any $\epsilon>0$ and $t\in(0,T]$, we establish the following bounds:
$\lVert{(z-P_{h}z)(t)}\rVert_{L^{p}(\Omega;L^{2})}\leqslant ch^{1-\epsilon}$ (by Eqs. 4.3 and 3.4),
[…]
Since the equality Eq. 4.40 implies
$\|(y-y_{h})(t)\|_{L^{p}(\Omega;L^{2})}$ $\leqslant\|(z-P_{h}z)(t)\|_{L^{p}(\Omega;L^{2})}+\|(G-G_{h})(t)\|_{L^{p}(\Omega;L^{2})}$
$\qquad{}+\|\xi_{h}(t)\|_{L^{p}(\Omega;L^{2})}+\|\eta_{h}(t)\|_{L^{p}(\Omega;L^{2})},\quad\forall t\in[0,T],$
the combination of these bounds yields the desired error estimate Eq. 3.6. This completes the proof of Theorem 3.2.
$\blacksquare$
Proof of Theorem 3.3. Let $G$ be defined by Eq. 2.11, and let $z$ be defined by Eq. 4.32. Extending the reasoning from the proof of Theorem 3.1, we establish the following enhanced regularity properties for $y$ and $z$:
$y\in L^{p}(\Omega;C([0,T];\dot{H}^{2-\epsilon,q})),$ $\forall p\in(2,\infty),\,\forall q\in[2,\infty),\,\forall\epsilon>0;$ (4.42)
$z\in L^{p}(\Omega\times(0,T);\dot{H}^{2,q}),$ $\forall p\in(2,\infty),\,\forall q\in[2,\infty);$ (4.43)
$z\in L^{p}(\Omega;C([0,T];\dot{H}^{2-\epsilon,q})),$ $\forall p\in(2,\infty),\,\forall q\in[2,\infty),\,\forall\epsilon>0.$ (4.44)
Let $G_{h}$ be defined by Eq. 2.12 and let $z_{h}$ be defined by Eq. 4.33. We adopt the same notation introduced in the proof of Theorem 3.2:
$e_{h}$ $:=z_{h}-P_{h}z,$
$\xi_{h}$ $:=S_{0,h}(\Delta_{h}P_{h}-P_{h}\Delta)z,$
$\eta_{h}$ $:=e_{h}-\xi_{h}.$
It is straightforward to confirm the following decomposition for $y-y_{h}$:
$y-y_{h}=(z-P_{h}z)+(G-G_{h})-(\xi_{h}+\eta_{h}).$ (4.45)
The remainder of the proof is structured into the following five steps.
Step 1. Let us present some key estimates that form the foundation of our subsequent analysis. We begin by establishing the following estimates:
$\left\|z-P_{h}z\right\|_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2},\quad\forall p\in(2,\infty),\,\forall q\in[2,\infty),$ (4.46)
$\left\|G-G_{h}\right\|_{C([0,T];L^{q})}+\left\|P_{h}G-G_{h}\right\|_{C([0,T];L^{q})}\leqslant ch^{2},\quad\forall q\in[2,\infty).$ (4.47)
Estimate (4.46) is a direct consequence of (4.3) combined with the regularity property (4.43), and estimate (4.47) follows from (4.6) and the condition $v\in W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$. Analogous to the estimate Eq. 4.27, we employ Lemma 4.5, along with (4.4) and the regularity assertion (4.43), to bound $\xi_{h}$ as follows:
$\left\|\xi_{h}\right\|_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2},\quad\forall p\in(2,\infty),\,\forall q\in[2,\infty).$ (4.48)
Adopting the reasoning from Theorem 3.2, we establish the following bound for $\eta_{h}$:
$\max_{t\in[0,T]}\left\|\eta_{h}(t)\right\|_{L^{p}(\Omega;L^{2})}\leqslant ch^{2},\quad\forall p\in(2,\infty).$ (4.49)
In light of the decomposition Eq. 4.45, combining Eqs. 4.48, 4.49, 4.47 and 4.46 leads to the error estimate
$\left\|y-y_{h}\right\|_{L^{p}(\Omega\times(0,T);L^{2})}\leqslant ch^{2},\quad\forall p\in(2,\infty).$ (4.50)
Furthermore, for any $p\in(2,\infty)$ and $q\in[2,\infty)$, we deduce the following chain of inequalities:
$\lVert{y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}$
[…]
$\stackrel{{\scriptstyle\mathrm{(i)}}}{{\leqslant}}{}$ $\lVert{P_{h}y}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}+ch^{3(1/q-1/2)}\lVert{P_{h}(y-y_{h})}\rVert_{L^{p}(\Omega\times(0,T);L^{2})}$
$\stackrel{{\scriptstyle\mathrm{(ii)}}}{{\leqslant}}{}$ $c\lVert{y}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}+ch^{3(1/q-1/2)}\lVert{y-y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{2})}$
$\stackrel{{\scriptstyle\mathrm{(iii)}}}{{\leqslant}}{}$ $c\lVert{y}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}+ch^{2+3(1/q-1/2)},$
where in (i) we apply the inverse estimate $\lVert{v_{h}}\rVert_{L^{q}}\leqslant ch^{3(1/q-1/2)}\lVert{v_{h}}\rVert_{L^{2}}$ for all $v_{h}\in X_{h}$; in (ii) we utilize the stability property of $P_{h}$ as stated in Eq. 4.2; and in (iii) we employ the error estimate Eq. 4.50. By invoking the regularity assertion Eq. 4.42, we subsequently conclude a uniform boundedness with respect to the spatial mesh size $h$:
$\lVert{y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\text{ is uniformly bounded w.r.t~{}$h$ for all $p,q\in[2,\infty)$}.$ (4.51)
Step 2. Let us prove
$\lVert{\eta_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{6})}\leqslant ch^{2},\quad\forall p\in(2,\infty).$ (4.52)
Using the equality Eq. 4.35, along with the definitions of the operators $S_{0,h}$ and $S_{1,h}$, we deduce the following decomposition:
$\eta_{h}=I_{1}+I_{2}+I_{3},$ (4.53)
where
$I_{1}$ $:=S_{0,h}\big{(}\xi_{h}+\eta_{h}+G_{h}-P_{h}G\big{)},$
$I_{2}$ $:=S_{0,h}P_{h}(y^{3}-y_{h}^{3}),$
$I_{3}$ $:=S_{1,h}P_{h}\big{(}F(y_{h})-F(y)\big{)}.$
Fix an arbitrary $p\in(2,\infty)$. For the term $I_{1}$, we establish the following bound for any $q\in[2,\infty)$:
$\lVert{I_{1}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}$ $\leqslant c\lVert{S_{0,h}\big{(}\xi_{h}+\eta_{h}+G_{h}-P_{h}G\big{)}}\rVert_{L^{p}(\Omega\times(0,T);\dot{H}_{h}^{2,2})}$
$\leqslant c\lVert{\xi_{h}+\eta_{h}+G_{h}-P_{h}G}\rVert_{L^{p}(\Omega\times(0,T);L^{2})}$
$\leqslant ch^{2},$
where the first inequality uses the discrete Sobolev’s inequality $\lVert{v_{h}}\rVert_{L^{q}}\leqslant c\lVert{v_{h}}\rVert_{\dot{H}_{h}^{2,2}}$ for all $v_{h}\in X_{h}$; the subsequent inequality relies on Lemma 4.5; and the final inequality is justified by Eqs. 4.49, 4.47 and 4.48. Turning to $I_{2}$, we establish the bound: for any $q\in[2,\infty)$,
$\lVert{I_{2}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}$ $\leqslant c\lVert{S_{0,h}P_{h}(y^{3}-y_{h}^{3})}\rVert_{L^{p}(\Omega\times(0,T);\dot{H}_{h}^{2,5/3})}$
$\leqslant c\lVert{y^{3}-y_{h}^{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{5/3})}$
$\leqslant c\lVert{y-y_{h}}\rVert_{L^{2p}(\Omega\times(0,T);L^{2})}\Big{(}\lVert{y}\rVert_{L^{4p}(\Omega\times(0,T);L^{20})}^{2}+\lVert{y_{h}}\rVert_{L^{4p}(\Omega\times(0,T);L^{20})}^{2}\Big{)}$
$\leqslant ch^{2},$
where the first inequality uses the discrete Sobolev’s inequality $\lVert{v_{h}}\rVert_{L^{q}}\leqslant c\lVert{v_{h}}\rVert_{\dot{H}_{h}^{2,5/3}}$ for all $v_{h}\in X_{h}$; the second inequality uses Lemma 4.5 and Eq. 4.2; the third inequality uses Hölder’s ineqality; and the last inequality uses Eqs. 4.42, 4.50 and 4.51. For $I_{3}$ we demonstrate the following bound:
$\lVert{I_{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{6})}$
[…]
where the first inequality employs the discrete Sobolev’s inequality $\lVert{v_{h}}\rVert_{L^{6}}\leqslant c\lVert{v_{h}}\rVert_{\dot{H}_{h}^{1,2}}$ for all $v_{h}\in X_{h}$, and the second inequality invokes Lemma 4.6, along with Eqs. 4.2 and 4.10. Given that the decomposition Eq. 4.53 implies
$\lVert{\eta_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{6})}$ $\leqslant\lVert{I_{1}}\rVert_{L^{p}(\Omega\times(0,T);L^{6})}+\lVert{I_{2}}\rVert_{L^{p}(\Omega\times(0,T);L^{6})}$
$\qquad{}+\lVert{I_{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{6})},$
combining the above bounds for $I_{1}$, $I_{2}$ and $I_{3}$ yields Eq. 4.52.
Step 3. Let us proceed to prove the error estimate Eq. 3.7. Fix any $p\in(2,\infty)$ and $q\in[2,\infty)$. We start by establishing the following error estimate:
$\lVert{y-y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{6})}$ $\leqslant\lVert{z-P_{h}z}\rVert_{L^{p}(\Omega\times(0,T);L^{6})}+\lVert{G-G_{h}}\rVert_{L^{p}(0,T;L^{6})}+{}$
[…]
Employing this error estimate and following the proof methodology for bounding $I_{3}$ in Step 2, we further obtain an enhanced bound for $I_{3}$:
$\lVert{I_{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2}.$
Integrating this refined bound with those previously derived bounds for $I_{1}$ and $I_{2}$ in Step 2, and considering the decomposition in Eq. 4.53, we conclude that
$\lVert{\eta_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2}.$ (4.54)
Furthermore, we infer that
$\lVert{y-y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}$ $\leqslant\lVert{z-P_{h}z}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}+\lVert{G-G_{h}}\rVert_{L^{p}(0,T;L^{q})}+{}$
[…]
Consequently, combining these estimates directly yields the desired error estimate Eq. 3.7.
Step 4. Let us demonstrate the following inequality:
$\lVert{\eta_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2},\quad\forall p\in(2,\infty),\,\forall q\in[2,\infty).$ (4.55)
To this end, let $p\in(2,\infty)$ and $q\in[2,\infty)$ be arbitrarily chosen. Invoking the definitions of $I_{1}$ and $I_{2}$ from (4.53), in conjunction with Lemma 4.10 and the continuous embedding of $\dot{H}_{h}^{\varepsilon,q}$ into $L^{q}$ for any $\varepsilon>0$, we deduce
$\lVert{I_{1}+I_{2}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant c\lVert{\xi_{h}+\eta_{h}+G_{h}-P_{h}G+P_{h}(y^{3}-y_{h}^{3})}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}.$
Relying on (4.47), (4.48), and (4.54), it follows that
$\lVert{I_{1}+I_{2}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2}+c\lVert{P_{h}(y^{3}-y_{h}^{3})}\rVert_{L^{p}(\Omega\times(0,T);L^{q})},$
which, combined with the stability property of $P_{h}$ as stated in (4.2), leads to
$\lVert{I_{1}+I_{2}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2}+c\lVert{y^{3}-y_{h}^{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}.$
Moreover, using Hölder’s inequality yields
$\lVert{y^{3}-y_{h}^{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}$
[…]
which, in light of the regularity result (4.42) and the stability assertion (4.51), simplifies to
$\lVert{y^{3}-y_{h}^{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant c\lVert{y-y_{h}}\rVert_{L^{2p}(\Omega\times(0,T);L^{2q})}.$
Consequently, we obtain the following bound for $I_{1}+I_{2}$:
$\lVert{I_{1}+I_{2}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2}+c\lVert{y-y_{h}}\rVert_{L^{2p}(\Omega\times(0,T);L^{2q})}.$
We proceed to bound the term $I_{3}$ from Eq. 4.53 as follows:
$\lVert{I_{3}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}$ $\leqslant c\lVert{P_{h}(F(y_{h})-F(y))}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}$
$\leqslant c\lVert{F(y_{h})-F(y)}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}$
$\leqslant c\lVert{y-y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})},$
where the first inequality is justified by Lemma 4.6 and the continuous embedding of $(\dot{H}_{h}^{0,q},\dot{H}_{h}^{2,q})_{1/2-1/p,p}$ in $L^{q}$; the second inequality is a consequence of the stability of the projection $P_{h}$ from Eq. 4.2; and the final inequality relies on the Lipschitz continuity of $F$ as stated in Eq. 4.10. Combining the above bounds for $I_{1}+I_{2}$ and $I_{3}$, we arrive at
$\lVert{I_{1}+I_{2}+I_{3}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}$
[…]
Given the decomposition Eq. 4.53 and the derived error estimate Eq. 3.7 for all $p\in(2,\infty)$ and $q\in[2,\infty)$, the desired inequality Eq. 4.55 is thereby established.
Step 5. We proceed to prove the error estimate Eq. 3.8. Fix arbitrarily $p\in(2,\infty)$, $q\in[2,\infty)$, and $\epsilon>0$. Using the regularity assertion Eq. 4.43 and following the proof technique used for Eq. 4.28, we deduce the following bound for $\xi_{h}$:
$\lVert{\xi_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2-\epsilon}.$
Employing Eq. 4.3 and the regularity assertion Eq. 4.44, we establish the following bound for $z-P_{h}z$:
$\|z-P_{h}z\|_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2-\epsilon}.$
Furthermore, combining Eqs. 4.47 and 4.55 leads to the following bound:
$\|G-G_{h}+\eta_{h}\|_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2}.$
Given that the decomposition Eq. 4.45 implies
$\|y-y_{h}\|_{L^{p}(\Omega;C([0,T];L^{q}))}$ $\leqslant\|\xi_{h}\|_{L^{p}(\Omega;C([0,T];L^{q}))}+\|z-P_{h}z\|_{L^{p}(\Omega;C([0,T];L^{q}))}$
$\qquad{}+\|G-G_{h}+\eta_{h}\|_{L^{p}(\Omega;C([0,T];L^{q}))},$
the desired error estimate Eq. 3.8 is derived by combining the above bounds. This concludes the proof of Theorem 3.3.
$\blacksquare$
5 Conclusions
In this work, we have rigorously investigated the regularity properties of the three-dimensional stochastic Allen-Cahn equation with multiplicative noise in the presence of non-smooth initial data. For the spatial semidiscretization using the finite element method, we have established a convergence rate of $O(h^{\alpha/p}+h^{2/p}t^{-1/p})$ with non-smooth initial data. Meanwhile, under the presence of smooth initial data, employing the discrete stochastic maximal $L^{p}$-regularity theory, we have achieved a convergence rate of $O(h^{2})$ in the norm of $L^{p}(\Omega\times(0,T);L^{q})$, alongside a pathwise uniform convergence rate of $O(h^{2-\epsilon})$. Significantly, these findings are inclusive of diverse spatial $L^{q}$-norms, enhancing their adaptability and universality. The derived theoretical results set a foundation for the numerical analysis of fully discrete approximations based on the finite element method for the three-dimensional stochastic Allen-Cahn equation with multiplicative noise.

### v7

@@
 1 Introduction
-Let $\mathcal{O}$ be a bounded convex polyhedral domain in $\mathbb{R}^{3}$, and let $\Delta$ be the Laplace operator with homogeneous Dirichlet boundary condition.
+Let $\mathcal{O}$ be a bounded convex domain in $\mathbb{R}^{3}$ with a smooth boundary $\partial\mathcal{O}$, and let $\Delta$ denote the Laplace operator with homogeneous Dirichlet boundary conditions.
 We consider the following stochastic Allen-Cahn equation:
-$\mathrm{d}y(t)=(\Delta y+y-y^{3})(t)\,\mathrm{d}t+F(y(t))\,\mathrm{d}W_{H}(t),\quad 0<t\leqslant T,$ (1.1a)
+$\mathrm{d}y(t)=(\Delta y+y-y^{3})(t)\,\mathrm{d}t+F(y(t))\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T,$ (1.1a)
 $y(0)=v,$ (1.1b)
-where the initial value $v$ is given and $W_{H}$ is an infinite-dimensional Wiener process.
+where the initial value $v$ is given, and $W_{H}$ is an infinite-dimensional $H$-cylindrical Brownian motion.
 The stochastic Allen-Cahn equations are widely used to model phase transitions in materials science; see e.g., [4, 10, 15, 27, 33].
@@
 The main motivation of this work includes two components:
-  • To the best of our knowledge, the existing works on the numerical analysis of the three-dimensional stochastic Allen-Cahn equation have not adequately addressed scenarios involving non-smooth initial data.
-This scenario is of practical significance, as initial phases may encompass randomly distributed particles.
-The inclusion of non-smooth initial data introduces heightened complexity to numerical analysis when compared to cases with smooth initial data.
+  • To the best of our knowledge, existing works on the numerical analysis of the three-dimensional stochastic Allen-Cahn equation have not adequately addressed scenarios involving non-smooth initial data.
+This scenario is of practical significance, as initial phases may include randomly distributed particles.
+The inclusion of non-smooth initial data introduces heightened complexity to numerical analysis compared to cases with smooth initial data.
 Firstly, the theoretical analysis of the three-dimensional stochastic Allen-Cahn equation with multiplicative noise and non-smooth initial data is relatively scarce (see, e.g., [18, 19]).
@@
   • In the existing literature, the numerical analysis of the three-dimensional stochastic Allen-Cahn equation with multiplicative noise primarily focuses on the spatial $L^{2}$-norm.
-However, error estimates employing the more general spatial $L^{q}$-norms are relatively scarce.
-Accordingly, the main contributions of this work are as follows:
+However, error estimates using the more general spatial $L^{q}$-norms are relatively scarce.
+The main contributions of this work are as follows:
   • For the model problem Eq.
-1.1 with non-smooth initial data $v\in L^{\infty}$, we derive some regularity estimates within the framework of the general spatial $L^{q}$-norms, which are pivotal for the numerical analysis.
+1.1 with non-smooth initial data $v\in L^{\infty}$, we derive some regularity estimates within the framework of the general spatial $L^{q}$-norms, which are crucial for the numerical analysis.
 Specifically, we affirm that
@@
     $\lVert{(y-y_{h})(t)}\rVert_{L^{p}(\Omega;L^{2})}\leqslant c\big{(}h^{\alpha/p}+h^{2/p}t^{-1/p}\big{)},\quad\forall t\in(0,T],$
-    where $p\in(2,\infty)$ and $\alpha\in(1,2)$ can be arbitrarily chosen. This result is a stochastic counterpart of the well-known error estimate in [28, Theoerem 3.2].
+    where $p\in(2,\infty)$ and $\alpha\in(1,2)$ can be arbitrarily chosen. This result is a stochastic counterpart of the well-known error estimate in [35, Theoerem 3.2].
   • Given smooth initial data $v\in W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$, we mainly use the standard theory of finite element methods (see, e.g., [1, 35]), the discrete maximal $L^{p}$-regularity [16], and the discrete stochastic maximal $L^{p}$-regularity [17] to derive two rigorous error estimates within the framework of general spatial $L^{q}$-norms.
@@
     $\lVert{y-y_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2-\epsilon}.$
-To achieve a second-order spatial convergence rate, the previous numerical analysis in the literature typically enforce stringent regularity assumptions on the diffusion term.
-However, in this study, we impose only some moderate regularity conditions on $F$ (see Section 2).
-Furthermore, this work introduces a novel approach to the numerical analysis of finite element method-based spatial semidiscretizations for a broad class of nonlinear stochastic parabolic equations.
-The rest of this paper is organized as follows.
-Section 2 introduces some conventions and definitions used in this paper, and the spatial semidiscretization analyzed in this research.
-Section 3 presents the main results of this paper, including the regularity of Eq.
-1.1 with non-smooth initial data and the convergence of the spatial semidiscretization under both non-smooth and smooth initial conditions.
-Section 4 provides a rigorous proof of the aforementioned results.
-Lastly, Section 5 summarizes the key findings of this work.
-The rest of this paper is organized as follows.
-Section 2 establishes the foundational notation, definitions, and preliminaries that underpin the analysis within this study, as well as detailing the spatial semidiscretization that is the focus of our investigation.
-Section 3 is dedicated to the principal findings of this work, which encompass the treatment of the regularity of the solution to (1.1) with nonsmooth initial data, along with the convergence analysis of the spatial semidiscretization scheme under both nonsmooth and smooth initial conditions.
-Section 4 furnishes a detailed and exacting proof of the results previously outlined.
-Finally, Section 5 offers a concise synthesis of the pivotal insights and contributions that this research brings to the field.
+To achieve a second-order spatial convergence rate, prior numerical analyses in the literature have typically enforced stringent regularity assumptions on the diffusion coefficients.
+In contrast, this study imposes relatively moderate regularity conditions on the diffusion coefficient $F$ (see Section 2).
+While our numerical analysis assumes deterministic initial data, extending the results to accommodate random initial data is straightforward.
+The remainder of this paper is organized as follows.
+Section 2 introduces the notational conventions and definitions that are used throughout the paper.
+Section 3 presents the main results of this work, including the regularity properties of the mild solution to model problem 1.1 with non-smooth initial data, and establishes the convergence of the finite element-based spatial semidiscretization for both non-smooth and smooth initial conditions.
+A rigorous proof of these results is provided in Section 4.
+Finally, Section 5 concludes the paper by summarizing the key findings of this study.
 2 Preliminaries
 Notational Conventions.
-Throughout this paper, we will adhere to the following notational conventions.
-For Banach spaces $E_{1}$ and $E_{2}$, we denote the space of all bounded linear operators from $E_{1}$ to $E_{2}$ as $\mathcal{L}(E_{1},E_{2})$, and it is abbreviated to $\mathcal{L}(E_{1})$ when $E_{1}=E_{2}$.
-The identity map is denoted by $I$.
-For any $\theta\in(0,1)$ and $p\in(1,\infty)$, we use $(\cdot,\cdot)_{\theta,p}$ to represent the real interpolation space constructed by the $K$-method, as described in [21, Chapter 1].
-The conjugate exponent of $p$ is denoted by $p^{\prime}$.
-For $q\in[1,\infty]$, we briefly write $L^{q}(\mathcal{O})$ as $L^{q}$.
-Let $W_{0}^{1,\infty}(\mathcal{O})$ and $W^{2,\infty}(\mathcal{O})$ denote two standard Sobolev spaces (see, e.g., [14]).
-Let $(\Omega,\mathcal{F},\mathbb{P})$ represent a complete probability space equipped with a filtration $\mathbb{F}:=(\mathcal{F}_{t})_{t\geqslant 0}$, on which a sequence of independent $\mathbb{F}$-adapted Brownian motions $(\beta_{n})_{n\in\mathbb{N}}$ is defined.
+Throughout this paper, we will use the following conventions.
+For Banach spaces $E_{1}$ and $E_{2}$, the space of all bounded linear operators from $E_{1}$ to $E_{2}$ is denoted by $\mathcal{L}(E_{1},E_{2})$, and we abbreviate this to $\mathcal{L}(E_{1})$ when $E_{1}=E_{2}$.
+The identity operator is indicated by $I$.
+Given $\theta\in(0,1)$ and $p\in(1,\infty)$, the notation $(\cdot,\cdot)_{\theta,p}$ is used to refer to the real interpolation space constructed via the $K$-method, as detailed in Chapter 1 of [22].
+The conjugate exponent of each $p\in[0,\infty]$ is denoted by $p^{\prime}$.
+The symbol $\mathcal{O}$ represents a bounded convex domain in $\mathbb{R}^{3}$ with a boundary $\partial\mathcal{O}$ that belongs to the $C^{2,0}$ class (see Section 6.2 of [11]).
+For $q\in[1,\infty]$, we abbreviate $L^{q}(\mathcal{O})$ to $L^{q}$, and use $W_{0}^{1,q}(\mathcal{O})$ and $W^{2,q}(\mathcal{O})$ to denote two standard Sobolev spaces (see, e.g., [34]).
+Let $(\Omega,\mathcal{F},\mathbb{P})$ represent a complete probability space equipped with a right-continuous filtration $\mathbb{F}:=(\mathcal{F}_{t})_{t\geqslant 0}$, on which a sequence of independent $\mathbb{F}$-adapted Brownian motions $(\beta_{n})_{n\in\mathbb{N}}$ is defined.
 The expectation operator associated with this probability space is denoted by $\mathbb{E}$.
-For $p\in(1,\infty)$ and a Banach space $X$, we use $L^{p}_{\mathbb{F}}(\Omega\times(0,T);X)$ to denote the space of all $\mathbb{F}$-adapted processes in $L^{p}(\Omega\times(0,T);X)$.
-For any Banach space $X$, the notation $\gamma(H,X)$ denotes the space of all $\gamma$-radonifying operators from $H$ to $X$ (see [11, Chapter 9]).
-Finally, for each $t\in\mathbb{R}_{+}$, define $W_{H}(t)\in\mathcal{L}(H,L^{2}(\Omega))$ by
-$W_{H}(t)h:=\sum_{n\in\mathbb{N}}\beta_{n}(t)(h,h_{n})_{H},\quad\forall h\in H,$
-where $H$ is a separable Hilbert space with inner product $(\cdot,\cdot)_{H}$ and an orthonormal basis $(h_{n})_{n\in\mathbb{N}}$.
-Definition of Operator $F$.
-Let $(f_{n})_{n\in\mathbb{N}}$ be a sequence of continuous functions mapping from $\overline{\mathcal{O}}\times\mathbb{R}$ to $\mathbb{R}$, satisfying the following conditions for every $n\in\mathbb{N}$:
+Let $H$ be a separable Hilbert space with inner product $(\cdot,\cdot)_{H}$ and an orthonormal basis $(h_{n})_{n\in\mathbb{N}}$.
+The $H$-cylindrical Brownian motion $W_{H}$ is defined such that, for each $t\in\mathbb{R}_{+}$, $W_{H}(t)\in\mathcal{L}(H,L^{2}(\Omega))$ is defined by
+$W_{H}(t)h:=\sum_{n\in\mathbb{N}}\beta_{n}(t)(h,h_{n})_{H},\quad\forall h\in H.$
+For $p\in(1,\infty)$ and a Banach space $X$, the space of all $\mathbb{F}$-adapted processes in $L^{p}(\Omega\times(0,T);X)$ is denoted by $L^{p}_{\mathbb{F}}(\Omega\times(0,T);X)$.
+For any Banach space $X$, the notation $\gamma(H,X)$ represents the space of all $\gamma$-radonifying operators from $H$ to $X$, as defined in [13, Chapter 9].
+For a comprehensive treatment of the theory of stochastic integration with respect to the $H$-cylindrical Wiener process $W_{H}$, we refer the reader to [24].
+Definition of the Operator $F$.
+Let $(f_{n})_{n\in\mathbb{N}}$ be a sequence of continuous functions mapping from $\overline{\mathcal{O}}\times\mathbb{R}$ to $\mathbb{R}$, which satisfy the following conditions for every $n\in\mathbb{N}$:
 $f_{n}(x,0)=0,$ $\forall x\in\partial\mathcal{O},$ (2.1)
@@
 $f_{n}(v):=f_{n}(\cdot,v(\cdot)),\quad\forall v\in L^{1}.$
-Let $\{\lambda_{n}\}_{n\in\mathbb{N}}$ be a sequence of non-negative numbers satisfying
+Let $(\lambda_{n})_{n\in\mathbb{N}}$ be a sequence of non-negative numbers satisfying the summability condition
 $\sum_{n\in\mathbb{N}}\lambda_{n}<\infty.$ (2.5)
-Given $v\in L^{2}$, we define $F(v)\in\gamma(H,L^{2})$ via
+For any $v\in L^{2}$, we define $F(v)\in\gamma(H,L^{2})$ as
 $F(v):=\sum_{n=0}^{\infty}\sqrt{\lambda}_{n}h_{n}\otimes f_{n}(v).$ (2.6)
-Consequently, the model equation Eq. 1.1 can be explicitly stated as follows:
-$\begin{cases}\mathrm{d}y(t)=(\Delta y+y-y^{3})(t)\,\mathrm{d}t+\sum_{n=0}^{\infty}\sqrt{\lambda_{n}}f_{n}(y(t))\,\mathrm{d}\beta_{n}(t),\quad 0<t\leqslant T,\\ y(0)=v.\end{cases}$
+Accordingly, the model equation (1.1) can be explicitly formulated as follows:
+$\begin{cases}\mathrm{d}y(t)=(\Delta y+y-y^{3})(t)\,\mathrm{d}t+\sum_{n=0}^{\infty}\sqrt{\lambda_{n}}f_{n}(y(t))\,\mathrm{d}\beta_{n}(t),\quad 0\leqslant t\leqslant T,\\ y(0)=v.\end{cases}$
+Mild Solutions. Let $p,q\in[2,\infty)$. We introduce the deterministic convolution operator $S_{0}$ acting on functions $g\in L^{p}(0,T;L^{q})$ as follows:
+$(S_{0}g)(t):=\int_{0}^{t}e^{(t-s)\Delta}g(s)\,\mathrm{d}s,\quad\text{for }0\leqslant t\leqslant T.$ (2.7)
+Additionally, for any process $g\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,L^{q}))$, we define the stochastic convolution operator $S_{1}$ by
+$(S_{1}g)(t):=\int_{0}^{t}e^{(t-s)\Delta}g(s)\,\mathrm{d}W_{H}(s),\quad\text{for }0\leqslant t\leqslant T.$ (2.8)
+Here, $e^{t\Delta}$, for $t\in[0,\infty)$, denotes the analytic semigroup generated by the Laplace operator $\Delta$ subject to homogeneous Dirichlet boundary conditions.
+A process $y\in L_{\mathbb{F}}^{6}(\Omega\times(0,T);L^{6})$ is called a mild solution to equation (1.1) if it satisfies the equality
+$y=G+S_{0}(y-y^{3})+S_{1}F(y),$ (2.9)
+where the function $G$ is given by
+$G(t):=e^{t\Delta}v,\quad\text{for }t\in[0,T].$ (2.10)
+Note that the conditions in the above definition of a mild solution can be relaxed (see, e.g., [29, Chapter 6]). However, the present formulation is adequate for the purposes of this study.
 Spatial Semidiscretization.
-Let $\mathcal{K}_{h}$ represent a conforming and quasi-uniform triangulation of the domain $\mathcal{O}$, consisting of three-dimensional simplexes.
-We define the finite element space $X_{h}$ as
-$X_{h}:=\left\{v_{h}\in C(\overline{\mathcal{O}})\,\middle|\,v_{h}|_{K}\in P_{r}(K)\text{ for each }K\in\mathcal{K}_{h}\text{ and }v_{h}\text{ vanishes on }\partial\mathcal{O}\right\},$
-where $C(\overline{\mathcal{O}})$ is the space of continuous functions over the closure of $\mathcal{O}$, $r\geqslant 1$ is an integer denoting the polynomial degree, and $P_{r}(K)$ is the space of polynomials defined on simplex $K$ up to degree $r$.
-Let $h$ denote the maximum diameter of the elements in $\mathcal{K}_{h}$.
-Let $P_{h}$ be the $L^{2}$-orthogonal projection operator onto $X_{h}$, and we define the discrete Laplace operator $\Delta_{h}:X_{h}\to X_{h}$ by
+Consider $\mathcal{K}_{h}$ to be a conforming and quasi-uniform triangulation of the domain $\mathcal{O}$, composed of three-dimensional simplexes.
+We introduce the finite element space $X_{h}$ as
+$X_{h}:=\Bigl{\{}v_{h}\in C(\overline{\mathcal{O}})\bigm{|}v_{h}=0\text{ on }(\mathcal{O}\setminus\mathcal{O}_{h})\cup\partial\mathcal{O}_{h},\text{ and }v_{h}|_{K}\in P_{r}(K)\text{ for each }K\in\mathcal{K}_{h}\Bigr{\}},$
+where $C(\overline{\mathcal{O}})$ denotes the space of continuous functions on the closure of $\mathcal{O}$, $\mathcal{O}_{h}$ is the closure of the union of the simplexes in $\mathcal{K}_{h}$, $r\geqslant 1$ is a positive integer, and $P_{r}(K)$ is the space of polynomials on simplex $K$ up to degree $r$.
+Let the spatial mesh size $h$ be the maximum diameter of the elements in $\mathcal{K}_{h}$.
+Define $P_{h}$ as the $L^{2}$-orthogonal projection operator onto $X_{h}$.
+The discrete Laplace operator $\Delta_{h}:X_{h}\to X_{h}$ is defined by
 $\int_{\mathcal{O}}(\Delta_{h}u_{h})\cdot v_{h}\,\mathrm{d}\mu=-\int_{\mathcal{O}}\nabla u_{h}\cdot\nabla v_{h}\,\mathrm{d}\mu,\quad\text{for all }u_{h},v_{h}\in X_{h},$
-where $\mu$ represents the three-dimensional Lebesgue measure on $\mathcal{O}$. The spatial semidiscretization considered in this work is given by
-$\mathrm{d}y_{h}(t)=\left(\Delta_{h}y_{h}+y_{h}-P_{h}(y_{h}^{3})\right)(t)\mathrm{d}t+P_{h}F(y_{h}(t))\,\mathrm{d}W_{H}(t),\quad t>0,$ (2.7a)
+where $\mu$ denotes the three-dimensional Lebesgue measure on $\mathcal{O}$. We consider the following spatial semidiscretization:
+$\mathrm{d}y_{h}(t)=\left(\Delta_{h}y_{h}+y_{h}-P_{h}(y_{h}^{3})\right)(t)\,\mathrm{d}t+P_{h}F(y_{h}(t))\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T,$ (2.11a)
 $y_{h}(0)=P_{h}v,$ (2.11b)
 where $v$ is the given initial value.
-Mild Solutions. Let $p\in(1,\infty)$ and $q\in[2,\infty)$. We define the convolution operator $S_{0}$ acting on any function $g\in L^{p}(0,T;L^{q})$ as follows:
-$(S_{0}g)(t):=\int_{0}^{t}e^{(t-s)\Delta}g(s)\,\mathrm{d}s,\quad 0\leqslant t\leqslant T.$ (2.8)
-Similarly, for any $g\in L_{\mathbb{F}}^{p}(\Omega;L^{2}(0,T;\gamma(H,L^{q})))$, we introduce the stochastic convolution operator $S_{1}$ defined by
-$(S_{1}g)(t):=\int_{0}^{t}e^{(t-s)\Delta}g(s)\,\mathrm{d}W_{H}(s),\quad 0\leqslant t\leqslant T.$ (2.9)
-An $\mathbb{F}$-adapted process $y$ is called a mild solution to Eq. 1.1 if $y\in L_{\mathbb{F}}^{6}(\Omega\times(0,T);L^{6})$ and satisfies
-$y=G+S_{0}(y-y^{3})+S_{1}F(y),$ (2.10)
-where $G$ is defined as
-$G(t):=e^{t\Delta}v,\quad t\in[0,T].$ (2.11)
-In the context of the spatial semidiscretization, for any $g_{h}\in L^{p}(0,T;X_{h})$, we define
-$(S_{0,h}g_{h})(t):=\int_{0}^{t}e^{(t-s)\Delta_{h}}g_{h}(t)\,\mathrm{d}t,\quad t\in[0,T].$
-For any $g_{h}\in L_{\mathbb{F}}^{p}(\Omega;L^{2}(0,T;\gamma(H,X_{h})))$, the discrete stochastic convolution operator $S_{1,h}$ is given by
-$(S_{1,h}g_{h})(t):=\int_{0}^{t}e^{(t-s)\Delta_{h}}g_{h}(t)\,\mathrm{d}W_{H}(t),\quad t\in[0,T].$
-An $\mathbb{F}$-adapted $X_{h}$-valued stochastic process $y_{h}$ is called a mild solution to the spatial semidiscretization Eq.
-2.7 if $y_{h}\in L_{\mathbb{F}}^{6}(\Omega\times(0,T);X_{h})$ and it satisfies that
-$y_{h}=G_{h}+S_{0,h}(y_{h}-P_{h}y_{h}^{3})+S_{1,h}P_{h}F(y_{h}),$
-where $G_{h}$ is given by
-$G_{h}(t):=e^{t\Delta_{h}}P_{h}v,\quad t\in[0,T].$ (2.12)
+Assume that $v\in L^{2}$.
+Since this spatial semidiscretization is a finite-dimensional stochastic differential equation (SDE), we can use the classical theory of finite-dimensional SDEs (see, e.g., [28, Theorem 3.21]) to conclude that Eq.
+2.11 has a unique strong solution.
 3 Main results
-In the remainder of this paper, the symbol $c$ represents a generic positive constant that depends solely on $T$, $\mathcal{O}$, $\sum_{n\in\mathbb{N}}\lambda_{h}$, the regularity parameters of the triangulation $\mathcal{K}_{h}$, the indices of the pertinent Lebesgue, Sobolev, and interpolation spaces, the initial value $v$ in Eqs.
-1.1 and 2.7, and the exponents related to the spatial mesh size $h$.
-The value of $c$ may differ from line to line.
-In addition, for any $\alpha\in[0,\infty)$ and $q\in(1,\infty)$, let $\dot{H}^{\alpha,q}$ denote the domain of $(-\Delta)^{\alpha/2}$ in $L^{q}$, endowed with the standard norm defined as follows:
-$\lVert{v}\rVert_{\dot{H}^{\alpha,q}}:=\lVert{(-\Delta)^{\alpha/2}v}\rVert_{L^{q}},\quad\forall v\in\dot{H}^{\alpha,q}.$
+In the remainder of this paper, the symbol $c$ denotes a generic positive constant that depends only on $T$, $\mathcal{O}$, $\sum_{n\in\mathbb{N}}\lambda_{n}$, the constant $C_{F}$ in Eqs.
+2.1, 2.2, 2.4 and 2.3, the indices of the relevant Lebesgue, Sobolev, and interpolation spaces, the initial value $v$ given in equations Eq.
+1.1 and Eq.
+2.11, the regularity parameters of the triangulation $\mathcal{K}_{h}$, and the exponents associated with the spatial mesh size $h$.
+The value of $c$ may differ at each occurrence.
+For any $\alpha\in[0,\infty)$ and $q\in(1,\infty)$, let $\dot{H}^{\alpha,q}$ denote the domain of $(-\Delta)^{\alpha/2}$ in $L^{q}$, endowed with the standard norm defined as follows:
+$\|v\|_{\dot{H}^{\alpha,q}}:=\|(-\Delta)^{\alpha/2}v\|_{L^{q}}\quad\text{for all $v\in\dot{H}^{\alpha,q}$}.$
+The dual space of $\dot{H}^{\alpha,q}$ is denoted by $\dot{H}^{-\alpha,q^{\prime}}$.
+It is well-established that $\Delta$ has a unique extension as a bounded linear operator from $\dot{H}^{\alpha,q}$ to $\dot{H}^{\alpha-2,q}$ for all $\alpha\in[0,2)$.
+Specifically, when $\alpha=2$, the space $\dot{H}^{2,q}$ is identical to the intersection $W_{0}^{1,q}(\mathcal{O})\cap W^{2,q}(\mathcal{O})$ and is endowed with a norm equivalent to that of $W^{2,q}(\mathcal{O})$ (see [11, Theorem 9.15]).
+Moreover, for $\alpha=1$, the space $\dot{H}^{1,q}$ coincides with the standard Sobolev space $W_{0}^{1,q}(\mathcal{O})$, and the two norms are equivalent (see [38, Theorem 16.14]).
 The main results of this paper are summarized in the following three theorems.
@@
 Assume that $v\in L^{\infty}$.
-Let $y$ and $y_{h}$ be the mild solutions of Eqs.
-1.1 and 2.7, respectively.
+Let $y$ be the mild solution of Eq.
+1.1, and let $y_{h}$ be the strong solution to the spatial semidiscretization Eq.
+2.11.
 Then, for any $\alpha\in(1,2)$ and $p\in(2,\infty)$, the following error estimate holds for all $t\in(0,T]$:
@@
 Assume that $v\in W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$.
-Let $y$ and $y_{h}$ be the mild solutions of Eqs.
-1.1 and 2.7, respectively.
+Let $y$ be the mild solution of Eq.
+1.1, and let $y_{h}$ be the strong solution of Eq.
+2.11.
 Then, for any $p\in(2,\infty)$, $q\in[2,\infty)$, and $\epsilon>0$, the following error estimates hold:
@@
 $\lVert{y-y_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2-\epsilon}.$ (3.8)
+It is worth noting that the aforementioned theoretical results remain valid under the assumption that the operator $F$ possesses the growth and Lipschitz continuity properties outlined in Lemma 4.5.
 Remark 3.1.
-For Theorem 3.1, we have the following remarks.
-  1. 1.
-    Liu [17] proposed a theoretical framework for a broad spectrum of nonlinear stochastic partial differential equations. Applying this framework to our model yields the result that
+Liu [18] proposed a theoretical framework for a broad spectrum of nonlinear stochastic partial differential equations. Applying this framework to our model yields
 $y\in L^{p}(\Omega;C([0,T];\dot{H}^{1,2})),\quad\forall p\in(2,\infty),$
 provided that $v\in\dot{H}^{1,2}$.
-However, the theory in [17] does not cover the case of non-smooth initial data.
-In the context of the stochastic Allen-Cahn equation with additive noise, Bréhier et al.
-[4, Lemma 3.1] provide some a priori estimates with the general spatial $L^{q}$ norms.
+Furthermore, Liu and Qiao [21] have established certain regularity estimates in the context of the spatial $L^{2}$-norm, and some regularity estimates with the spatial $L^{\infty}$-norm were derived by using Sobolev’s embedding theorem.
+Regarding the stochastic Allen-Cahn equation with additive noise, Bréhier et al.
+[5, Lemma 3.1] provided some a priori estimates involving general spatial $L^{q}$ norms.
+Remark 3.2.
+For Theorems 3.2 and 3.3, we have the following comments.
+  1. 1.
+    In the numerical analysis of the three-dimensional stochastic Allen-Cahn equation with multiplicative noise, the existing literature often imposes strict regularity conditions on the diffusion coefficients.
+For instance, the study in [23] demands that the diffusion coefficients possess bounded second-order derivatives.
+Our choice of diffusion coefficient $F$ adheres to the criteria outlined in Assumption 2.2 of [21]; however, it does not meet the more stringent requirements stated in Assumption 2.3 of the same reference.
+Investigating the numerical approximations of Eq.
+1.1 under less restrictive conditions on $F$, especially the loosening of condition (2.1), has important theoretical and practical implications.
   2. 2.
-    Assume that the condition Eq.
-2.1 is removed.
-By an approximation argument we conclude that Eq.
-3.1 remains valid.
-Moreover, for any $p\in(2,\infty)$ and $q\in[2,\infty)$, the following assertions hold:
-    $y-G\in L^{p}(\Omega\times(0,T);\dot{H}^{1,q}),$
-    $y-G\in L^{p}(\Omega;C([0,T];(L^{q},\dot{H}^{2,q})_{1/2-1/p,p})).$
-Remark 3.2.
-For Theorems 3.2 and 3.3, we have the following remarks.
-  1. 1.
-    In the numerical analysis of the three-dimensional stochastic Allen-Cahn equation with multiplicative infinite-dimensional Wiener process, the existing literature typically imposes stringent regularity conditions on the diffusion coefficient.
-Our chosen formulation of the function $F$ adheres to the criteria specified in Assumption 2.2 of Liu and Qiao’s work [20]; however, it fails to satisfy the more stringent requirements imposed by their Assumption 2.3.
-Examining the numerical approximations of Eq.
-1.1 under less restrictive conditions on $F$, particularly the relaxation of condition (2.1), holds significant theoretical and practical implications.
-  2. 2.
-    In Section 4.3, we demonstrate that the norm of $y_{h}$ in $L^{p}(\Omega\times(0,T);L^{q})$ is uniformly bounded with respect to the spatial mesh size $h$ for all $p,q\in[2,\infty)$, provided that $v$ is an element of $W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$ (see Eq.
-4.51).
-It can be rigorously shown that this result remains valid even if $v$ is an element of $W_{0}^{\epsilon,\infty}(\mathcal{O})$, where $\epsilon>0$ can be any arbitrarily small positive number.
-This conclusion is crucial for the numerical analysis of the fully discrete approximations of Eq.
+    In Section 4.3, we show that the norm of $y_{h}$ in $L^{p}(\Omega\times(0,T);L^{q})$ is uniformly bounded with respect to the spatial mesh size $h$ for all $p,q\in[2,\infty)$, assuming $v\in W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$ (see Eq.
+4.53).
+It can be rigorously demonstrated that this result also holds when $v\in W_{0}^{\epsilon,\infty}(\mathcal{O})$, where $\epsilon>0$ can be any arbitrarily small positive number.
+This finding is critical for the numerical analysis of the fully discrete approximations of Eq.
 1.1.
-However, the question of whether this conclusion extends to the case where $v$ is an element of $L^{\infty}$ remains unresolved.
+However, it remains an open question whether this conclusion applies when $v\in L^{\infty}$.
 4 Proofs
@@
 4.1 Preliminary Lemmas
+It is a well-established result that for any $q\in(1,\infty)$ and $s\in(0,2)$, the space $\dot{H}^{s,q}$ coincides with the complex interpolation space $[L^{q},\dot{H}^{2,q}]_{s/2}$, equipped with equivalent norms.
+This result is documented, for example, in Chapter 16 of [38].
+The combination of Proposition 1.4 from [22] and Theorem C.4.1 in [12] yields the following continuous embeddings.
 Lemma 4.1.
@@
   • (i)
-    $\dot{H}^{\gamma,q}$ is continuously embedded in $(L^{q},\dot{H}^{1,q})_{\beta,q}$;
+    The space $\dot{H}^{\gamma,q}$ is continuously embedded in $(L^{q},\dot{H}^{1,q})_{\beta,q}$;
   • (ii)
-    $(L^{q},\dot{H}^{1,q})_{\gamma,q}$ is continuously embedded in $\dot{H}^{\beta,q}$.
+    The space $(L^{q},\dot{H}^{1,q})_{\gamma,q}$ is continuously embedded in $\dot{H}^{\beta,q}$.
+Theorem 2.12 and inequality (2.128) in [38] provide the following standard estimate for the growth of the analytic semigroup generated by the Laplace operator with homogeneous Dirichlet boundary conditions.
 Lemma 4.2.
-Let $0\leqslant\alpha\leqslant\beta<\infty$ and $q\in(1,\infty)$. Then for all $t\in(0,T]$ we have
+Let $0\leqslant\alpha\leqslant\beta<\infty$ and $q\in(1,\infty)$. For all $t\in(0,T]$, the following estimate holds:
 $\lVert{e^{t\Delta}}\rVert_{\mathcal{L}(\dot{H}^{\alpha,q},\dot{H}^{\beta,q})}\leqslant ct^{(\alpha-\beta)/2}.$
+For the deterministic convolution operator $S_{0}$, defined by Eq.
+2.7, we have the following celebrated deterministic maximal $L^{p}$-regularity estimate; see Theorem 6.1.11 in [30] and Corollary 1.14 from [22].
 Lemma 4.3.
-Let $p,q\in(1,\infty)$. Then for any $g\in L^{p}(0,T;L^{q})$, the following maximal $L^{p}$-regularity estimate holds:
-$\Big{\lVert}{\frac{\mathrm{d}}{\mathrm{d}t}S_{0}g}\Big{\rVert}_{L^{p}(0,T;L^{q})}+\lVert{S_{0}g}\rVert_{L^{p}(0,T;\dot{H}^{2,q})}+\lVert{S_{0}g}\rVert_{C([0,T];(L^{q},\dot{H}^{2,q})_{1-1/p,p})}$
-[…]
+Let $p,q\in(1,\infty)$. Then for any $g\in L^{p}(0,T;L^{q})$, the following inequality holds:
+$\big{\lVert}\frac{\mathrm{d}}{\mathrm{d}t}S_{0}g\big{\rVert}_{L^{p}(0,T;L^{q})}+\lVert{S_{0}g}\rVert_{L^{p}(0,T;\dot{H}^{2,q})}+\lVert{S_{0}g}\rVert_{C([0,T];(L^{q},\dot{H}^{2,q})_{1-1/p,p})}\leqslant c\lVert{g}\rVert_{L^{p}(0,T;L^{q})}.$
+For the stochastic convolution operator $S_{1}$, defined by Eq. 2.8, we present the following fundamental stochastic maximal $L^{p}$-regularity estimate; see Theorems 1.1 and 1.2 in [26].
 Lemma 4.4.
-Let $p\in(2,\infty)$ and $q\in[2,\infty)$. Then for any $g\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,L^{q}))$, the following stochastic maximal $L^{p}$-regularity estimate holds:
-$\lVert{S_{1}g}\rVert_{L^{p}(\Omega;C([0,T];(L^{q},\dot{H}^{2,q})_{1/2-1/p,p}))}+\lVert{S_{1}g}\rVert_{L^{p}(\Omega\times(0,T);\dot{H}^{1,q})}$
-[…]
-For any $\alpha\in\mathbb{R}$ and $q\in(1,\infty)$, let $\dot{H}_{h}^{\alpha,q}$ denote the space $X_{h}$ equipped with the norm
-$\lVert{v_{h}}\rVert_{\dot{H}_{h}^{\alpha,q}}:=\lVert{(-\Delta_{h})^{\alpha/2}v_{h}}\rVert_{L^{q}},\quad\forall v_{h}\in X_{h}.$
+Let $p\in(2,\infty)$ and $q\in[2,\infty)$. Then for any $g\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,L^{q}))$, the following inequality holds:
+$\lVert{S_{1}g}\rVert_{L^{p}(\Omega;C([0,T];(L^{q},\dot{H}^{2,q})_{1/2-1/p,p}))}+\lVert{S_{1}g}\rVert_{L^{p}(\Omega\times(0,T);\dot{H}^{1,q})}\leqslant c\lVert{g}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}.$
+The operator $F$, defined by Eq. 2.6, possesses the following growth and Lipschitz continuity properties.
 Lemma 4.5.
-Let $p,q\in(1,\infty)$. Then for any $g_{h}\in L^{p}(0,T;X_{h})$, the following discrete maximal $L^{p}$-regularity estimate holds:
-$\Big{\lVert}{\frac{\mathrm{d}}{\mathrm{d}t}S_{0,h}g_{h}}\Big{\rVert}_{L^{p}(0,T;\dot{H}_{h}^{0,q})}+\lVert{S_{0,h}g_{h}}\rVert_{L^{p}(0,T;\dot{H}_{h}^{2,q})}\\ \leqslant c\lVert{g_{h}}\rVert_{L^{p}(0,T;\dot{H}_{h}^{0,q})}.$
-Lemma 4.6.
-Let $p\in(2,\infty)$ and $q\in[2,\infty)$. Then for any $g_{h}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,X_{h}))$, the following discrete stochastic maximal $L^{p}$-regularity estimate holds:
-$\lVert{S_{1,h}g_{h}}\rVert_{L^{p}(\Omega;C([0,T];(\dot{H}_{h}^{0,q},\dot{H}_{h}^{2,q})_{1/2-1/p,p}))}+\lVert{S_{1,h}g_{h}}\rVert_{L^{p}(\Omega\times(0,T);\dot{H}_{h}^{1,q})}$ (4.1)
-[…]
-Lemma 4.7.
-The following assertions are valid:
-  1. (i)
-    For all $q\in[1,\infty]$, it holds that
-    $\lVert{P_{h}}\rVert_{\mathcal{L}(L^{q})}\text{ is uniformly bounded with respect to $h$}.$ (4.2)
-  2. (ii)
-    For all $\alpha\in[0,2]$ and $q\in(1,\infty)$, $P_{h}$ has the following approximation property:
-    $\lVert{I-P_{h}}\rVert_{\mathcal{L}(\dot{H}^{\alpha,q},L^{q})}\leqslant ch^{\alpha}.$ (4.3)
-  3. (iii)
-    For all $0\leqslant\alpha<\beta\leqslant 2$ and $q\in(1,\infty)$, the following inequality holds:
-    $\lVert{\Delta_{h}P_{h}-P_{h}\Delta}\rVert_{\mathcal{L}(\dot{H}^{\beta,q},\dot{H}_{h}^{\alpha-2,q})}\leqslant ch^{\beta-\alpha}.$ (4.4)
-  4. (iv)
-    For all $\alpha\in[0,2]$, $q\in(1,\infty)$, and $t>0$, the following inequalities are satisfied:
-    $\lVert{e^{t\Delta_{h}}}\rVert_{\mathcal{L}(\dot{H}_{h}^{0,q},\dot{H}_{h}^{\alpha,q})}\leqslant ct^{-\alpha/2},$ (4.5)
-    $\lVert{e^{t\Delta_{h}}P_{h}-P_{h}e^{t\Delta}}\rVert_{\mathcal{L}(\dot{H}^{\alpha,q},L^{q})}+\lVert{(I-P_{h})e^{t\Delta}}\rVert_{\mathcal{L}(\dot{H}^{\alpha,q},L^{q})}$
-    $\leqslant c\min\Big{\{}h^{\alpha},\frac{h^{2}}{t^{1-\alpha/2}}\Big{\}}.$ (4.6)
-Remark 4.1.
-Lemma 4.1 is established by Proposition 1.4 from [21] and Theorem C.4.1 in [10].
-The inequality asserted in Lemma 4.2 is a direct consequence of Theorem 2.12 and Equation (2.128) as presented in [33].
-The proof of Lemma 4.3 is provided in [32] and Corollary 1.14 from [21].
-The proof of Lemma 4.4 is detailed in [30].
-The proof of Lemma 4.5 is accessible in [15], and the proof of Lemma 4.6 can be found in [16].
-The assertions in Lemma 4.7 are widely recognized and can be found in standard references such as [6, 28].
-Lemma 4.8.
-For any $q\in(1,\infty)$, the following inequalities hold:
-$\lVert{F(v)}\rVert_{\gamma(H,L^{q})}\leqslant c(1+\lVert{v}\rVert_{L^{q}}),$ $\forall v\in L^{q};$ (4.7)
-$\lVert{F(v)}\rVert_{\gamma(H,\dot{H}^{1,q})}\leqslant c(1+\lVert{v}\rVert_{\dot{H}^{1,q}}),$ $\forall v\in\dot{H}^{1,q};$ (4.8)
-$\lVert{F(v)}\rVert_{\gamma(H,\dot{H}^{s_{1},q})}\leqslant c(1+\lVert{v}\rVert_{\dot{H}^{s_{2},q}}),$ $\forall v\in\dot{H}^{s_{2},q},\,\forall 0\leqslant s_{1}<s_{2}<1;$ (4.9)
-$\lVert{F(u)-F(v)}\rVert_{\gamma(H,L^{q})}\leqslant c\lVert{u-v}\rVert_{L^{q}},$ $\forall u,v\in L^{q}.$ (4.10)
+Let $q\in(1,\infty)$. The following inequalities are valid:
+$\lVert{F(v)}\rVert_{\gamma(H,L^{q})}\leqslant c\left(1+\lVert{v}\rVert_{L^{q}}\right),$ for all $\text{for all }v\in L^{q};$ (4.1)
+$\lVert{F(v)}\rVert_{\gamma(H,\dot{H}^{1,q})}\leqslant c\left(1+\lVert{v}\rVert_{\dot{H}^{1,q}}\right),$ for all $\text{for all }v\in\dot{H}^{1,q};$ (4.2)
+$\lVert{F(v)}\rVert_{\gamma(H,\dot{H}^{s_{1},q})}\leqslant c\left(1+\lVert{v}\rVert_{\dot{H}^{s_{2},q}}\right),$ for $\text{for all }v\in\dot{H}^{s_{2},q},\text{ where }0\leqslant s_{1}<s_{2}<1;$ (4.3)
+$\lVert{F(u)-F(v)}\rVert_{\gamma(H,L^{q})}\leqslant c\lVert{u-v}\rVert_{L^{q}},$ for all $\text{for all }u,v\in L^{q}.$ (4.4)
 Proof.
-We only prove Eq.
-4.9, the proof of Eqs.
-4.7, 4.8 and 4.10 being easier.
-Let $\Omega_{\gamma}$ denote a probability space on which a sequence $(\gamma_{n})_{n\in\mathbb{N}}$ of independent standard Gaussian random variables is defined, and let $\mathbb{E}_{\gamma}$ denote the expectation operator with respect to $\Omega_{\gamma}$.
-The rest of the proof is divided into two steps.
+We focus our demonstration on Eq.
+4.3; the proofs of Eq.
+4.1, Eq.
+4.2, and Eq.
+4.4 follow similar, yet more straightforward, lines of reasoning.
+For the definition of $\gamma$-radonifying operators, we direct the reader to Section 9.1 of [13].
+Consider $\Omega_{\gamma}$ as a probability space supporting a sequence $(\gamma_{n})_{n\in\mathbb{N}}$ of independent standard Gaussian random variables.
+Let $\mathbb{E}_{\gamma}$ denote the expectation operator with respect to the probability space $\Omega_{\gamma}$.
+The subsequent argument is structured in two main steps.
 Step 1. We proceed to demonstrate that for any $\theta\in(0,1)$ and $N\in\mathbb{N}_{>0}$, the following inequality holds:
@@
 $A_{N}v:=\sum_{n=0}^{N}\gamma_{n}\sqrt{\lambda_{n}}f_{n}(v),\quad\forall v\in L^{q}.$
-For any $v\in\dot{H}^{1,q}$, we infer that
-$\lVert{A_{N}v}\rVert_{L^{q}(\Omega_{\gamma};\dot{H}^{1,q})}$ $\leqslant c\Big{(}\mathbb{E}_{\gamma}\int_{\mathcal{O}}\Big{\lvert}{\sum_{n=0}^{N}\gamma_{n}\sqrt{\lambda_{n}}\nabla f_{n}(v)}\Big{\rvert}^{q}\,\mathrm{d}\mu\Big{)}^{1/q}$
-$\stackrel{{\scriptstyle\mathrm{(i)}}}{{\leqslant}}c\Big{(}\int_{\mathcal{O}}\Big{(}\mathbb{E}_{\gamma}\Big{\lvert}{\sum_{n=0}^{N}\gamma_{n}\sqrt{\lambda_{n}}\nabla f_{n}(v)}\Big{\rvert}^{2}\Big{)}^{q/2}\,\mathrm{d}\mu\Big{)}^{1/q}$
-$=c\Big{(}\int_{\mathcal{O}}\Big{(}\sum_{n=0}^{N}\lambda_{n}\lvert{\nabla f_{n}(v)}\rvert^{2}\Big{)}^{q/2}\,\mathrm{d}\mu\Big{)}^{1/q}$
-$\stackrel{{\scriptstyle\mathrm{(ii)}}}{{\leqslant}}c\Big{(}\int_{\mathcal{O}}\Big{(}\sum_{n=0}^{N}\lambda_{n}(1+\lvert{v}\rvert+\lvert{\nabla v}\rvert)^{2}\Big{)}^{q/2}\,\mathrm{d}\mu\Big{)}^{1/q},$
-where step $(\mathrm{i})$ invokes Theorem 6.2.6 from [11], and step $(\mathrm{ii})$ relies on Lipschitz conditions Eqs. 2.3 and 2.4. Using the summability condition Eq. 2.5 and the standard inequality
+For any $v\in\dot{H}^{1,q}$, we deduce that
+$\lVert{A_{N}v}\rVert_{L^{q}(\Omega_{\gamma};\dot{H}^{1,q})}$ $\leqslant c\bigg{[}\mathbb{E}_{\gamma}\int_{\mathcal{O}}\Big{\lvert}{\sum_{n=0}^{N}\gamma_{n}\sqrt{\lambda_{n}}\nabla f_{n}(v)}\Big{\rvert}^{q}\,\mathrm{d}\mu\bigg{]}^{1/q}$
+$\stackrel{{\scriptstyle\mathrm{(i)}}}{{\leqslant}}c\bigg{[}\int_{\mathcal{O}}\Big{(}\mathbb{E}_{\gamma}\Big{\lvert}{\sum_{n=0}^{N}\gamma_{n}\sqrt{\lambda_{n}}\nabla f_{n}(v)}\Big{\rvert}^{2}\Big{)}^{q/2}\,\mathrm{d}\mu\bigg{]}^{1/q}$
+$=c\bigg{[}\int_{\mathcal{O}}\Big{(}\sum_{n=0}^{N}\lambda_{n}\lvert{\nabla f_{n}(v)}\rvert^{2}\Big{)}^{q/2}\,\mathrm{d}\mu\bigg{]}^{1/q}$
+$\stackrel{{\scriptstyle\mathrm{(ii)}}}{{\leqslant}}c\bigg{[}\int_{\mathcal{O}}\Big{(}\sum_{n=0}^{N}\lambda_{n}(1+\lvert{v}\rvert+\lvert{\nabla v}\rvert)^{2}\Big{)}^{q/2}\,\mathrm{d}\mu\bigg{]}^{1/q},$
+where step $(\mathrm{i})$ uses the famous Kahane-Khintchine inequality (see Theorem 6.2.6 from [13]), and step $(\mathrm{ii})$ relies on Lipschitz conditions Eqs.
+2.3 and 2.4.
+Using the summability condition Eq.
+2.5 and the standard inequality
 $\left(\int_{\mathcal{O}}|v|^{q}+|\nabla v|^{q}\,\mathrm{d}\mu\right)^{1/q}\leqslant c\lVert{v}\rVert_{\dot{H}^{1,q}},$
@@
 $\lVert{A_{N}v}\rVert_{L^{q}(\Omega_{\gamma};\dot{H}^{1,q})}\leqslant c(1+\lVert{v}\rVert_{\dot{H}^{1,q}}).$
-Analogously, employing Eq. 2.2 yields
+Analogously, employing the growth condition Eq. 2.2 yields
 $\lVert{A_{N}v}\rVert_{L^{q}(\Omega_{\gamma};L^{q})}\leqslant c(1+\lVert{v}\rVert_{L^{q}}),\quad\forall v\in L^{q},$
-and utilizing Eq. 2.3 results in
+and utilizing the Lipschitz continuous condition Eq. 2.3 results in
 $\lVert{A_{N}u-A_{N}v}\rVert_{L^{q}(\Omega_{\gamma};L^{q})}\leqslant c\lVert{u-v}\rVert_{L^{q}},\quad\forall u,v\in L^{q}.$
@@
 $\lVert{A_{N}v}\rVert_{(L^{q}(\Omega_{\gamma};L^{q}),L^{q}(\Omega_{\gamma};\dot{H}^{1,q}))_{\theta,q}}\leqslant c(1+\lVert{v}\rVert_{(L^{q},\dot{H}^{1,q})_{\theta,q}}),\quad\forall v\in(L^{q},\dot{H}^{1,q})_{\theta,q}.$
-This result, when combined with the continuous embedding property
+Combining this result with the continuous embedding property
 $\left(L^{q}(\Omega_{\gamma};L^{q}),L^{q}(\Omega_{\gamma};\dot{H}^{1,q})\right)_{\theta,q}\hookrightarrow L^{q}\left(\Omega_{\gamma};(L^{q},\dot{H}^{1,q})_{\theta,q}\right),$
-yields
+which can be easily verified by definition, yields
 $\lVert{A_{N}v}\rVert_{L^{q}(\Omega_{\gamma};(L^{q},\dot{H}^{1,q})_{\theta,q})}\leqslant c\big{(}1+\lVert{v}\rVert_{(L^{q},\dot{H}^{1,q})_{\theta,q}}\big{)},\quad\forall v\in(L^{q},\dot{H}^{1,q})_{\theta,q}.$
-Applying Theorem 6.2.6 from [11] once again, we arrive at the assertion in Eq. 4.11.
+Applying the Kahane-Khintchine inequality once again, we arrive at the assertion in Eq. 4.5.
 Step 2. For any $\theta\in(0,1)$ and $v\in(L^{q},\dot{H}^{1,q})_{\theta,q}$, we consider the equality
-$\lVert{F(v)}\rVert_{\gamma(H,(L^{q},\dot{H}^{1,q})_{\theta,q})}=\lVert{\lim_{N\to\infty}A_{N}v}\rVert_{L^{2}(\Omega_{\gamma};(L^{q},\dot{H}^{1,q})_{\theta,q})},$
-which stems directly from the definition of $F$ as stated in (2.6) and is justified by Theorem 9.1.19 in [11]. By invoking the inequality (4.11), we apply Fatou’s Lemma to deduce that
+$\lVert{F(v)}\rVert_{\gamma(H,(L^{q},\dot{H}^{1,q})_{\theta,q})}=\Big{\lVert}\lim_{N\to\infty}A_{N}v\Big{\rVert}_{L^{2}(\Omega_{\gamma};(L^{q},\dot{H}^{1,q})_{\theta,q})},$
+which follows directly from the definition of $F$ as stated in (2.6) and is justified by Theorem 9.1.19 in [13]. Using the inequality (4.5), we apply Fatou’s Lemma to deduce that
 $\lVert{F(v)}\rVert_{\gamma(H,(L^{q},\dot{H}^{1,q})_{\theta,q})}\leqslant c(1+\lVert{v}\rVert_{(L^{q},\dot{H}^{1,q})_{\theta,q}}).$
@@
 Remark 4.1.
-As established by Theorem 4.36 in [21], the interpolation space $(L^{2},\dot{H}^{1,q})_{s,2}$ coincides with $\dot{H}^{s,2}$, sharing equivalent norms across the interval $s\in(0,1)$.
-Consequently, we refine the estimate Eq.
+As established by Theorem 4.36 in [22], the interpolation space $(L^{2},\dot{H}^{1,2})_{s,2}$ coincides with $\dot{H}^{s,2}$, sharing equivalent norms for all $s\in(0,1)$.
+Consequently, we refine the inequality Eq.
 4.3 to the following form:
-$\|F(v)\|_{\gamma(H,\dot{H}^{s,2})}\leqslant c(1+\|v\|_{\dot{H}^{s,2}}),\quad\forall v\in\dot{H}^{s,2},\,\forall s\in(0,1).$
+$\|F(v)\|_{\gamma(H,\dot{H}^{s,2})}\leqslant c(1+\|v\|_{\dot{H}^{s,2}}),\quad\text{for all $v\in\dot{H}^{s,2}$, where $s\in(0,1)$}.$
 4.2 Proof of Theorem 3.1
-It is a well-established fact that the space $L^{2}$ admits an orthonormal basis $\{\phi_{n}\}_{n\in\mathbb{N}}$, comprised of eigenvectors of the negative Laplace operator $-\Delta$.
-Furthermore, this basis $\{\phi_{n}\}_{n\in\mathbb{N}}$ forms a complete orthogonal system within the space $\dot{H}^{s,2}$ for all $s\in(0,\infty)$.
+It is well known that the space $L^{2}$ admits an orthonormal basis $(\phi_{n})_{n\in\mathbb{N}}$, consisting of eigenvectors of the negative Laplace operator $-\Delta$ with homogeneous Dirichlet boundary conditions.
+This basis $(\phi_{n})_{n\in\mathbb{N}}$ forms a complete orthogonal system in the space $\dot{H}^{s,2}$ for all $s\in(0,\infty)$.
 Lemma 4.6.
 For each $n\in\mathbb{N}$ and $v_{n}\in\mathrm{span}\{\phi_{0},\ldots,\phi_{n}\}$, the stochastic Allen-Cahn equation
-$\mathrm{d}y_{n}(t)=\big{(}\Delta y_{n}+y_{n}-y_{n}^{3}\big{)}(t)\,\mathrm{d}t+F(y_{n}(t))\,\mathrm{d}W_{H}(t),\quad 0<t\leqslant T,$ (4.12a)
+$\mathrm{d}y_{n}(t)=\big{(}\Delta y_{n}+y_{n}-y_{n}^{3}\big{)}(t)\,\mathrm{d}t+F(y_{n}(t))\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T,$ (4.6a)
 $y_{n}(0)=v_{n}$ (4.6b)
@@
 We split the proof into the following two steps.
-Step 1. Let $G_{n}(t):=e^{t\Delta}v_{n}$, $t\in[0,T]$. Since $v_{n}\in\dot{H}^{2,q}$ for each $q\in(2,\infty)$, by Lemma 4.2 with $\alpha=\beta=2$, we have
+Step 1. Define $G_{n}(t):=e^{t\Delta}v_{n}$ for $t\in[0,T]$. Given that $v_{n}\in\dot{H}^{2,q}$ for every $q\in(2,\infty)$, it follows from Lemma 4.2 with $\alpha=\beta=2$ that
 $G_{n}\in L^{p}(0,T;\dot{H}^{2,q}),\quad\forall p,q\in(2,\infty).$ (4.8)
-In view of Eq. 4.8 and the fact that $v_{n}\in\dot{H}^{1,2}$, we use [17, Theorem 1.1] to deduce that Eq. 4.12 has a unique mild solution $y_{n}$ and
+In view of Eq. 4.2 and the fact that $v_{n}\in\dot{H}^{1,2}$, we use [18, Theorem 1.1] (see also, [21, Lemma 2.3 and Theorem 3.1]) to deduce that Eq. 4.6 has a unique mild solution $y_{n}$ and
 $y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{1,2}),\quad\forall p\in(2,\infty).$ (4.9)
@@
 $y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{6}),\quad\forall p\in(2,\infty),$
-which implies $y_{n}-y_{n}^{3}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{2})$ for each $p\in(2,\infty)$. Hence, using Lemma 4.3 gives
+which in turn implies that $y_{n}-y_{n}^{3}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{2})$ for all $p\in(2,\infty)$. Applying Lemma 4.3, we obtain
 $S_{0}(y_{n}-y_{n}^{3})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,2}),\quad\forall p\in(2,\infty).$ (4.10)
-Since Eqs. 4.8 and 4.15 imply that $F(y_{n})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,\dot{H}^{1,2}))$ for each $p\in(2,\infty)$, by Lemma 4.4 we obtain
+Since (4.2) and (4.9) ensure that $F(y_{n})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,\dot{H}^{1,2}))$ for every $p\in(2,\infty)$, Lemma 4.4 allows us to conclude that
 $S_{1}F(y_{n})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,2}),\quad\forall p\in(2,\infty).$ (4.11)
 Considering the mild solution representation
-$y_{n}=G_{n}+S_{0}(y_{n}-y_{n}^{3})+S_{1}F(y_{n}),$ (4.18)
-and combining Eqs. 4.14, 4.16 and 4.17, we conclude that
+$y_{n}=G_{n}+S_{0}(y_{n}-y_{n}^{3})+S_{1}F(y_{n}),$
+and combining (4.8), (4.10), and (4.11), we deduce that
 $y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,2}),\quad\forall p\in(2,\infty).$
-From the continuous embeddings of $\dot{H}^{2,2}$ into $\dot{H}^{1,6}$ and $\dot{H}^{2,2}$ into $L^{q}$ for all $q\in(2,\infty)$, it follows that
+The continuous embeddings of $\dot{H}^{2,2}$ into $\dot{H}^{1,6}$ and $\dot{H}^{2,2}$ into $L^{q}$ for all $q\in(2,\infty)$ yield
 $y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}),\quad\forall p,q\in(2,\infty),$
 $y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{1,6}),\quad\forall p\in(2,\infty).$
-Hence, Eqs. 4.16 and 4.17 can be improved to
+Hence, Eqs. 4.10 and 4.11 can be refined to
 $S_{0}(y_{n}-y_{n}^{3})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,q}),\quad\forall p,q\in(2,\infty),$ (4.12)
-$S_{1}F(y_{n})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,6}),\quad\forall p\in(2,\infty).$ (4.20)
-Combining Eqs. 4.14, 4.18, 4.19 and 4.20 yields
+$S_{1}F(y_{n})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,6}),\quad\forall p\in(2,\infty).$
+Combining these refined regularity results with (4.8), we arrive at
 $y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,6}),\quad\forall p\in(2,\infty),$
-thereby implying, through the continuous embedding of $\dot{H}^{2,6}$ into $\dot{H}^{1,q}$ for all $q\in(2,\infty)$, that
+which, through the continuous embedding of $\dot{H}^{2,6}$ into $\dot{H}^{1,q}$ for all $q\in(2,\infty)$, implies that
 $y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{1,q}),\quad\forall p,q\in(2,\infty).$
-This regularity result allows us to enhance Eq. 4.17 to
+This regularity result allows us to further improve (4.11) to
 $S_{1}F(y_{n})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,q}),\quad\forall p,q\in(2,\infty),$
-which, combined with Eqs. 4.14, 4.18 and 4.19, implies
+which, in conjunction with (4.8) and (4.12), yields
 $y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,q}),\quad\forall p,q\in(2,\infty).$ (4.13)
@@
 we use Itô’s formula in UMD Banach spaces (see [7, Theorem 2.4]) to conclude that, for any $t\in(0,T]$,
-$\mathbb{E}\Big{(}\lVert{y_{n}(t)}\rVert_{L^{q}}^{p}-\lVert{v_{n}}\rVert_{L^{q}}^{p}\Big{)}$
+$\mathbb{E}\Big{[}\lVert{y_{n}(t)}\rVert_{L^{q}}^{p}-\lVert{v_{n}}\rVert_{L^{q}}^{p}\Big{]}$
 […]
-$\quad{}+\frac{p(q-1)}{2}\,\mathbb{E}\int_{0}^{t}\lVert{y_{n}(s)}\rVert_{L^{q}}^{p-q}\,\int_{\mathcal{O}}\lvert{y_{n}(s)}\rvert^{q-2}\sum_{k=0}^{\infty}\lambda_{k}\lvert{f_{k}(y_{n}(s))}\rvert^{2}\,\mathrm{d}\mu\,\mathrm{d}s$
-$\quad{}+\frac{p(p-q)}{2}\,\mathbb{E}\int_{0}^{t}\lVert{y_{n}(s)}\rVert_{L^{q}}^{p-2q}\sum_{k=0}^{\infty}\lambda_{k}\Big{[}\int_{\mathcal{O}}\lvert{y_{n}(s)}\rvert^{q-2}y_{n}(s)f_{k}(y_{n}(s))\,\mathrm{d}\mu\Big{]}^{2}\,\mathrm{d}s$
+$\quad{}+\frac{p(q-1)}{2}\,\mathbb{E}\Biggl{[}\int_{0}^{t}\lVert{y_{n}(s)}\rVert_{L^{q}}^{p-q}\,\int_{\mathcal{O}}\lvert{y_{n}(s)}\rvert^{q-2}\sum_{k=0}^{\infty}\lambda_{k}\lvert{f_{k}(y_{n}(s))}\rvert^{2}\,\mathrm{d}\mu\,\mathrm{d}s\Biggr{]}$
+$\quad{}+\frac{p(p-q)}{2}\,\mathbb{E}\Biggl{\{}\int_{0}^{t}\lVert{y_{n}(s)}\rVert_{L^{q}}^{p-2q}\sum_{k=0}^{\infty}\lambda_{k}\bigg{[}\int_{\mathcal{O}}\lvert{y_{n}(s)}\rvert^{q-2}y_{n}(s)f_{k}(y_{n}(s))\,\mathrm{d}\mu\bigg{]}^{2}\,\mathrm{d}s\Biggr{\}}$
 […]
@@
 we infer the bound for $I_{1}$:
-$I_{1}\leqslant p\,\mathbb{E}\int_{0}^{t}\lVert{y_{n}(s)}\rVert_{L^{q}}^{p}\,\mathrm{d}s.$
-Regarding $I_{2}+I_{3}$, by using Hölder’s inequality in conjunction with Young’s inequality, and employing the summability condition Eq.
-2.5 alongside the growth condition Eq.
-2.2, a straightforward calculation leads to the following bound for $I_{2}+I_{3}$:
-$I_{2}+I_{3}\leqslant c\,\mathbb{E}\int_{0}^{t}1+\lVert{y_{n}(s)}\rVert_{L^{q}}^{p}\,\mathrm{d}s.$
+$I_{1}\leqslant p\,\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{y_{n}(s)}\rVert_{L^{q}}^{p}\,\mathrm{d}s\bigg{]}.$
+For $I_{2}+I_{3}$, using Hölder’s inequality in conjunction with Young’s inequality, and employing the summability condition Eq.
+2.5 along with the growth condition Eq.
+2.2, we derive the following bound:
+$I_{2}+I_{3}\leqslant c\,\mathbb{E}\bigg{[}\int_{0}^{t}1+\lVert{y_{n}(s)}\rVert_{L^{q}}^{p}\,\mathrm{d}s\bigg{]}.$
 Combining the above bounds for $I_{1}$ and $I_{2}+I_{3}$, we arrive at
-$\mathbb{E}\lVert{y_{n}(t)}\rVert_{L^{q}}^{p}\leqslant\lVert{v_{n}}\rVert_{L^{q}}^{p}+c\,\mathbb{E}\int_{0}^{t}1+\lVert{y_{n}(s)}\rVert_{L^{q}}^{p}\,\mathrm{d}s,\quad\forall t\in[0,T].$
+$\mathbb{E}\lVert{y_{n}(t)}\rVert_{L^{q}}^{p}\leqslant\lVert{v_{n}}\rVert_{L^{q}}^{p}+c\,\mathbb{E}\bigg{[}\int_{0}^{t}1+\lVert{y_{n}(s)}\rVert_{L^{q}}^{p}\,\mathrm{d}s\bigg{]},\quad\forall t\in[0,T].$
 Applying Gronwall’s lemma to the above inequality yields
-$\lVert{y_{n}(t)}\rVert_{L^{p}(\Omega;L^{q})}\leqslant c\big{(}1+\lVert{v_{n}}\rVert_{L^{q}}\big{)},\quad\forall t\in[0,T],$
+$\lVert{y_{n}}\rVert_{C([0,T];L^{p}(\Omega;L^{q}))}\leqslant c\big{(}1+\lVert{v_{n}}\rVert_{L^{q}}\big{)},$
 which implies the desired inequality Eq. 4.7. This completes the proof. ∎
@@
 Step 1.
-Let us prove that Eq.
-1.1 admits a unique mild solution $y$ and Eq.
-3.1 holds.
-Let $p,q\in(2,\infty)$ be arbitrary but fixed.
-For each $n\in\mathbb{N}$, let $y_{n}$ be the mild solution of Eq.
-4.12 with $v_{n}$ defined by
+We proceed to show that equation Eq.
+1.1 has a unique mild solution $y$, and that Eq.
+3.1 is satisfied.
+Let $p,q\in(2,\infty)$ be fixed but arbitrary.
+For each $n\in\mathbb{N}$, denote by $y_{n}$ the mild solution of equation Eq.
+4.6, where $v_{n}$ is defined as
 $v_{n}:=\mathop{\mathrm{argmin}}_{w\in\mathrm{span}\{\phi_{0},\ldots,\phi_{n}\}}\lVert{v-w}\rVert_{L^{3q}}.$
-For any $m,n\in\mathbb{N}$, put $e_{m,n}:=y_{m}-y_{n}$. By definition, we infer that, for any $t\in[0,T]$,
-$\mathrm{d}e_{m,n}(t)=(\Delta e_{m,n}+e_{m,n}+y_{n}^{3}-y_{m}^{3})(t)\,\mathrm{d}t+\big{(}F(y_{m}(t))-F(y_{n}(t))\big{)}\,\mathrm{d}W_{H}(t).$
+For any $m,n\in\mathbb{N}$, put $e_{m,n}:=y_{m}-y_{n}$. By definition, we infer that
+$\mathrm{d}e_{m,n}(t)=(\Delta e_{m,n}+e_{m,n}+y_{n}^{3}-y_{m}^{3})(t)\,\mathrm{d}t+\big{(}F(y_{m}(t))-F(y_{n}(t))\big{)}\,\mathrm{d}W_{H}(t),\quad t\in[0,T].$
 Observing that the integral over $\mathcal{O}$ of $\lvert{e_{m,n}(t)}\rvert^{q-2}e_{m,n}(t)(y_{n}^{3}-y_{m}^{3})(t)$ is non-positive for all $t\in[0,T]$, and following the proof of Eq.
-4.13, we can deduce that
-$\lVert{e_{m,n}}\rVert_{L_{\mathbb{F}}^{3p}(\Omega\times(0,T);L^{3q})}\leqslant c\lVert{v_{m}-v_{n}}\rVert_{L^{3q}}.$
-On the other hand, from the fact that $\dot{H}^{2,2}$ is dense in $L^{3q}$ and that $\{\phi_{n}\}_{n\in\mathbb{N}}$ is a complete orthogonal system in $\dot{H}^{2,2}$, we conclude that $\{v_{n}\}_{n\in\mathbb{N}}$ is a Cauchy sequence in $L^{3q}$, converging to $v$.
-Consequently, there exists a unique $y\in L_{\mathbb{F}}^{3p}(\Omega\times(0,T);L^{3q})$ such that
+4.7, we can deduce that
+$\lVert{e_{m,n}}\rVert_{L_{\mathbb{F}}^{3p}(\Omega\times(0,T);L^{3q})}\leqslant c\lVert{v_{m}-v_{n}}\rVert_{L^{3q}}.$ (4.14)
+On the other hand, from the fact that $\dot{H}^{2,2}$ is dense in $L^{3q}$ and that $(\phi_{n})_{n\in\mathbb{N}}$ is a complete orthogonal system in $\dot{H}^{2,2}$, we conclude that $(v_{n})_{n\in\mathbb{N}}$ is a Cauchy sequence in $L^{3q}$, converging to $v$.
+Consequently, there exists a unique process $y\in L_{\mathbb{F}}^{3p}(\Omega\times(0,T);L^{3q})$ such that
 $\lim_{n\to\infty}y_{n}=y\quad\text{ in }L_{\mathbb{F}}^{3p}(\Omega\times(0,T);L^{3q}).$ (4.15)
-For any $n\in\mathbb{N}$, application of Hölder’s inequality yields
+For any $n\in\mathbb{N}$, applying Hölder’s inequality yields
 $\lVert{y_{n}^{3}-y^{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}$
@@
 $\lim_{n\to\infty}y_{n}^{3}=y^{3}\quad\text{ in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}),$
-consequently leading to
+which consequently leads to
 $\lim_{n\to\infty}y_{n}-y_{n}^{3}=y-y^{3}\quad\text{ in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}).$
-Using Lemma 4.3, this directly implies the following convergence for $S_{0}(y_{n}-y_{n}^{3})$:
+Using Lemma 4.3, this directly implies the following convergence of $S_{0}(y_{n}-y_{n}^{3})$:
 $\lim_{n\to\infty}S_{0}(y_{n}-y_{n}^{3})=S_{0}(y-y^{3})\quad\text{ in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}).$
-Furthermore, invoking Eqs. 4.10 and 4.22 yields
+Furthermore, invoking Eqs. 4.4 and 4.15 yields
 $\lim_{n\to\infty}F(y_{n})=F(y)\quad\text{ in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,L^{q})),$
@@
 where $G$ is defined by Eq.
-2.11, and $G_{n}(t):=e^{t\Delta}v_{n}$ for each $t\in[0,T]$.
-Finally, combining the derived convergences of $y_{n}$, $G_{n}$, $S_{0}(y_{n}-y_{n}^{3})$, and $S_{1}F(y_{n})$, and recalling that $y_{n}=G_{n}+S_{0}(y_{n}-y_{n}^{3})+S_{1}F(y_{n})$, by passing to the limit $n\to\infty$ we obtain
+2.10, and $G_{n}(t):=e^{t\Delta}v_{n}$ for each $t\in[0,T]$.
+Finally, combining the derived convergences of $y_{n}$, $G_{n}$, $S_{0}(y_{n}-y_{n}^{3})$, and $S_{1}F(y_{n})$, and recalling that $y_{n}=G_{n}+S_{0}(y_{n}-y_{n}^{3})+S_{1}F(y_{n})$, we obtain by passing to the limit $n\to\infty$ that
 $y=G+S_{0}(y-y^{3})+S_{1}F(y)\quad\text{in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}).$
-This confirms that $y$ is a mild solution of Eq. 1.1, and it is evident that Eq. 3.1 holds. The uniqueness of the mild solution can be proved similarly as in [17, Theorem 1.1].
+This confirms that $y$ is a mild solution of Eq.
+1.1, and it is evident that Eq.
+3.1 holds.
+To prove the uniqueness of the mild solution, assume $\widetilde{y}$ is another mild solution.
+Consider the difference $e=y-\widetilde{y}$.
+Then $e$ satisfies the equation
+$\begin{cases}\mathrm{d}e(t)=(\Delta e+e+\widetilde{y}^{3}-y^{3})(t)\,\mathrm{d}t+\big{(}F(y(t))-F(\widetilde{y}(t))\big{)}\,\mathrm{d}W_{H}(t),&t\in[0,T],\\ e(0)=0.\end{cases}$
+By an argument similar to that used in Eq. 4.14, we find that
+$\lVert{e}\rVert_{L^{2}(\Omega\times(0,T);L^{2})}\leqslant 0,$
+which implies that $e=0$ almost surely, confirming the uniqueness of the mild solution to Eq. 1.1.
 Step 2.
@@
 Consequently, by the mild solution representation Eq.
-2.10, we arrive at
+2.9, we arrive at
 $y\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2/p-\epsilon,q}),\quad\forall\epsilon>0,$
@@
 Finally, combining Eqs.
-2.10, 4.23 and 4.24 establishes the regularity assertions in Eqs.
+2.9, 4.16 and 4.17 establishes the regularity assertions in Eqs.
 3.4 and 3.5.
@@
 4.3 Proof of Theorems 3.2 and 3.3
-We first summarize some standard results as follows.
+We begin by summarizing several well-established results. As known from [8], it holds that
+$\|P_{h}\|_{\mathcal{L}(L^{q})}\text{ is uniformly bounded with respect to }h\text{ for all }q\in(1,\infty).$ (4.18)
+Consequently, the operator $P_{h}$ possesses the following approximation property:
+$\|I-P_{h}\|_{\mathcal{L}(\dot{H}^{\alpha,q},L^{q})}\leqslant ch^{\alpha}\quad\text{for all }\alpha\in[0,2]\text{ and }q\in(1,\infty).$ (4.19)
+Inequality (2.5) from [1], combined with the observation that
+$\lVert{\Delta_{h}^{-1}P_{h}\Delta}\rVert_{\mathcal{L}(\dot{H}^{1,2})}\leqslant 1,$
+yields, through interpolation, that for any $q\in[2,\infty)$, the operator norm
+$\lVert{\Delta_{h}^{-1}P_{h}\Delta}\rVert_{\mathcal{L}(\dot{H}^{1,q})}$
+is uniformly bounded with respect to $h$.
+Utilizing this fundamental stability estimate, the approximation property Eq.
+4.19, the interpolation theory as stated in [22, Theorem 2.6], and the fact that $\Delta^{-1}$ is a bounded linear operator mapping from $\dot{H}^{-s,q}$ to $\dot{H}^{2-s,q}$ for all $s\in[0,2]$, standard arguments then lead to the conclusion that
+$\|P_{h}\Delta^{-1}-\Delta_{h}^{-1}P_{h}\|_{\mathcal{L}(\dot{H}^{\beta-2,q},\dot{H}^{\alpha,q})}\leqslant ch^{\beta-\alpha}\quad\text{for all }0\leqslant\alpha\leqslant 1\leqslant\beta\leqslant 2.$ (4.20)
+Let $e^{t\Delta_{h}}$, for $t\in[0,\infty)$, denote the analytic semigroup generated by the discrete Laplace operator $\Delta_{h}$.
+Using the established estimates for the resolvents of $\Delta_{h}$ from [1] and applying (2.128) from [38], we obtain the well-known growth estimate: for any $\alpha\in[0,1]$, $q\in(1,\infty)$, and $t\in(0,T]$,
+$\lVert{(-\Delta_{h})^{\alpha}e^{t\Delta_{h}}}\rVert_{\mathcal{L}(L^{q})}\leqslant ct^{-\alpha}.$ (4.21)
+Furthermore, utilizing techniques analogous to those found in Theorem 6.10 and Lemma 20.7 of [35], one can readily derive the following well-known inequality: for any $\alpha\in[0,2]$, $q\in(1,\infty)$, and $t\in[0,T]$,
+$\lVert{e^{t\Delta_{h}}P_{h}-P_{h}e^{t\Delta}}\rVert_{\mathcal{L}(\dot{H}^{\alpha,q},L^{q})}+\lVert{(I-P_{h})e^{t\Delta}}\rVert_{\mathcal{L}(\dot{H}^{\alpha,q},L^{q})}\leqslant c\min\Big{\{}h^{\alpha},\frac{h^{2}}{t^{1-\alpha/2}}\Big{\}}.$ (4.22)
+For the special case where $q=2$, the reader is referred to Theorem 3.2 of [35] for a detailed treatment of this result.
+Let $q\in(1,\infty)$. For any $\alpha\in\mathbb{R}$, denote by $\dot{H}_{h}^{\alpha,q}$ the space $X_{h}$ equipped with the norm
+$\|v_{h}\|_{\dot{H}_{h}^{\alpha,q}}:=\|(-\Delta_{h})^{\alpha/2}v_{h}\|_{L^{q}}\quad\text{for all }v_{h}\in X_{h}.$
+According to Theorem 3.1 (ii) in [17], the boundedness of the purely imaginary powers of $-\Delta_{h}$ in $\mathcal{L}(\dot{H}_{h}^{0,q})$ is independent of the spatial mesh size $h$.
+Consequently, by applying Theorem 4.17 from [22], we deduce that for any $\alpha\in(0,2)$,
+$c_{0}\|v_{h}\|_{\dot{H}_{h}^{\alpha,q}}\leqslant\|v_{h}\|_{[\dot{H}_{h}^{0,q},\dot{H}_{h}^{2,q}]_{\alpha/2}}\leqslant c_{1}\|v_{h}\|_{\dot{H}_{h}^{\alpha,q}}\quad\text{for all }v_{h}\in X_{h},$ (4.23)
+where $c_{0}$ and $c_{1}$ are positive constants independent of $h$.
+Here, $[\dot{H}_{h}^{0,q},\dot{H}_{h}^{2,q}]_{\alpha/2}$ denotes the complex interpolation space between $\dot{H}_{h}^{0,q}$ and $\dot{H}_{h}^{2,q}$ (see [22, Chapter 2]).
+This result is important as it enables the application of interpolation theory to the space $\dot{H}_{h}^{s,q}$ for all $s\in\mathbb{R}$.
+For example, given the standard inverse estimate
+$\|v_{h}\|_{\dot{H}_{h}^{2,q}}\leqslant ch^{-2}\|v_{h}\|_{L^{q}}\quad\text{for all }v_{h}\in X_{h},$
+we can invoke [22, Theorem 2.6] to derive the following inverse estimate: for any $0\leqslant\alpha<\beta<\infty$,
+$\|v_{h}\|_{\dot{H}_{h}^{\beta,q}}\leqslant ch^{\alpha-\beta}\|v_{h}\|_{\dot{H}_{h}^{\alpha,q}}\quad\text{for all }v_{h}\in X_{h}.$
+Furthermore, for any $\alpha\in[0,1]$, there exist positive constants $c_{0}$ and $c_{1}$, independent of $h$, such that
+$c_{0}\|v_{h}\|_{\dot{H}^{\alpha,q}}\leqslant\|v_{h}\|_{\dot{H}_{h}^{\alpha,q}}\leqslant c_{1}\|v_{h}\|_{\dot{H}^{\alpha,q}}\quad\text{for all }v_{h}\in X_{h}.$
+Henceforth, the aforementioned theoretical results will be utilized without further explicit reference.
+We now introduce a discrete deterministic maximal $L^{p}$-regularity estimate and its stochastic counterpart.
+For a function $g_{h}\in L^{1}(0,T;\dot{H}_{h}^{0,2})$, the discrete deterministic convolution operator $S_{0,h}$ is defined as
+$(S_{0,h}g_{h})(t):=\int_{0}^{t}e^{(t-s)\Delta_{h}}g_{h}(s)\,\mathrm{d}s,\quad\text{for }t\in[0,T].$ (4.24)
+For a process $g_{h}\in L_{\mathbb{F}}^{2}(\Omega\times(0,T);\gamma(H,\dot{H}_{h}^{0,2}))$, the discrete stochastic convolution operator $S_{1,h}$ is given by
+$(S_{1,h}g_{h})(t):=\int_{0}^{t}e^{(t-s)\Delta_{h}}g_{h}(s)\,\mathrm{d}W_{H}(s),\quad\text{for }t\in[0,T].$ (4.25)
+The proofs of the following two lemmas can be found in [16] for the first and in [17] for the second.
 Lemma 4.7.
-Let $p,q\in(1,\infty)$. For any $g_{h}\in L^{p}(0,T;X_{h})$, the following inequality holds for all $\epsilon>0$:
-$\max_{t\in[0,T]}\|S_{0,h}g_{h}(t)\|_{\dot{H}_{h}^{2-\frac{2}{p}-\epsilon,q}}\leqslant c\|g_{h}\|_{L^{p}(0,T;L^{q})}.$
+Suppose $p,q\in(1,\infty)$. For any $g_{h}\in L^{p}(0,T;X_{h})$, the following discrete deterministic maximal $L^{p}$-regularity estimate is valid:
+$\Big{\lVert}{\frac{\mathrm{d}}{\mathrm{d}t}S_{0,h}g_{h}}\Big{\rVert}_{L^{p}(0,T;\dot{H}_{h}^{0,q})}+\lVert{S_{0,h}g_{h}}\rVert_{L^{p}(0,T;\dot{H}_{h}^{2,q})}\leqslant c\lVert{g_{h}}\rVert_{L^{p}(0,T;\dot{H}_{h}^{0,q})}.$
+Lemma 4.8.
+Let $p\in(2,\infty)$ and $q\in[2,\infty)$.
+For any $g_{h}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,\dot{H}_{h}^{0,q}))$, the following discrete stochastic maximal $L^{p}$-regularity estimate holds:
+$\lVert{S_{1,h}g_{h}}\rVert_{L^{p}(\Omega;C([0,T];(\dot{H}_{h}^{0,q},\dot{H}_{h}^{2,q})_{1/2-1/p,p}))}+\lVert{S_{1,h}g_{h}}\rVert_{L^{p}(\Omega\times(0,T);\dot{H}_{h}^{1,q})}$ (4.26)
+[…]
+Additionally, for the discrete deterministic convolution operator $S_{0,h}$, we establish the following property.
+Lemma 4.9.
+Let $p,q\in(1,\infty)$. For any $g_{h}\in L^{p}(0,T;\dot{H}_{h}^{0,q})$, the following inequality holds for all $\epsilon>0$:
+$\|S_{0,h}g_{h}(t)\|_{C([0,T];\dot{H}_{h}^{2-2/p-\epsilon,q})}\leqslant c\|g_{h}\|_{L^{p}(0,T;L^{q})}.$
 Proof.
-For any $t\in[0,T]$, recognizing that $S_{0,h}g_{h}(t)=\int_{0}^{t}e^{(t-s)\Delta_{h}}g_{h}(s)\,\mathrm{d}s$, we proceed as follows:
-$\|S_{0,h}g_{h}(t)\|_{\dot{H}_{h}^{2-\frac{2}{p}-\epsilon,q}}$ $=\left\|\int_{0}^{t}e^{(t-s)\Delta_{h}}g_{h}(s)\,\mathrm{d}s\right\|_{\dot{H}_{h}^{2-2/p-\epsilon,q}}$
-$\leqslant\int_{0}^{t}\left\|e^{(t-s)\Delta_{h}}g_{h}(s)\right\|_{\dot{H}_{h}^{2-2/p-\epsilon,q}}\,\mathrm{d}s$
+For any $t\in[0,T]$, from the identity $S_{0,h}g_{h}(t)=\int_{0}^{t}e^{(t-s)\Delta_{h}}g_{h}(s)\,\mathrm{d}s$, we deduce that
+$\|S_{0,h}g_{h}(t)\|_{\dot{H}_{h}^{2-2/p-\epsilon,q}}$ $\leqslant\int_{0}^{t}\left\|e^{(t-s)\Delta_{h}}g_{h}(s)\right\|_{\dot{H}_{h}^{2-2/p-\epsilon,q}}\,\mathrm{d}s$
 $\stackrel{{\scriptstyle\text{(i)}}}{{\leqslant}}c\int_{0}^{t}(t-s)^{-1+\frac{1}{p}+\frac{\epsilon}{2}}\|g_{h}(s)\|_{L^{q}}\,\mathrm{d}s$
 $\stackrel{{\scriptstyle\text{(ii)}}}{{\leqslant}}c\Big{(}\int_{0}^{t}(t-s)^{(-1+\frac{1}{p}+\frac{\epsilon}{2})\frac{p}{p-1}}\,\mathrm{d}s\Big{)}^{1-\frac{1}{p}}\|g_{h}\|_{L^{p}(0,T;L^{q})},$
-where step (i) invokes Eq. 4.5 with $\alpha=2-2/p-\epsilon$, and step (ii) uses Hölder’s inequality. The uniform boundedness of the integral
+where step (i) uses the estimate Eq. 4.21 with $\alpha=2-2/p-\epsilon$, and step (ii) applies Hölder’s inequality. Since the integral
 $\int_{0}^{t}(t-s)^{\left(-1+\frac{1}{p}+\frac{\epsilon}{2}\right)\frac{p}{p-1}}\,\mathrm{d}s$
-with respect to $t\in[0,T]$ then establishes the lemma. ∎
+is uniformly bounded with respect to $t\in[0,T]$, the desired inequality follows, thereby completing the proof of the lemma. ∎
+We proceed to present some auxiliary estimates.
 Lemma 4.10.
-Assume that $v\in L^{\infty}$.
-Let $y$ and $y_{h}$ be the mild solutions of Eqs.
-1.1 and 2.7, respectively, and let $G$ and $G_{h}$ be defined by Eqs.
-2.11 and 2.12, respectively.
-For simplicity, let $z:=y-G$ and $\xi_{h}:=S_{0,h}(\Delta_{h}P_{h}-P_{h}\Delta)z$.
-Then, for any $p\in(2,\infty)$, $q\in[2,\infty)$, and $\varepsilon>0$, the following inequalities hold:
+Suppose that $v\in L^{\infty}$.
+Consider the mild solution $y$ to equation (1.1), the strong solution $y_{h}$ to equation (2.11), and the function $G$ defined by equation (2.10).
+Additionally, let $G_{h}(t):=e^{t\Delta_{h}}P_{h}v$ for $t\in[0,T]$.
+For notational simplicity, let $z:=y-G$ and $\xi_{h}:=S_{0,h}(\Delta_{h}P_{h}-P_{h}\Delta)z$.
+For any $p\in(2,\infty)$, $q\in[2,\infty)$, and $\varepsilon>0$, the following inequalities hold:
 $\lVert{G-G_{h}}\rVert_{L^{p}(0,T;L^{q})}+\lVert{P_{h}G-G_{h}}\rVert_{L^{p}(0,T;L^{q})}\leqslant ch^{\frac{2}{p}},$ (4.27)
@@
 $\lVert{\xi_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{1-\varepsilon},$ (4.30)
-$\mathbb{E}\int_{0}^{T}\lVert{(y-P_{h}y)(t)}\rVert_{L^{4}}^{p}\lVert{y(t)}\rVert_{L^{8}}^{2p}\,\mathrm{d}t\leqslant ch^{2-\varepsilon}.$ (4.29)
+$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{(y-P_{h}y)(t)}\rVert_{L^{4}}^{p}\lVert{y(t)}\rVert_{L^{8}}^{2p}\,\mathrm{d}t\bigg{]}\leqslant ch^{2-\varepsilon}.$ (4.31)
 Furthermore, for any $\beta\in\Big{(}1,\frac{3p}{3p-4}\Big{)}$, the following inequalities hold:
-$\mathbb{E}\int_{0}^{T}\lVert{(P_{h}G-G_{h})(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\big{(}\lVert{y(t)}\rVert_{L^{2\beta}}^{p}+\lVert{y_{h}(t)}\rVert_{L^{2\beta}}^{p}\big{)}\,\mathrm{d}t\leqslant ch^{2-\frac{3}{2}\big{(}1-\frac{1}{\beta}\big{)}p},$ (4.30)
-$\mathbb{E}\int_{0}^{T}\lVert{\xi_{h}(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\big{(}\lVert{y(t)}\rVert_{L^{2\beta}}^{p}+\lVert{y_{h}(t)}\rVert_{L^{2\beta}}^{p}\big{)}\,\mathrm{d}t\leqslant ch^{p+2-\frac{3}{2}\big{(}1-\frac{1}{\beta}\big{)}p-\varepsilon}.$ (4.31)
+$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{(P_{h}G-G_{h})(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\big{(}\lVert{y(t)}\rVert_{L^{2\beta}}^{p}+\lVert{y_{h}(t)}\rVert_{L^{2\beta}}^{p}\big{)}\,\mathrm{d}t\bigg{]}\leqslant ch^{2-\frac{3}{2}\big{(}1-\frac{1}{\beta}\big{)}p},$ (4.32)
+$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{\xi_{h}(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\big{(}\lVert{y(t)}\rVert_{L^{2\beta}}^{p}+\lVert{y_{h}(t)}\rVert_{L^{2\beta}}^{p}\big{)}\,\mathrm{d}t\bigg{]}\leqslant ch^{p+2-\frac{3}{2}\big{(}1-\frac{1}{\beta}\big{)}p-\varepsilon}.$ (4.33)
+Here, $\beta^{\prime}$ denotes the conjugate exponent of $\beta$.
 Proof.
@@
 4.19 and the regularity assertions in Theorem 3.1.
-Now let us prove Eq. 4.30. By invoking Eq. 4.2 and replicating the reasoning behind [18, Equation (2.2)], we establish that
-$\mathbb{E}\int_{0}^{T}\lVert{y_{h}(t)}\rVert_{L^{2}}^{\frac{4\beta}{3\beta-3}-2}\lVert{y_{h}(t)}\rVert_{\dot{H}^{1,2}}^{2}\,\mathrm{d}t$
+Now let us prove Eq. 4.32. By invoking Eq. 4.18 and replicating the reasoning behind [19, Equation (2.2)], we establish that
+$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{y_{h}(t)}\rVert_{L^{2}}^{\frac{4\beta}{3\beta-3}-2}\lVert{y_{h}(t)}\rVert_{\dot{H}^{1,2}}^{2}\,\mathrm{d}t\bigg{]}$
 is uniformly bounded with respect to $h$. Using the continuous embedding of $\dot{H}^{1,2}$ into $L^{6}$, this uniform bound extends to
-$\mathbb{E}\int_{0}^{T}\lVert{y_{h}(t)}\rVert_{L^{2}}^{\frac{4\beta}{3\beta-3}-2}\lVert{y_{h}(t)}\rVert_{L^{6}}^{2}\,\mathrm{d}t.$
+$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{y_{h}(t)}\rVert_{L^{2}}^{\frac{4\beta}{3\beta-3}-2}\lVert{y_{h}(t)}\rVert_{L^{6}}^{2}\,\mathrm{d}t\bigg{]}.$
 Employing Hölder’s inequality, we deduce for all $t\in[0,T]$ that
-$\lVert{y_{h}(t)}\rVert_{L^{2\beta}}^{\frac{4\beta}{3\beta-3}}\leq\lVert{y_{h}(t)}\rVert_{L^{2}}^{\frac{4\beta}{3\beta-3}-2}\lVert{y_{h}(t)}\rVert_{L^{6}}^{2},$
+$\lVert{y_{h}(t)}\rVert_{L^{2\beta}}^{\frac{4\beta}{3\beta-3}}\leqslant\lVert{y_{h}(t)}\rVert_{L^{2}}^{\frac{4\beta}{3\beta-3}-2}\lVert{y_{h}(t)}\rVert_{L^{6}}^{2},$
 which in turn confirms the uniform bound on
@@
 Subsequently, applying Hölder’s inequality once more yields
-$\mathbb{E}\int_{0}^{T}\lVert{(P_{h}G-G_{h})(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\lVert{y_{h}(t)}\rVert_{2\beta}^{p}\,\mathrm{d}t$
+$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{(P_{h}G-G_{h})(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\lVert{y_{h}(t)}\rVert_{2\beta}^{p}\,\mathrm{d}t\bigg{]}$
 $\leqslant\lVert{P_{h}G-G_{h}}\rVert_{L^{\frac{4p\beta}{4\beta-(3\beta-3)p}}(\Omega\times(0,T);L^{2\beta^{\prime}})}^{p}\lVert{y_{h}}\rVert_{L^{\frac{4\beta}{3\beta-3}}(\Omega\times(0,T);L^{2\beta})}^{p}$
 $\leqslant c\lVert{P_{h}G-G_{h}}\rVert_{L^{\frac{4p\beta}{4\beta-(3\beta-3)p}}(\Omega\times(0,T);L^{2\beta^{\prime}})}^{p}.$
-Invoking the inequality Eq. 4.25 with $p$ replaced by $4p\beta(4\beta-(3\beta-3)p)$, we find
-$\mathbb{E}\int_{0}^{T}\lVert{(P_{h}G-G_{h})(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\lVert{y_{h}(t)}\rVert_{2\beta}^{p}\,\mathrm{d}t\leqslant ch^{2-\frac{3}{2}(1-\frac{1}{\beta})p}.$
-Given that Theorem 3.1 ensures $y$ resides within $L^{\frac{4\beta}{3\beta-3}}(\Omega\times(0,T);L^{2\beta})$, we can mirror the preceding argument to deduce that
-$\mathbb{E}\int_{0}^{T}\lVert{(P_{h}G-G_{h})(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\lVert{y(t)}\rVert_{2\beta}^{p}\,\mathrm{d}t\leqslant ch^{2-\frac{3}{2}(1-1/\beta)p}.$
-The inequality Eq. 4.30 is thus substantiated by combining the above derived inequalities.
-Finally, by invoking Eq. 4.27 and adhering to the same logical progression as demonstrated in the proof of Eq. 4.30, we deduce Eq. 4.31. This concludes the proof. ∎
+Invoking the inequality Eq. 4.27 with $p$ replaced by $4p\beta/(4\beta-(3\beta-3)p)$, we find
+$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{(P_{h}G-G_{h})(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\lVert{y_{h}(t)}\rVert_{2\beta}^{p}\,\mathrm{d}t\bigg{]}\leqslant ch^{2-\frac{3}{2}(1-\frac{1}{\beta})p}.$
+Given that Theorem 3.1 ensures $y$ belongs to $L^{\frac{4\beta}{3\beta-3}}(\Omega\times(0,T);L^{2\beta})$, we can replicate the previous argument to conclude that
+$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{(P_{h}G-G_{h})(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\lVert{y(t)}\rVert_{2\beta}^{p}\,\mathrm{d}t\bigg{]}\leqslant ch^{2-\frac{3}{2}(1-1/\beta)p}.$
+Hence, Eq. 4.32 follows by combining the aforementioned inequalities. To deduce Eq. 4.33, we rely on Eq. 4.29 and proceed analogously to the proof of Eq. 4.32. This completes the proof. ∎
 Proof of Theorem 3.2. Fix any $\alpha\in(1,2)$ and $p\in(2,\infty)$. Let
@@
 $z_{h}(t)$ $:=y_{h}(t)-G_{h}(t),\quad 0\leqslant t\leqslant T,$ (4.35)
-where $G$ and $G_{h}$ are defined by Eqs. 2.11 and 2.12, respectively. From Eq. 1.1 and the fact $G^{\prime}(t)=\Delta G(t)$ for $t\in[0,T]$, we deduce that
+where $G$ is defined by Eq. 2.10, and $G_{h}$ is given by
+$G_{h}(t):=e^{t\Delta_{h}}P_{h}v,\quad t\in[0,T].$ (4.36)
+From Eq. 1.1 and the fact $G^{\prime}(t)=\Delta G(t)$ for $t\in[0,T]$, we deduce that
 $\mathrm{d}z(t)=\big{(}\Delta z+z+G-y^{3}\big{)}(t)\,\mathrm{d}t+F(y(t))\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T.$
@@
 $\mathrm{d}e_{h}(t)$ $=\Big{(}\Delta_{h}e_{h}+\Delta_{h}P_{h}z-P_{h}\Delta z+e_{h}+G_{h}-P_{h}G+P_{h}\big{(}y^{3}-y_{h}^{3}\big{)}\Big{)}(t)\,\mathrm{d}t+{}$ (4.37)
-$\quad P_{h}\Big{(}F\big{(}y_{h}(t))-F\big{(}y(t)\big{)}\Big{)}\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T.$
+$\quad P_{h}\Big{(}F\big{(}y_{h}(t)\big{)}-F\big{(}y(t)\big{)}\Big{)}\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T.$
 Defining $\eta_{h}:=e_{h}-\xi_{h}$, where $\xi_{h}:=S_{0,h}(\Delta_{h}P_{h}z-P_{h}\Delta z)$, and considering the differential equation
@@
 By applying Itô’s formula and considering the initial condition $\eta_{h}(0)=0$ and the definition of $F$ given by Eq. 2.6, we deduce that for any $0\leqslant t\leqslant T$,
-$\mathbb{E}\lVert{\eta_{h}(t)}\rVert_{L^{2}}^{p}$ $=p\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot\Big{[}\Delta_{h}\eta_{h}(s)+\xi_{h}(s)+\eta_{h}(s)$
-$\qquad\qquad\qquad\qquad\qquad{}+(G_{h}-P_{h}G)(s)+\big{(}y^{3}-y_{h}^{3}\big{)}(s)\Big{]}\,\mathrm{d}\mu\,\mathrm{d}s$
-$\qquad{}+\frac{p}{2}\,\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\sum_{k=0}^{\infty}\lambda_{k}\lVert{P_{h}\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}}\rVert_{L^{2}}^{2}\,\mathrm{d}s$
-$\qquad{}+\frac{p(p-2)}{2}\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-4}\sum_{k=0}^{\infty}\lambda_{k}\Big{[}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}\,\mathrm{d}\mu\Big{]}^{2}\,\mathrm{d}s$
+$\mathbb{E}\lVert{\eta_{h}(t)}\rVert_{L^{2}}^{p}$ $=p\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot\Big{(}\Delta_{h}\eta_{h}(s)+\xi_{h}(s)+\eta_{h}(s)$
+$\qquad\qquad\qquad\qquad\qquad{}+(G_{h}-P_{h}G)(s)+\big{(}y^{3}-y_{h}^{3}\big{)}(s)\Big{)}\,\mathrm{d}\mu\,\mathrm{d}s\biggr{]}$
+$\qquad{}+\frac{p}{2}\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\sum_{k=0}^{\infty}\lambda_{k}\lVert{P_{h}\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}}\rVert_{L^{2}}^{2}\,\mathrm{d}s\biggr{]}$
+$\qquad{}+\frac{p(p-2)}{2}\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-4}\sum_{k=0}^{\infty}\lambda_{k}\Big{(}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}\,\mathrm{d}\mu\Big{)}^{2}\,\mathrm{d}s\biggr{]}$
 $=I_{1}+I_{2}+I_{3}+I_{4}+I_{5}+I_{6},$
 where
-$I_{1}$ $:=p\,\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot\Delta_{h}\eta_{h}(s)\,\mathrm{d}\mu\,\mathrm{d}s,$
-$I_{2}$ $:=p\,\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot(\xi_{h}+\eta_{h}+G_{h}-P_{h}G)(s)\,\mathrm{d}\mu\,\mathrm{d}s,$
-$I_{3}$ $:=p\,\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{[}y^{3}-(P_{h}y)^{3}\big{]}(s)\,\mathrm{d}\mu\,\mathrm{d}s,$
-$I_{4}$ $:=p\,\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{[}(P_{h}y)^{3}-y_{h}^{3}\big{]}(s)\,\mathrm{d}\mu\,\mathrm{d}s,$
-$I_{5}$ $:=\frac{p}{2}\,\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\sum_{k=0}^{\infty}\lambda_{k}\lVert{P_{h}\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}}\rVert_{L^{2}}^{2}\,\mathrm{d}s,$
-$I_{6}$ $:=\frac{p(p-2)}{2}\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-4}\sum_{k=0}^{\infty}\lambda_{k}\Big{[}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}\,\mathrm{d}\mu\Big{]}^{2}\,\mathrm{d}s.$
+$\allowdisplaybreaks I_{1}$ $:=p\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot\Delta_{h}\eta_{h}(s)\,\mathrm{d}\mu\,\mathrm{d}s\biggr{]},$
+$I_{2}$ $:=p\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot(\xi_{h}+\eta_{h}+G_{h}-P_{h}G)(s)\,\mathrm{d}\mu\,\mathrm{d}s\biggr{]},$
+$I_{3}$ $:=p\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{[}y^{3}-(P_{h}y)^{3}\big{]}(s)\,\mathrm{d}\mu\,\mathrm{d}s\biggr{]},$
+$I_{4}$ $:=p\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{[}(P_{h}y)^{3}-y_{h}^{3}\big{]}(s)\,\mathrm{d}\mu\,\mathrm{d}s\biggr{]},$
+$I_{5}$ $:=\frac{p}{2}\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\sum_{k=0}^{\infty}\lambda_{k}\lVert{P_{h}\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}}\rVert_{L^{2}}^{2}\,\mathrm{d}s\biggr{]},$
+$I_{6}$ $:=\frac{p(p-2)}{2}\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-4}\sum_{k=0}^{\infty}\lambda_{k}\Big{(}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}\,\mathrm{d}\mu\Big{)}^{2}\,\mathrm{d}s\biggr{]}.$
 Next, let us estimate $I_{1}$, $I_{2}$, $I_{3}$, and $I_{4}$. For the term $I_{1}$, by the definition of $\Delta_{h}$, we establish that
-$I_{1}=-p\mathbb{E}\int_{0}^{t}\|{\eta_{h}(s)}\|_{L^{2}}^{p-2}\|{\eta_{h}(s)}\|_{\dot{H}^{1,2}}^{2}\,\mathrm{d}s.$ (4.36)
-Regarding $I_{2}$, we can deduce that
-$I_{2}$ $\leqslant\,c\mathbb{E}\int_{0}^{t}\|{\eta_{h}(s)}\|_{L^{2}}^{p-2}\big{(}\|{\xi_{h}(s)}\|_{L^{2}}^{2}+\|{\eta_{h}(s)}\|_{L^{2}}^{2}+\|{(G_{h}-P_{h}G)(s)}\|_{L^{2}}^{2}\big{)}\,\mathrm{d}s$
-$\stackrel{{\scriptstyle\mathrm{(i)}}}{{\leqslant}}c\,\mathbb{E}\int_{0}^{t}\|{\xi_{h}(s)}\|_{L^{2}}^{p}+\|{\eta_{h}(s)}\|_{L^{2}}^{p}+\|{(G_{h}-P_{h}G)(s)}\|_{L^{2}}^{p}\,\mathrm{d}s$
-=$\stackrel{{\scriptstyle\mathrm{(ii)}}}{{=}}c\int_{0}^{t}\|{\xi_{h}(s)}\|_{L^{p}(\Omega;L^{2})}^{p}+\|{\eta_{h}(s)}\|_{L^{p}(\Omega;L^{2})}^{p}+\|{(G_{h}-P_{h}G)(s)}\|_{L^{2}}^{p}\,\mathrm{d}s$
-$\stackrel{{\scriptstyle\mathrm{(iii)}}}{{\leqslant}}ch^{2}+c\int_{0}^{t}\|{\eta_{h}(s)}\|_{L^{p}(\Omega;L^{2})}^{p}\,\mathrm{d}s,$ (4.37)
-where (i) follows from Young’s inequality, (ii) uses the determinism of $G$ and $G_{h}$, and (iii) is justified by the inequalities (4.25) and (4.27).
-For $I_{3}$, by employing Hölder’s inequality, Young’s inequality, and the stability property of $P_{h}$ as stated in Eq.
-4.2, a straightforward calculation yields
-$I_{3}$ $\leqslant c\,\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-1}\lVert{(y-P_{h}y)(s)}\rVert_{L^{4}}\Big{(}\lVert{y(s)}\rVert_{L^{8}}^{2}+\lVert{P_{h}y(s)}\rVert_{L^{8}}^{2}\Big{)}\,\mathrm{d}s$
-$\leqslant c\,\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-1}\lVert{(y-P_{h}y)(s)}\rVert_{L^{4}}\lVert{y(s)}\rVert_{L^{8}}^{2}\,\mathrm{d}s$
-$\leqslant c\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p}+\lVert{(y-P_{h}y)(s)}\rVert_{L^{4}}^{p}\lVert{y(s)}\rVert_{L^{8}}^{2p}\,\mathrm{d}s.$
+$I_{1}=-p\mathbb{E}\biggl{[}\int_{0}^{t}\|{\eta_{h}(s)}\|_{L^{2}}^{p-2}\|{\eta_{h}(s)}\|_{\dot{H}^{1,2}}^{2}\,\mathrm{d}s\biggr{]}.$ (4.39)
+Regarding $I_{2}$, by Hölder’s inequality, Young’s inequality, and the inequalities Eq. 4.27 and Eq. 4.29, we can deduce that
+$I_{2}\leqslant ch^{2}+c\int_{0}^{t}\|{\eta_{h}(s)}\|_{L^{p}(\Omega;L^{2})}^{p}\,\mathrm{d}s,$ (4.40)
+For $I_{3}$, by employing Hölder’s inequality, Young’s inequality, and the stability property of $P_{h}$ as stated in Eq. 4.18, a straightforward calculation yields
+$I_{3}$ $\leqslant c\,\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-1}\lVert{(y-P_{h}y)(s)}\rVert_{L^{4}}\Big{(}\lVert{y(s)}\rVert_{L^{8}}^{2}+\lVert{P_{h}y(s)}\rVert_{L^{8}}^{2}\Big{)}\,\mathrm{d}s\bigg{]}$
+$\leqslant c\,\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-1}\lVert{(y-P_{h}y)(s)}\rVert_{L^{4}}\lVert{y(s)}\rVert_{L^{8}}^{2}\,\mathrm{d}s\bigg{]}$
+$\leqslant c\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p}+\lVert{(y-P_{h}y)(s)}\rVert_{L^{4}}^{p}\lVert{y(s)}\rVert_{L^{8}}^{2p}\,\mathrm{d}s\bigg{]}.$
 By utilizing Eq. 4.31, we arrive at the following bound:
@@
 ×$\qquad\qquad\qquad{}\times\Big{(}\|y(s)\|^{p}_{L^{2\beta}}+\|y_{h}(s)\|^{p}_{L^{2\beta}}\Big{)}\bigg{]}\,\mathrm{d}s,\quad\forall\beta\in\big{(}1,\frac{3p}{3p-4}\big{)}.$
-Using (4.30), we further refine this bound to:
+Using (4.32) and (4.33), we further simplifies this bound to:
 $I_{4}\leqslant c\int_{0}^{t}\|\eta_{h}(s)\|^{p}_{L^{p}(\Omega;L^{2})}\,\mathrm{d}s+ch^{2-\frac{3}{2}p\left(1-\frac{1}{\beta}\right)},\quad\forall\beta\in\left(1,\frac{3p}{3p-4}\right).$
@@
 Now, let us proceed to estimate $I_{5}+I_{6}$. Using Hölder’s inequality and the stability property of $P_{h}$ from Eq. 4.18, we obtain the following bound:
-$I_{5}+I_{6}\leqslant c\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\sum_{k=0}^{\infty}\lambda_{k}\lVert{f_{k}(y_{h}(s))-f_{k}(y(s))}\rVert_{L^{2}}^{2}\,\mathrm{d}s.$
+$I_{5}+I_{6}\leqslant c\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\sum_{k=0}^{\infty}\lambda_{k}\lVert{f_{k}(y_{h}(s))-f_{k}(y(s))}\rVert_{L^{2}}^{2}\,\mathrm{d}s\bigg{]}.$
 Invoking the Lipschitz continuity condition Eq. 2.3 and the summability condition Eq. 2.5, we further deduce that
-$I_{5}+I_{6}\leqslant c\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\lVert{(y_{h}-y)(s)}\rVert_{L^{2}}^{2}\,\mathrm{d}s.$
+$I_{5}+I_{6}\leqslant c\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\lVert{(y_{h}-y)(s)}\rVert_{L^{2}}^{2}\,\mathrm{d}s\bigg{]}.$
 Subsequently, by Young’s inequality, we obtain
-$I_{5}+I_{6}\leqslant c\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p}+\lVert{(y_{h}-y)(s)}\rVert_{L^{2}}^{p}\,\mathrm{d}s.$
-Recalling the definitions in Eqs. 4.32 and 4.33 and considering the relation $z_{h}-P_{h}z=\xi_{h}+\eta_{h}$, we infer that
+$I_{5}+I_{6}\leqslant c\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p}+\lVert{(y_{h}-y)(s)}\rVert_{L^{2}}^{p}\,\mathrm{d}s\bigg{]}.$
+Recalling the definitions in Eqs. 4.34 and 4.35 and considering the relation $z_{h}-P_{h}z=\xi_{h}+\eta_{h}$, we infer that
 $(y_{h}-y)(s)=\xi_{h}(s)+\eta_{h}(s)-(z-P_{h}z)(s)-(G-G_{h})(s),\quad\forall s\in[0,T],$ (4.43)
-which, in conjunction with Eq. 4.25, Eq. 4.26, and Eq. 4.27, yields
-$\mathbb{E}\int_{0}^{t}\lVert{(y_{h}-y)(s)}\rVert_{L^{2}}^{p}\,\mathrm{d}s\leqslant ch^{2}+\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{p}(\Omega;L^{2})}^{p}\,\mathrm{d}s.$
+which, in conjunction with Eqs. 4.27, 4.28 and 4.29, yields
+$\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{(y_{h}-y)(s)}\rVert_{L^{2}}^{p}\,\mathrm{d}s\bigg{]}\leqslant ch^{2}+\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{p}(\Omega;L^{2})}^{p}\,\mathrm{d}s.$
 Consequently, we obtain the following bound for $I_{5}+I_{6}$:
 $I_{5}+I_{6}\leqslant ch^{2}+c\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{p}(\Omega;L^{2})}^{p}\,\mathrm{d}s.$ (4.44)
-Finally, by integrating the estimates for $I_{1}$, $I_{2}$, $I_{3}$, $I_{4}$, and $I_{5}+I_{6}$, as given by Eqs. 4.36, 4.37, 4.38, 4.39 and 4.41, we obtain the following inequality:
-$\|\eta_{h}(t)\|_{L^{p}(\Omega;L^{2})}^{p}+\mathbb{E}\int_{0}^{t}\|\eta_{h}(s)\|_{L^{2}}^{p-2}\|\eta_{h}(s)\|_{\dot{H}_{h}^{1,2}}^{2}\,\mathrm{d}s$
+Finally, by integrating the estimates for $I_{1}$, $I_{2}$, $I_{3}$, $I_{4}$, and $I_{5}+I_{6}$, as given by Eqs. 4.39, 4.40, 4.41, 4.42 and 4.44, we obtain the following inequality:
+$\|\eta_{h}(t)\|_{L^{p}(\Omega;L^{2})}^{p}+\mathbb{E}\bigg{[}\int_{0}^{t}\|\eta_{h}(s)\|_{L^{2}}^{p-2}\|\eta_{h}(s)\|_{\dot{H}_{h}^{1,2}}^{2}\,\mathrm{d}s\bigg{]}$
 […]
@@
 Let $G$ be defined by Eq.
-2.11, and let $z$ be defined by Eq.
-4.32.
-Extending the reasoning from the proof of Theorem 3.1, we establish the following enhanced regularity properties for $y$ and $z$:
+2.10, and let $z$ be defined by Eq.
+4.34.
+Extending the reasoning from the proof of Theorem 3.1, we establish the following enhanced regularity properties for the processes $y$ and $z$:
 $y\in L^{p}(\Omega;C([0,T];\dot{H}^{2-\epsilon,q})),$ $\forall p\in(2,\infty),\,\forall q\in[2,\infty),\,\forall\epsilon>0;$ (4.45)
@@
 $z\in L^{p}(\Omega;C([0,T];\dot{H}^{2-\epsilon,q})),$ $\forall p\in(2,\infty),\,\forall q\in[2,\infty),\,\forall\epsilon>0.$ (4.47)
-Let $G_{h}$ be defined by Eq. 2.12 and let $z_{h}$ be defined by Eq. 4.33. We adopt the same notation introduced in the proof of Theorem 3.2:
-$e_{h}$ $:=z_{h}-P_{h}z,$
-$\xi_{h}$ $:=S_{0,h}(\Delta_{h}P_{h}-P_{h}\Delta)z,$
-$\eta_{h}$ $:=e_{h}-\xi_{h}.$
-It is straightforward to confirm the following decomposition for $y-y_{h}$:
-$y-y_{h}=(z-P_{h}z)+(G-G_{h})-(\xi_{h}+\eta_{h}).$ (4.45)
+Let $G_{h}$ be defined by Eq. 4.36. We adopt the same notation as introduced in the proof of Theorem 3.2:
+$z_{h}$ $:=y_{h}-G_{h},$ $e_{h}:=z_{h}-P_{h}z,$
+$\xi_{h}$ $:=S_{0,h}(\Delta_{h}P_{h}-P_{h}\Delta)z,$ $\eta_{h}:=e_{h}-\xi_{h}.$
 The remainder of the proof is structured into the following five steps.
@@
 $\left\|G-G_{h}\right\|_{C([0,T];L^{q})}+\left\|P_{h}G-G_{h}\right\|_{C([0,T];L^{q})}\leqslant ch^{2},\quad\forall q\in[2,\infty).$ (4.49)
-Estimate (4.46) is a direct consequence of (4.3) combined with the regularity property (4.43), and estimate (4.47) follows from (4.6) and the condition $v\in W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$.
-Analogous to the estimate Eq.
-4.27, we employ Lemma 4.5, along with (4.4) and the regularity assertion (4.43), to bound $\xi_{h}$ as follows:
+Estimate (4.48) is a direct consequence of (4.19) combined with the regularity property (4.46).
+Estimate (4.49) follows from (4.22) and the condition $v\in W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$.
+In analogy to the estimate (4.29), we utilize Lemma 4.7, along with (4.20) and the regularity assertion (4.46), to derive the following bound for $\xi_{h}$:
 $\left\|\xi_{h}\right\|_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2},\quad\forall p\in(2,\infty),\,\forall q\in[2,\infty).$ (4.50)
-Adopting the reasoning from Theorem 3.2, we establish the following bound for $\eta_{h}$:
+Following the reasoning from Theorem 3.2, we obtain the following estimate for $\eta_{h}$:
 $\max_{t\in[0,T]}\left\|\eta_{h}(t)\right\|_{L^{p}(\Omega;L^{2})}\leqslant ch^{2},\quad\forall p\in(2,\infty).$ (4.51)
-In light of the decomposition Eq. 4.45, combining Eqs. 4.48, 4.49, 4.47 and 4.46 leads to the error estimate
+In light of the decomposition Eq. 4.43, combining Eqs. 4.50, 4.51, 4.49 and 4.48 leads to the error estimate
 $\left\|y-y_{h}\right\|_{L^{p}(\Omega\times(0,T);L^{2})}\leqslant ch^{2},\quad\forall p\in(2,\infty).$ (4.52)
@@
 where in (i) we apply the inverse estimate $\lVert{v_{h}}\rVert_{L^{q}}\leqslant ch^{3(1/q-1/2)}\lVert{v_{h}}\rVert_{L^{2}}$ for all $v_{h}\in X_{h}$; in (ii) we utilize the stability property of $P_{h}$ as stated in Eq.
-4.2; and in (iii) we employ the error estimate Eq.
-4.50.
+4.18; and in (iii) we employ the error estimate Eq.
+4.52.
 By invoking the regularity assertion Eq.
-4.42, we subsequently conclude a uniform boundedness with respect to the spatial mesh size $h$:
-$\lVert{y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\text{ is uniformly bounded w.r.t~{}$h$ for all $p,q\in[2,\infty)$}.$ (4.51)
+4.45, we subsequently conclude a uniform boundedness with respect to the spatial mesh size $h$:
+$\lVert{y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\text{ is uniformly bounded with respect to $h$ for all $p,q\in[2,\infty)$}.$ (4.53)
 Step 2. Let us prove
 $\lVert{\eta_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{6})}\leqslant ch^{2},\quad\forall p\in(2,\infty).$ (4.54)
-Using the equality Eq. 4.35, along with the definitions of the operators $S_{0,h}$ and $S_{1,h}$, we deduce the following decomposition:
-$\eta_{h}=I_{1}+I_{2}+I_{3},$ (4.53)
+Using the equality Eq. 4.38, along with the definitions of the discrete convolution operators $S_{0,h}$ and $S_{1,h}$ given in Eqs. 4.24 and 4.25, we deduce the following decomposition:
+$\eta_{h}=II_{1}+II_{2}+II_{3},$ (4.55)
 where
-$I_{1}$ $:=S_{0,h}\big{(}\xi_{h}+\eta_{h}+G_{h}-P_{h}G\big{)},$
-$I_{2}$ $:=S_{0,h}P_{h}(y^{3}-y_{h}^{3}),$
-$I_{3}$ $:=S_{1,h}P_{h}\big{(}F(y_{h})-F(y)\big{)}.$
-Fix an arbitrary $p\in(2,\infty)$. For the term $I_{1}$, we establish the following bound for any $q\in[2,\infty)$:
-$\lVert{I_{1}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}$ $\leqslant c\lVert{S_{0,h}\big{(}\xi_{h}+\eta_{h}+G_{h}-P_{h}G\big{)}}\rVert_{L^{p}(\Omega\times(0,T);\dot{H}_{h}^{2,2})}$
+$II_{1}$ $:=S_{0,h}\big{(}\xi_{h}+\eta_{h}+G_{h}-P_{h}G\big{)},$
+$II_{2}$ $:=S_{0,h}P_{h}(y^{3}-y_{h}^{3}),$
+$II_{3}$ $:=S_{1,h}P_{h}\big{(}F(y_{h})-F(y)\big{)}.$
+Fix an arbitrary $p\in(2,\infty)$. For the term $II_{1}$, we establish the following bound for any $q\in[2,\infty)$:
+$\lVert{II_{1}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}$ $\leqslant c\lVert{S_{0,h}\big{(}\xi_{h}+\eta_{h}+G_{h}-P_{h}G\big{)}}\rVert_{L^{p}(\Omega\times(0,T);\dot{H}_{h}^{2,2})}$
 $\leqslant c\lVert{\xi_{h}+\eta_{h}+G_{h}-P_{h}G}\rVert_{L^{p}(\Omega\times(0,T);L^{2})}$
 $\leqslant ch^{2},$
-where the first inequality uses the discrete Sobolev’s inequality $\lVert{v_{h}}\rVert_{L^{q}}\leqslant c\lVert{v_{h}}\rVert_{\dot{H}_{h}^{2,2}}$ for all $v_{h}\in X_{h}$; the subsequent inequality relies on Lemma 4.5; and the final inequality is justified by Eqs.
-4.49, 4.47 and 4.48.
-Turning to $I_{2}$, we establish the bound: for any $q\in[2,\infty)$,
-$\lVert{I_{2}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}$ $\leqslant c\lVert{S_{0,h}P_{h}(y^{3}-y_{h}^{3})}\rVert_{L^{p}(\Omega\times(0,T);\dot{H}_{h}^{2,5/3})}$
+where the first inequality uses the discrete Sobolev’s inequality $\lVert{v_{h}}\rVert_{L^{\infty}}\leqslant c\lVert{v_{h}}\rVert_{\dot{H}_{h}^{2,2}}$ for all $v_{h}\in X_{h}$; the subsequent inequality relies on Lemma 4.7; and the final inequality is justified by Eqs.
+4.51, 4.49 and 4.50.
+Turning to $II_{2}$, we establish the bound: for any $q\in[2,\infty)$,
+$\lVert{II_{2}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}$ $\leqslant c\lVert{S_{0,h}P_{h}(y^{3}-y_{h}^{3})}\rVert_{L^{p}(\Omega\times(0,T);\dot{H}_{h}^{2,5/3})}$
 $\leqslant c\lVert{y^{3}-y_{h}^{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{5/3})}$
@@
 $\leqslant ch^{2},$
-where the first inequality uses the discrete Sobolev’s inequality $\lVert{v_{h}}\rVert_{L^{q}}\leqslant c\lVert{v_{h}}\rVert_{\dot{H}_{h}^{2,5/3}}$ for all $v_{h}\in X_{h}$; the second inequality uses Lemma 4.5 and Eq.
-4.2; the third inequality uses Hölder’s ineqality; and the last inequality uses Eqs.
-4.42, 4.50 and 4.51.
-For $I_{3}$ we demonstrate the following bound:
-$\lVert{I_{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{6})}$
-[…]
-[…]
+where the first inequality uses the discrete Sobolev’s inequality $\lVert{v_{h}}\rVert_{L^{\infty}}\leqslant c\lVert{v_{h}}\rVert_{\dot{H}_{h}^{2,5/3}}$ for all $v_{h}\in X_{h}$; the second inequality uses Lemma 4.7 and Eq.
+4.18; the third inequality uses Hölder’s inequality; and the last inequality uses Eqs.
+4.45, 4.52 and 4.53.
+For $II_{3}$ we establish the following bound:
+$\lVert{II_{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{6})}$ $\leqslant c\lVert{S_{1,h}P_{h}\big{(}F(y_{h})-F(y)\big{)}}\rVert_{L^{p}(\Omega\times(0,T);\dot{H}_{h}^{1,2})}$
+$\leqslant c\lVert{y-y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{2})}$
 […]
 where the first inequality employs the discrete Sobolev’s inequality $\lVert{v_{h}}\rVert_{L^{6}}\leqslant c\lVert{v_{h}}\rVert_{\dot{H}_{h}^{1,2}}$ for all $v_{h}\in X_{h}$, and the second inequality invokes Lemma 4.8, along with Eqs.
-4.2 and 4.10.
-Given that the decomposition Eq.
-4.53 implies
-$\lVert{\eta_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{6})}$ $\leqslant\lVert{I_{1}}\rVert_{L^{p}(\Omega\times(0,T);L^{6})}+\lVert{I_{2}}\rVert_{L^{p}(\Omega\times(0,T);L^{6})}$
-$\qquad{}+\lVert{I_{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{6})},$
-combining the above bounds for $I_{1}$, $I_{2}$ and $I_{3}$ yields Eq. 4.52.
-Step 3. Let us proceed to prove the error estimate Eq. 3.7. Fix any $p\in(2,\infty)$ and $q\in[2,\infty)$. We start by establishing the following error estimate:
-$\lVert{y-y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{6})}$ $\leqslant\lVert{z-P_{h}z}\rVert_{L^{p}(\Omega\times(0,T);L^{6})}+\lVert{G-G_{h}}\rVert_{L^{p}(0,T;L^{6})}+{}$
-[…]
-[…]
-Employing this error estimate and following the proof methodology for bounding $I_{3}$ in Step 2, we further obtain an enhanced bound for $I_{3}$:
-$\lVert{I_{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2}.$
-Integrating this refined bound with those previously derived bounds for $I_{1}$ and $I_{2}$ in Step 2, and considering the decomposition in Eq. 4.53, we conclude that
-$\lVert{\eta_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2}.$ (4.54)
-Furthermore, we infer that
-$\lVert{y-y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}$ $\leqslant\lVert{z-P_{h}z}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}+\lVert{G-G_{h}}\rVert_{L^{p}(0,T;L^{q})}+{}$
-[…]
-[…]
-Consequently, combining these estimates directly yields the desired error estimate Eq. 3.7.
+4.18 and 4.4.
+In view of the decomposition Eq.
+4.55, combining the above bounds for $II_{1}$, $II_{2}$, and $II_{3}$ yields Eq.
+4.54.
+Step 3.
+We proceed to establish the error estimate (3.7).
+Considering the decomposition of $y_{h}-y$ given in (4.43), we integrate the inequalities (4.48), (4.49), (4.50), and (4.54) to validate (3.7) for the specific case where $p\in(2,\infty)$ and $q=6$.
+Building upon this outcome and employing a similar bounding technique as utilized for $II_{3}$ in Step 2, we obtain an enhanced bound for $II_{3}$:
+$\lVert{II_{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2},\quad\forall p\in(2,\infty),\,\forall q\in[2,\infty).$
+Integrating this refined bound with the previously established bounds for $II_{1}$ and $II_{2}$ from Step 2, and incorporating the decomposition (4.55), we arrive at the following conclusion:
+$\lVert{\eta_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2},\quad\forall p\in(2,\infty),\,\forall q\in[2,\infty).$ (4.56)
+Ultimately, revisiting the decomposition (4.43) and synthesizing the inequalities (4.48), (4.49), (4.50), and (4.56), we affirm the validity of the error estimate (3.7) for all $p\in(2,\infty)$ and $q\in[2,\infty)$.
 Step 4. Let us demonstrate the following inequality:
@@
 To this end, let $p\in(2,\infty)$ and $q\in[2,\infty)$ be arbitrarily chosen.
-Invoking the definitions of $I_{1}$ and $I_{2}$ from (4.53), in conjunction with Lemma 4.10 and the continuous embedding of $\dot{H}_{h}^{\varepsilon,q}$ into $L^{q}$ for any $\varepsilon>0$, we deduce
-$\lVert{I_{1}+I_{2}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant c\lVert{\xi_{h}+\eta_{h}+G_{h}-P_{h}G+P_{h}(y^{3}-y_{h}^{3})}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}.$
+Invoking the definitions of $II_{1}$ and $II_{2}$ from (4.55), in conjunction with Lemma 4.9 and the continuous embedding of $\dot{H}_{h}^{1,q}$ into $L^{q}$, we deduce
+$\lVert{II_{1}+II_{2}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant c\lVert{\xi_{h}+\eta_{h}+G_{h}-P_{h}G+P_{h}(y^{3}-y_{h}^{3})}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}.$
 Relying on (4.49), (4.50), and (4.56), it follows that
-$\lVert{I_{1}+I_{2}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2}+c\lVert{P_{h}(y^{3}-y_{h}^{3})}\rVert_{L^{p}(\Omega\times(0,T);L^{q})},$
+$\lVert{II_{1}+II_{2}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2}+c\lVert{P_{h}(y^{3}-y_{h}^{3})}\rVert_{L^{p}(\Omega\times(0,T);L^{q})},$
 which, combined with the stability property of $P_{h}$ as stated in (4.18), leads to
-$\lVert{I_{1}+I_{2}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2}+c\lVert{y^{3}-y_{h}^{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}.$
+$\lVert{II_{1}+II_{2}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2}+c\lVert{y^{3}-y_{h}^{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}.$
 Moreover, using Hölder’s inequality yields
@@
 $\lVert{y^{3}-y_{h}^{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant c\lVert{y-y_{h}}\rVert_{L^{2p}(\Omega\times(0,T);L^{2q})}.$
-Consequently, we obtain the following bound for $I_{1}+I_{2}$:
-$\lVert{I_{1}+I_{2}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2}+c\lVert{y-y_{h}}\rVert_{L^{2p}(\Omega\times(0,T);L^{2q})}.$
-We proceed to bound the term $I_{3}$ from Eq. 4.53 as follows:
-$\lVert{I_{3}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}$ $\leqslant c\lVert{P_{h}(F(y_{h})-F(y))}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}$
+Consequently, we obtain the following bound for $II_{1}+II_{2}$:
+$\lVert{II_{1}+II_{2}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2}+c\lVert{y-y_{h}}\rVert_{L^{2p}(\Omega\times(0,T);L^{2q})}.$
+We proceed to bound the term $II_{3}$ from Eq. 4.55 as follows:
+$\lVert{II_{3}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}$ $\leqslant c\lVert{P_{h}(F(y_{h})-F(y))}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}$
 $\leqslant c\lVert{F(y_{h})-F(y)}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}$
@@
 where the first inequality is justified by Lemma 4.8 and the continuous embedding of $(\dot{H}_{h}^{0,q},\dot{H}_{h}^{2,q})_{1/2-1/p,p}$ in $L^{q}$; the second inequality is a consequence of the stability of the projection $P_{h}$ from Eq.
-4.2; and the final inequality relies on the Lipschitz continuity of $F$ as stated in Eq.
-4.10.
-Combining the above bounds for $I_{1}+I_{2}$ and $I_{3}$, we arrive at
-$\lVert{I_{1}+I_{2}+I_{3}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}$
+4.18; and the final inequality relies on the Lipschitz continuity of $F$ as stated in Eq.
+4.4.
+Combining the above bounds for $II_{1}+II_{2}$ and $II_{3}$, we arrive at
+$\lVert{II_{1}+II_{2}+II_{3}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}$
 […]
@@
 4.46 and following the proof technique used for Eq.
-4.28, we deduce the following bound for $\xi_{h}$:
+4.30, we deduce the following bound for $\xi_{h}$:
 $\lVert{\xi_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2-\epsilon}.$
@@
 $\|z-P_{h}z\|_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2-\epsilon}.$
-Furthermore, combining Eqs. 4.47 and 4.55 leads to the following bound:
+Furthermore, by combining Eqs. 4.49 and 4.57, we obtain the following bound:
 $\|G-G_{h}+\eta_{h}\|_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2}.$
-Given that the decomposition Eq. 4.45 implies
+Since the decomposition Eq. 4.43 implies
 $\|y-y_{h}\|_{L^{p}(\Omega;C([0,T];L^{q}))}$ $\leqslant\|\xi_{h}\|_{L^{p}(\Omega;C([0,T];L^{q}))}+\|z-P_{h}z\|_{L^{p}(\Omega;C([0,T];L^{q}))}$
 $\qquad{}+\|G-G_{h}+\eta_{h}\|_{L^{p}(\Omega;C([0,T];L^{q}))},$
-the desired error estimate Eq. 3.8 is derived by combining the above bounds. This concludes the proof of Theorem 3.3.
+the desired error estimate Eq. 3.8 follows by combining the above bounds. This concludes the proof of Theorem 3.3.
 $\blacksquare$
 5 Conclusions
-In this work, we have rigorously investigated the regularity properties of the three-dimensional stochastic Allen-Cahn equation with multiplicative noise in the presence of non-smooth initial data.
+In this work, we have rigorously investigated the regularity properties of the three-dimensional stochastic Allen-Cahn equation with multiplicative noise, particularly in the presence of non-smooth initial data.
 For the spatial semidiscretization using the finite element method, we have established a convergence rate of $O(h^{\alpha/p}+h^{2/p}t^{-1/p})$ with non-smooth initial data.
-Meanwhile, under the presence of smooth initial data, employing the discrete stochastic maximal $L^{p}$-regularity theory, we have achieved a convergence rate of $O(h^{2})$ in the norm of $L^{p}(\Omega\times(0,T);L^{q})$, alongside a pathwise uniform convergence rate of $O(h^{2-\epsilon})$.
-Significantly, these findings are inclusive of diverse spatial $L^{q}$-norms, enhancing their adaptability and universality.
-The derived theoretical results set a foundation for the numerical analysis of fully discrete approximations based on the finite element method for the three-dimensional stochastic Allen-Cahn equation with multiplicative noise.
+Meanwhile, under the condition of smooth initial data, employing the discrete stochastic maximal $L^{p}$-regularity theory, we have achieved a convergence rate of $O(h^{2})$ in the $L^{p}(\Omega\times(0,T);L^{q})$ norm, along with a pathwise uniform convergence rate of $O(h^{2-\epsilon})$.
+Significantly, these findings cover a variety of spatial $L^{q}$-norms, enhancing their adaptability and universality.
+The theoretical results derived here lay a foundation for the numerical analysis of fully discrete approximations based on the finite element method for the three-dimensional stochastic Allen-Cahn equation with multiplicative noise, especially within the context of general spatial $L^{q}$-norms.

### v8

@@
 1 Introduction
-Let $\mathcal{O}$ be a bounded convex domain in $\mathbb{R}^{3}$ with a smooth boundary $\partial\mathcal{O}$, and let $\Delta$ denote the Laplace operator with homogeneous Dirichlet boundary conditions.
-We consider the following stochastic Allen-Cahn equation:
-$\mathrm{d}y(t)=(\Delta y+y-y^{3})(t)\,\mathrm{d}t+F(y(t))\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T,$ (1.1a)
-$y(0)=v,$ (1.1b)
-where the initial value $v$ is given, and $W_{H}$ is an infinite-dimensional $H$-cylindrical Brownian motion.
+Let $\mathcal{O}$ be a bounded, convex domain in $\mathbb{R}^{3}$ with a smooth boundary $\partial\mathcal{O}$, and let $\Delta$ denote the Laplace operator equipped with homogeneous Dirichlet boundary conditions.
+We consider the stochastic Allen–Cahn equation given by
+$\mathrm{d}y(t)=\left(\Delta y+y-y^{3}\right)(t)\,\mathrm{d}t+F(y(t))\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T,$ (1.1a)
+$y(0)=y_{0},$ (1.1b)
+where $y_{0}$ is the given initial condition, $W_{H}$ is an $H$-cylindrical Brownian motion, and the operator $F$ will be specified later.
 The stochastic Allen-Cahn equations are widely used to model phase transitions in materials science; see e.g., [4, 11, 18, 26, 29].
@@
 [10] analyzed finite element approximations of a stochastic Allen-Cahn equation with gradient-type multiplicative noise.
-Majee and Prohl [23] derived strong convergence rates for a stochastic Allen-Cahn equation driven by a multiplicative finite-dimensional Wiener process.
+Majee and Prohl [25] derived strong convergence rates for a stochastic Allen-Cahn equation driven by a multiplicative Wiener process.
 Bréhier et al.
-[5] analyzed an explicit temporal splitting numerical scheme for a stochastic Allen-Cahn equation driven by an additive infinite-dimensional Wiener process.
+[5] analyzed an explicit temporal splitting numerical scheme for a stochastic Allen-Cahn equation driven by an additive Wiener process.
 Jentzen and Pusnik [17] proposed a numerical method for approximating a class of semilinear stochastic evolution equations with non-globally Lipschitz continuous nonlinearities.
-Qi and Wang [31] derived strong approximation errors for a stochastic Allen-Cahn equation driven by an additive infinite-dimensional Wiener process.
-Liu and Qiao [21] established a general theory of the numerical analysis for the stochastic Allen-Cahn equations driven by a multiplicative infinite-dimensional Wiener process.
-Qi et al.
-[32] proposed an efficient approximation method to the stochastic Allen-Cahn equation with random diffusion coefficient field and multiplicative noise.
-The main motivation of this work includes two components:
-  • To the best of our knowledge, existing works on the numerical analysis of the three-dimensional stochastic Allen-Cahn equation have not adequately addressed scenarios involving non-smooth initial data.
-This scenario is of practical significance, as initial phases may include randomly distributed particles.
-The inclusion of non-smooth initial data introduces heightened complexity to numerical analysis compared to cases with smooth initial data.
-Firstly, the theoretical analysis of the three-dimensional stochastic Allen-Cahn equation with multiplicative noise and non-smooth initial data is relatively scarce (see, e.g., [18, 19]).
-Secondly, the singularity introduced by non-smooth initial data significantly complicates the process of numerical analysis.
-  • In the existing literature, the numerical analysis of the three-dimensional stochastic Allen-Cahn equation with multiplicative noise primarily focuses on the spatial $L^{2}$-norm.
-However, error estimates using the more general spatial $L^{q}$-norms are relatively scarce.
-The main contributions of this work are as follows:
+Qi and Wang [28] derived strong approximation errors for a stochastic Allen-Cahn equation driven by an additive Wiener process.
+Liu and Qiao [23] established a general theory of the numerical analysis for the stochastic Allen-Cahn equations with multiplicative noise.
+The motivation for this work stems from two important gaps in the existing literature:
+  • To the best of our knowledge, the numerical analysis of the three-dimensional stochastic Allen–Cahn equation with non-smooth initial data remains underdeveloped, despite its practical significance.
+Existing studies often impose restrictive regularity assumptions on the initial data.
+For instance, the analysis in [23] requires that the initial condition $y_{0}$ belongs to the space $\dot{H}^{\alpha,2}$ for some $\alpha\in[1,2]$, while [28] assumes $\dot{H}^{1,2}$-regularity; the precise definition of these spaces is given in Section 2.
+The treatment of non-smooth initial data introduces substantial theoretical and numerical challenges.
+On the one hand, there is a lack of rigorous results for the three-dimensional stochastic Allen–Cahn equation with multiplicative noise and rough initial conditions (see, e.g., [20, 21]).
+On the other hand, the presence of singularities due to non-smooth initial data significantly complicates the numerical analysis.
+  • Current numerical studies on the three-dimensional stochastic Allen–Cahn equation with multiplicative noise predominantly focus on error estimates in the $L^{2}$-norm in space.
+However, the analysis of numerical approximations in more general spatial $L^{q}$-norms (for $q\geqslant 2$) remains largely unaddressed, which limits the scope and applicability of the theoretical results.
+The main contributions of this work are summarized as follows:
   • For the model problem Eq.
-1.1 with non-smooth initial data $v\in L^{\infty}$, we derive some regularity estimates within the framework of the general spatial $L^{q}$-norms, which are crucial for the numerical analysis.
-Specifically, we affirm that
-    $y\in L^{p}(\Omega;C([0,T];L^{q})),\quad\forall p,q\in(2,\infty).$
-    Our analysis uses a suite of primary tools, including: (i) stochastic integration and Itô’s formula in UMD spaces [7]; (ii) the stochastic maximal $L^{p}$-regularity estimates developed in [25, 26]; (iii) the theory of stochastic partial differential equations with Lyapunov conditions from [18, 19]; (iv) the maximal $L^{p}$-regularity theory for deterministic parabolic equations [37].
-  • To handle the singularity caused by the non-smooth initial data, by decomposing the error into a singular part and a relatively non-singular part, we derive the error estimate
-    $\lVert{(y-y_{h})(t)}\rVert_{L^{p}(\Omega;L^{2})}\leqslant c\big{(}h^{\alpha/p}+h^{2/p}t^{-1/p}\big{)},\quad\forall t\in(0,T],$
-    where $p\in(2,\infty)$ and $\alpha\in(1,2)$ can be arbitrarily chosen. This result is a stochastic counterpart of the well-known error estimate in [35, Theoerem 3.2].
-  • Given smooth initial data $v\in W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$, we mainly use the standard theory of finite element methods (see, e.g., [1, 35]), the discrete maximal $L^{p}$-regularity [16], and the discrete stochastic maximal $L^{p}$-regularity [17] to derive two rigorous error estimates within the framework of general spatial $L^{q}$-norms.
-Specifically, for arbitrary parameters $p\in(2,\infty)$ and $q\in[2,\infty)$, and for any chosen $\epsilon>0$, we establish the following error estimates:
-    $\lVert{y-y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2},$
-    $\lVert{y-y_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2-\epsilon}.$
-To achieve a second-order spatial convergence rate, prior numerical analyses in the literature have typically enforced stringent regularity assumptions on the diffusion coefficients.
-In contrast, this study imposes relatively moderate regularity conditions on the diffusion coefficient $F$ (see Section 2).
-While our numerical analysis assumes deterministic initial data, extending the results to accommodate random initial data is straightforward.
-The remainder of this paper is organized as follows.
-Section 2 introduces the notational conventions and definitions that are used throughout the paper.
-Section 3 presents the main results of this work, including the regularity properties of the mild solution to model problem 1.1 with non-smooth initial data, and establishes the convergence of the finite element-based spatial semidiscretization for both non-smooth and smooth initial conditions.
-A rigorous proof of these results is provided in Section 4.
-Finally, Section 5 concludes the paper by summarizing the key findings of this study.
+1.1 with non-smooth initial data $y_{0}\in L^{\infty}(\mathcal{O})$, we establish essential regularity estimates within the general framework of spatial $L^{q}$-norms.
+These estimates are crucial for the subsequent numerical analysis.
+Specifically, we show that the mild solution satisfies
+    $y\in L_{\mathbb{F}}^{p}(\Omega;C([0,T];L^{q}(\mathcal{O})))\quad\text{for all }p,q\in(2,\infty).$
+    Our analysis combines key tools from both stochastic and deterministic PDE theory, including: stochastic integration and Itô’s formula in UMD Banach spaces [8]; stochastic maximal $L^{p}$-regularity estimates [32, 33]; the Lyapunov framework for stochastic partial differential equations [20, 21]; and maximal $L^{p}$-regularity estimates for deterministic evolution equations [36].
+  • To address the singularity induced by the non-smooth initial data, we employ a decomposition technique that separates the error into a singular and a more regular component.
+This allows us to derive the error estimate
+    $\lVert{(y-y_{h})(t)}\rVert_{L^{p}(\Omega;L^{2}(\mathcal{O}))}\leqslant c\big{(}h^{2/p-\epsilon}+h^{2/p}t^{-1/p}\big{)},\quad\forall t\in(0,T],\,p\in(2,\infty),$
+    where $y_{h}$ is the strong solution of the spatial semidiscretization, $\epsilon>0$ can be arbitrarily small, and $c>0$ is a constant independent of $t$ and the spatial mesh size $h$.
+This result can be viewed as a stochastic counterpart of the classical deterministic error estimate given in [31, Theorem 3.2].
+  • For smooth initial data $y_{0}\in W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$, we establish the optimal error estimate
+    $\lVert{y-y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q}(\mathcal{O}))}\leqslant ch^{2},$
+    and the nearly optimal pathwise uniform error estimate
+    $\lVert{y-y_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{\infty}(\mathcal{O}))}\leqslant ch^{2-\epsilon}\quad\forall\epsilon>0,$
+    where $p,q\in(2,\infty)$ and $c>0$ is a generic constant independent of the spatial mesh size $h$.
+The analysis combines classical finite element theory [1, 31] with discrete maximal $L^{p}$-regularity [12] and discrete stochastic maximal $L^{p}$-regularity [19].
+To the best of our knowledge, this work provides the first result on the spatial convergence of numerical approximations for the three-dimensional stochastic Allen–Cahn equation with multiplicative noise, established within the general framework of spatial $L^{q}$-norms.
+Although we assume deterministic initial data, the analysis can be naturally extended to the case of random initial conditions.
+The main contribution of this study lies in developing a rigorous and general numerical framework for spatial semidiscretizations of nonlinear stochastic parabolic equations under the framework of general spatial $L^{q}$-norms, thereby distinguishing it from previous works.
+The remainder of this paper is structured as follows.
+Section 2 introduces notation, defines the operator $F$, presents the mild solution to the model problem Eq.
+1.1, and describes its spatial semidiscretization.
+Section 3 states the main results of this work, analyzing the regularity properties of the mild solution under non-smooth initial data, and establishes the convergence of the spatial semidiscretization for both non-smooth and smooth initial conditions.
+Section 4 provides the rigorous proofs of these results.
+Finally, Section 5 concludes the paper with a summary of the key findings.
 2 Preliminaries
 Notational Conventions.
-Throughout this paper, we will use the following conventions.
-For Banach spaces $E_{1}$ and $E_{2}$, the space of all bounded linear operators from $E_{1}$ to $E_{2}$ is denoted by $\mathcal{L}(E_{1},E_{2})$, and we abbreviate this to $\mathcal{L}(E_{1})$ when $E_{1}=E_{2}$.
-The identity operator is indicated by $I$.
-Given $\theta\in(0,1)$ and $p\in(1,\infty)$, the notation $(\cdot,\cdot)_{\theta,p}$ is used to refer to the real interpolation space constructed via the $K$-method, as detailed in Chapter 1 of [22].
-The conjugate exponent of each $p\in[0,\infty]$ is denoted by $p^{\prime}$.
-The symbol $\mathcal{O}$ represents a bounded convex domain in $\mathbb{R}^{3}$ with a boundary $\partial\mathcal{O}$ that belongs to the $C^{2,0}$ class (see Section 6.2 of [11]).
-For $q\in[1,\infty]$, we abbreviate $L^{q}(\mathcal{O})$ to $L^{q}$, and use $W_{0}^{1,q}(\mathcal{O})$ and $W^{2,q}(\mathcal{O})$ to denote two standard Sobolev spaces (see, e.g., [34]).
-Let $(\Omega,\mathcal{F},\mathbb{P})$ represent a complete probability space equipped with a right-continuous filtration $\mathbb{F}:=(\mathcal{F}_{t})_{t\geqslant 0}$, on which a sequence of independent $\mathbb{F}$-adapted Brownian motions $(\beta_{n})_{n\in\mathbb{N}}$ is defined.
-The expectation operator associated with this probability space is denoted by $\mathbb{E}$.
+For Banach spaces $E_{1}$ and $E_{2}$, we denote by $\mathcal{L}(E_{1},E_{2})$ the space of all bounded linear operators from $E_{1}$ to $E_{2}$; in the case $E_{1}=E_{2}$, we write $\mathcal{L}(E_{1})$ for brevity.
+The identity operator on any given space is denoted by $I$.
+For $\theta\in(0,1)$ and $p\in(1,\infty)$, the notation $(\cdot,\cdot)_{\theta,p}$ refers to the real interpolation space defined via the $K$-method, as presented in Chapter 1 of [24].
+Let $\mathcal{O}\subset\mathbb{R}^{3}$ denote a bounded convex domain with $\mathcal{C}^{2}$-regular boundary $\partial\mathcal{O}$, as defined in [37, p.
+39].
+For $q\in[1,\infty]$, we adopt the abbreviation $L^{q}:=L^{q}(\mathcal{O})$ for Lebesgue spaces and denote by $W_{0}^{1,q}(\mathcal{O})$ and $W^{2,q}(\mathcal{O})$ the standard Sobolev spaces, following the conventions established in [14, Chapter 7].
+The Laplace operator $\Delta$ is equipped with homogeneous Dirichlet boundary conditions.
+For $\alpha\geqslant 0$ and $q\in(1,\infty)$, define $\dot{H}^{\alpha,q}$ as the domain of the operator $(-\Delta)^{\alpha/2}$ in $L^{q}$, endowed with the natural norm:
+$\|v\|_{\dot{H}^{\alpha,q}}:=\left\|(-\Delta)^{\alpha/2}v\right\|_{L^{q}},\quad v\in\dot{H}^{\alpha,q}.$
+The dual space of $\dot{H}^{\alpha,q}$ is denoted by $\dot{H}^{-\alpha,q^{\prime}}$, where $q^{\prime}$ is the Hölder conjugate of $q$.
+When $\alpha=2$, $\dot{H}^{2,q}$ coincides with $W_{0}^{1,q}(\mathcal{O})\cap W^{2,q}(\mathcal{O})$, and its norm is equivalent to the $W^{2,q}(\mathcal{O})$-norm (cf.
+[14, Theorem 9.15]).
+For $\alpha=1$, $\dot{H}^{1,q}$ agrees with $W_{0}^{1,q}(\mathcal{O})$, with equivalent norms (see [37, Theorem 16.14]).
+Let $(\Omega,\mathcal{F},\mathbb{P})$ be a complete probability space endowed with a filtration $\mathbb{F}:=(\mathcal{F}_{t})_{t\geqslant 0}$ satisfying the usual conditions.
+On this stochastic basis, we define a sequence $(W_{n})_{n\in\mathbb{N}}$ of independent real-valued $\mathbb{F}$-adapted Brownian motions.
+The associated expectation operator is denoted by $\mathbb{E}$.
 Let $H$ be a separable Hilbert space with inner product $(\cdot,\cdot)_{H}$ and an orthonormal basis $(h_{n})_{n\in\mathbb{N}}$.
-The $H$-cylindrical Brownian motion $W_{H}$ is defined such that, for each $t\in\mathbb{R}_{+}$, $W_{H}(t)\in\mathcal{L}(H,L^{2}(\Omega))$ is defined by
-$W_{H}(t)h:=\sum_{n\in\mathbb{N}}\beta_{n}(t)(h,h_{n})_{H},\quad\forall h\in H.$
-For $p\in(1,\infty)$ and a Banach space $X$, the space of all $\mathbb{F}$-adapted processes in $L^{p}(\Omega\times(0,T);X)$ is denoted by $L^{p}_{\mathbb{F}}(\Omega\times(0,T);X)$.
-For any Banach space $X$, the notation $\gamma(H,X)$ represents the space of all $\gamma$-radonifying operators from $H$ to $X$, as defined in [13, Chapter 9].
-For a comprehensive treatment of the theory of stochastic integration with respect to the $H$-cylindrical Wiener process $W_{H}$, we refer the reader to [24].
+The $H$-cylindrical Brownian motion $W_{H}$ is defined such that, for each $t\in\mathbb{R}_{+}$, $W_{H}(t)\in\mathcal{L}(H,L^{2}(\Omega))$ is given by
+$W_{H}(t)x:=\sum_{n\in\mathbb{N}}W_{n}(t)(x,h_{n})_{H},\quad\forall x\in H.$
+We refer the reader to [34] for the stochastic integration with respect to $W_{H}$.
+Let $p\in(1,\infty)$ and let $E$ be a Banach space. Denote by $L^{p}_{\mathbb{F}}(\Omega\times(0,T);E)$ the closure in the Bochner space $L^{p}(\Omega\times(0,T);E)$ of the linear span of the set
+[…]
+where $\mathbbm{1}_{(s,t]\times F}$ denotes the indicator function of the set $(s,t]\times F$.
+Let $L^{0}_{\mathbb{F}}(\Omega;C([0,T];E))$ denote the space of all continuous, $\mathbb{F}$-adapted, $E$-valued processes.
+Two processes $g_{1}$ and $g_{2}$ in $L_{\mathbb{F}}^{0}(\Omega;C([0,T];E))$ are considered identical if $g_{1}=g_{2}$ in $C([0,T];E)$ holds $\mathbb{P}$-almost surely.
+We denote by $L^{p}_{\mathbb{F}}(\Omega;C([0,T];E))$ the subspace of all processes $g\in L^{0}_{\mathbb{F}}(\Omega;C([0,T];E))$ satisfying
+$\mathbb{E}\big{[}\|g\|_{C([0,T];E)}^{p}\big{]}<\infty,$
+where $C([0,T];E)$ is the space of continuous $E$-valued functions on $[0,T]$, equipped with the supremum norm.
+The space $\gamma(H,E)$ denotes the space of $\gamma$-radonifying operators from the Hilbert space $H$ to $E$ (see, e.g., [16, Chapter 9]).
+If $E$ is itself a Hilbert space, then $\gamma(H,E)$ is isometrically isomorphic to the space of Hilbert–Schmidt operators from $H$ to $E$.
 Definition of the Operator $F$.
-Let $(f_{n})_{n\in\mathbb{N}}$ be a sequence of continuous functions mapping from $\overline{\mathcal{O}}\times\mathbb{R}$ to $\mathbb{R}$, which satisfy the following conditions for every $n\in\mathbb{N}$:
-$f_{n}(x,0)=0,$ $\forall x\in\partial\mathcal{O},$ (2.1)
-$\lvert{f_{n}(x,y)}\rvert\leqslant C_{F}(1+\lvert{y}\rvert),$ $\forall x\in\mathcal{O},\,\forall y\in\mathbb{R},$ (2.2)
-$\lvert{f_{n}(x,y_{1})-f_{n}(x,y_{2})}\rvert\leqslant C_{F}\lvert{y_{1}-y_{2}}\rvert,$ $\forall x\in\mathcal{O},\,\forall y_{1},y_{2}\in\mathbb{R},$ (2.3)
-$\lvert{f_{n}(x_{1},y)-f_{n}(x_{2},y)}\rvert\leqslant C_{F}(1+\lvert{y}\rvert)\lvert{x_{1}-x_{2}}\rvert,$ $\forall x_{1},x_{2}\in\mathcal{O},\,\forall y\in\mathbb{R},$ (2.4)
-where $C_{F}$ is a positive constant independent of $n$, $x$, $y$, $x_{1}$, $x_{2}$, $y_{1}$, and $y_{2}$.
-Hereafter, with a slight abuse of notation, we will also employ $f_{n}$ to denote the Nemytskii operator defined by
-$f_{n}(v):=f_{n}(\cdot,v(\cdot)),\quad\forall v\in L^{1}.$
-Let $(\lambda_{n})_{n\in\mathbb{N}}$ be a sequence of non-negative numbers satisfying the summability condition
-$\sum_{n\in\mathbb{N}}\lambda_{n}<\infty.$ (2.5)
-For any $v\in L^{2}$, we define $F(v)\in\gamma(H,L^{2})$ as
-$F(v):=\sum_{n=0}^{\infty}\sqrt{\lambda}_{n}h_{n}\otimes f_{n}(v).$ (2.6)
-Accordingly, the model equation (1.1) can be explicitly formulated as follows:
-$\begin{cases}\mathrm{d}y(t)=(\Delta y+y-y^{3})(t)\,\mathrm{d}t+\sum_{n=0}^{\infty}\sqrt{\lambda_{n}}f_{n}(y(t))\,\mathrm{d}\beta_{n}(t),\quad 0\leqslant t\leqslant T,\\ y(0)=v.\end{cases}$
-Mild Solutions. Let $p,q\in[2,\infty)$. We introduce the deterministic convolution operator $S_{0}$ acting on functions $g\in L^{p}(0,T;L^{q})$ as follows:
-$(S_{0}g)(t):=\int_{0}^{t}e^{(t-s)\Delta}g(s)\,\mathrm{d}s,\quad\text{for }0\leqslant t\leqslant T.$ (2.7)
-Additionally, for any process $g\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,L^{q}))$, we define the stochastic convolution operator $S_{1}$ by
-$(S_{1}g)(t):=\int_{0}^{t}e^{(t-s)\Delta}g(s)\,\mathrm{d}W_{H}(s),\quad\text{for }0\leqslant t\leqslant T.$ (2.8)
-Here, $e^{t\Delta}$, for $t\in[0,\infty)$, denotes the analytic semigroup generated by the Laplace operator $\Delta$ subject to homogeneous Dirichlet boundary conditions.
-A process $y\in L_{\mathbb{F}}^{6}(\Omega\times(0,T);L^{6})$ is called a mild solution to equation (1.1) if it satisfies the equality
-$y=G+S_{0}(y-y^{3})+S_{1}F(y),$ (2.9)
-where the function $G$ is given by
-$G(t):=e^{t\Delta}v,\quad\text{for }t\in[0,T].$ (2.10)
-Note that the conditions in the above definition of a mild solution can be relaxed (see, e.g., [29, Chapter 6]). However, the present formulation is adequate for the purposes of this study.
+Let $(f_{n})_{n\in\mathbb{N}}$ be a sequence of real-valued, continuously differentiable functions defined on $\overline{\mathcal{O}}\times\mathbb{R}$, satisfying the following conditions:
+$\sum_{n\in\mathbb{N}}|f_{n}(x,0)|^{2}=0,$ $\forall x\in\partial\mathcal{O},$ (2.1)
+$\sum_{n\in\mathbb{N}}\big{(}|f_{n}(x,y)|^{2}+|\nabla_{x}f_{n}(x,y)|^{2}\big{)}\leqslant C_{F}(1+|y|^{2}),$ $\forall x\in\mathcal{O},\,\forall y\in\mathbb{R},$ (2.2)
+$\sum_{n\in\mathbb{N}}|\nabla_{y}f_{n}(x,y)|^{2}\leqslant C_{F},$ $\forall x\in\mathcal{O},\,\forall y\in\mathbb{R},$ (2.3)
+where $C_{F}>0$ is a constant independent of $x$ and $y$. Here and in the sequel, we slightly abuse notation by identifying each $f_{n}$ with the corresponding Nemytskii operator defined pointwise by
+$f_{n}(v):=f_{n}(\cdot,v(\cdot)),\quad\text{for all }v\in L^{2}.$
+Recall that $(h_{n})_{n\in\mathbb{N}}$ denotes a fixed orthonormal basis of the separable Hilbert space $H$.
+Given condition Eq.
+2.2, for each $v\in L^{2}$, define the operator $F(v)\in\gamma(H,L^{2})$ by
+$F(v):=\sum_{n\in\mathbb{N}}h_{n}\otimes f_{n}(v),$ (2.4)
+where the elementary tensor $h_{n}\otimes f_{n}(v)\in\gamma(H,L^{2})$ acts on $x\in H$ via
+$(h_{n}\otimes f_{n}(v))(x):=(x,h_{n})_{H}f_{n}(v).$
+In this setting, the model equation (1.1) takes the explicit form of the stochastic evolution equation:
+$\begin{cases}\mathrm{d}y(t)=\left(\Delta y+y-y^{3}\right)(t)\,\mathrm{d}t+\displaystyle\sum_{n\in\mathbb{N}}f_{n}(y(t))\,\mathrm{d}W_{n}(t),\quad t\in[0,T],\\ y(0)=y_{0},\end{cases}$
+where $\Delta$ denotes the Laplace operator with homogeneous Dirichlet boundary conditions.
+Mild Solutions. We define the two convolution operators:
+$(S_{0}g)(t)$ $:=\int_{0}^{t}e^{(t-s)\Delta}g(s)\,\mathrm{d}s,\quad 0\leqslant t\leqslant T,$ (2.5)
+$(S_{1}g)(t)$ $:=\int_{0}^{t}e^{(t-s)\Delta}g(s)\,\mathrm{d}W_{H}(s),\quad 0\leqslant t\leqslant T,$ (2.6)
+whenever the integrals are well-defined. Here, $(e^{t\Delta})_{t\geqslant 0}$ denotes the analytic semigroup generated by the Laplace operator $\Delta$ under homogeneous Dirichlet boundary conditions.
+Assume that $y_{0}:(\Omega,\mathcal{F}_{0},\mathbb{P})\to L^{4}$ is strongly measurable.
+A process $y\in L_{\mathbb{F}}^{0}(\Omega;C([0,T];L^{4}))$ is called a mild solution to equation (1.1) if it satisfies, $\mathbb{P}$-almost surely for all $t\in[0,T]$,
+$y(t)=G(t)+\big{[}S_{0}(y-y^{3})\big{]}(t)+\big{[}S_{1}F(y)\big{]}(t),$ (2.7)
+where
+$G(t):=e^{t\Delta}y_{0},\quad t\in[0,T].$ (2.8)
 Spatial Semidiscretization.
-Consider $\mathcal{K}_{h}$ to be a conforming and quasi-uniform triangulation of the domain $\mathcal{O}$, composed of three-dimensional simplexes.
-We introduce the finite element space $X_{h}$ as
-$X_{h}:=\Bigl{\{}v_{h}\in C(\overline{\mathcal{O}})\bigm{|}v_{h}=0\text{ on }(\mathcal{O}\setminus\mathcal{O}_{h})\cup\partial\mathcal{O}_{h},\text{ and }v_{h}|_{K}\in P_{r}(K)\text{ for each }K\in\mathcal{K}_{h}\Bigr{\}},$
-where $C(\overline{\mathcal{O}})$ denotes the space of continuous functions on the closure of $\mathcal{O}$, $\mathcal{O}_{h}$ is the closure of the union of the simplexes in $\mathcal{K}_{h}$, $r\geqslant 1$ is a positive integer, and $P_{r}(K)$ is the space of polynomials on simplex $K$ up to degree $r$.
-Let the spatial mesh size $h$ be the maximum diameter of the elements in $\mathcal{K}_{h}$.
-Define $P_{h}$ as the $L^{2}$-orthogonal projection operator onto $X_{h}$.
-The discrete Laplace operator $\Delta_{h}:X_{h}\to X_{h}$ is defined by
-$\int_{\mathcal{O}}(\Delta_{h}u_{h})\cdot v_{h}\,\mathrm{d}\mu=-\int_{\mathcal{O}}\nabla u_{h}\cdot\nabla v_{h}\,\mathrm{d}\mu,\quad\text{for all }u_{h},v_{h}\in X_{h},$
-where $\mu$ denotes the three-dimensional Lebesgue measure on $\mathcal{O}$. We consider the following spatial semidiscretization:
-$\mathrm{d}y_{h}(t)=\left(\Delta_{h}y_{h}+y_{h}-P_{h}(y_{h}^{3})\right)(t)\,\mathrm{d}t+P_{h}F(y_{h}(t))\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T,$ (2.11a)
-$y_{h}(0)=P_{h}v,$ (2.11b)
-where $v$ is the given initial value.
-Assume that $v\in L^{2}$.
-Since this spatial semidiscretization is a finite-dimensional stochastic differential equation (SDE), we can use the classical theory of finite-dimensional SDEs (see, e.g., [28, Theorem 3.21]) to conclude that Eq.
-2.11 has a unique strong solution.
+Let $\mathcal{K}_{h}$ be a conforming and quasi-uniform triangulation of the domain $\mathcal{O}$, consisting of three-dimensional simplices.
+We define the finite element space $X_{h}$ as
+$X_{h}:=\Bigl{\{}v_{h}\in C(\overline{\mathcal{O}})\Bigm{|}v_{h}=0\text{ on }\overline{\mathcal{O}}\setminus\mathcal{O}_{h},\text{ and }v_{h}\text{ is piecewise linear with respect to each }K\in\mathcal{K}_{h}\Bigr{\}},$
+where $C(\overline{\mathcal{O}})$ denotes the space of continuous functions on $\overline{\mathcal{O}}$, and $\mathcal{O}_{h}$ is the closure of the union of all elements in $\mathcal{K}_{h}$.
+We assume that every boundary vertex of $\mathcal{O}_{h}$ lies on $\partial\mathcal{O}$.
+The spatial mesh size $h$ is defined as the maximum diameter over all elements in $\mathcal{K}_{h}$.
+Let $P_{h}$ denote the $L^{2}$-orthogonal projection onto $X_{h}$, and define the discrete Laplace operator $\Delta_{h}:X_{h}\to X_{h}$ via the identity
+$\int_{\mathcal{O}}(\Delta_{h}u_{h})v_{h}\,\mathrm{d}x=-\int_{\mathcal{O}}\nabla u_{h}\cdot\nabla v_{h}\,\mathrm{d}x,\quad\text{for all }u_{h},v_{h}\in X_{h}.$
+We consider the following spatial semidiscretization of the model equation:
+$\mathrm{d}y_{h}(t)=\left(\Delta_{h}y_{h}+y_{h}-P_{h}y_{h}^{3}\right)(t)\,\mathrm{d}t+P_{h}F(y_{h}(t))\,\mathrm{d}W_{H}(t),\quad t\in[0,T],$ (2.9a)
+$y_{h}(0)=P_{h}y_{0}.$ (2.9b)
+Suppose that $y_{0}:(\Omega,\mathcal{F}_{0},\mathbb{P})\to L^{2}(\mathcal{O})$ is strongly measurable.
+A process $y_{h}\in L_{\mathbb{F}}^{0}(\Omega;C([0,T];X_{h}))$, where $X_{h}$ is equipped with the $L^{2}$-norm, is called a strong solution to (2.9) if, $\mathbb{P}$-almost surely, the identity
+$y_{h}(t)=y_{h}(0)+\int_{0}^{t}\left(\Delta_{h}y_{h}+y_{h}-P_{h}y_{h}^{3}\right)(s)\,\mathrm{d}s+\int_{0}^{t}P_{h}F(y_{h}(s))\,\mathrm{d}W_{H}(s)$
+holds for all $t\in[0,T]$.
+By the standard stability estimate for $P_{h}$ (see, e.g., Eq.
+4.19), together with the growth and Lipschitz conditions on $F$ given in Lemma 4.5, the classical theory of stochastic differential equations (see, e.g., [27, Theorem 3.21]) ensures the existence and uniqueness of a strong solution $y_{h}$ to the spatial semidiscretization (2.9).
 3 Main results
-In the remainder of this paper, the symbol $c$ denotes a generic positive constant that depends only on $T$, $\mathcal{O}$, $\sum_{n\in\mathbb{N}}\lambda_{n}$, the constant $C_{F}$ in Eqs.
-2.1, 2.2, 2.4 and 2.3, the indices of the relevant Lebesgue, Sobolev, and interpolation spaces, the initial value $v$ given in equations Eq.
-1.1 and Eq.
-2.11, the regularity parameters of the triangulation $\mathcal{K}_{h}$, and the exponents associated with the spatial mesh size $h$.
+In the remainder of this paper, we denote by $c$ a generic positive constant that is independent of the spatial mesh size $h$.
+It may, however, depend on the final time $T$, the domain $\mathcal{O}$, the constant $C_{F}$ from Eqs.
+2.2 and 2.3, the indices of the relevant Lebesgue, Sobolev, and interpolation spaces, the initial data $y_{0}$, the shape-regularity parameters of the triangulation $\mathcal{K}_{h}$, and the exponents associated with $h$, unless otherwise specified.
 The value of $c$ may differ at each occurrence.
-For any $\alpha\in[0,\infty)$ and $q\in(1,\infty)$, let $\dot{H}^{\alpha,q}$ denote the domain of $(-\Delta)^{\alpha/2}$ in $L^{q}$, endowed with the standard norm defined as follows:
-$\|v\|_{\dot{H}^{\alpha,q}}:=\|(-\Delta)^{\alpha/2}v\|_{L^{q}}\quad\text{for all $v\in\dot{H}^{\alpha,q}$}.$
-The dual space of $\dot{H}^{\alpha,q}$ is denoted by $\dot{H}^{-\alpha,q^{\prime}}$.
-It is well-established that $\Delta$ has a unique extension as a bounded linear operator from $\dot{H}^{\alpha,q}$ to $\dot{H}^{\alpha-2,q}$ for all $\alpha\in[0,2)$.
-Specifically, when $\alpha=2$, the space $\dot{H}^{2,q}$ is identical to the intersection $W_{0}^{1,q}(\mathcal{O})\cap W^{2,q}(\mathcal{O})$ and is endowed with a norm equivalent to that of $W^{2,q}(\mathcal{O})$ (see [11, Theorem 9.15]).
-Moreover, for $\alpha=1$, the space $\dot{H}^{1,q}$ coincides with the standard Sobolev space $W_{0}^{1,q}(\mathcal{O})$, and the two norms are equivalent (see [38, Theorem 16.14]).
-The main results of this paper are summarized in the following three theorems.
 Theorem 3.1.
-Assume that $v\in L^{\infty}$. Then equation Eq. 1.1 admits a unique mild solution $y$, which satisfies that
+Assume that $y_{0}\in L^{\infty}$ is deterministic. Then equation Eq. 1.1 admits a unique mild solution $y$, which satisfies that
 $y\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}),\qquad\forall p,q\in(2,\infty).$ (3.1)
-Furthermore, for any $p\in(2,\infty)$, $q\in[2,\infty)$, and $\epsilon>0$, the solution $y$ possesses the following regularity properties:
-$y\in L^{p}\big{(}\Omega;C([0,T];L^{q})\big{)},$ (3.2)
-$y\in L^{p}\big{(}\Omega\times(0,T);\dot{H}^{2/p-\epsilon,q}\big{)},$ (3.3)
-$y-G\in L^{p}\big{(}\Omega;C([0,T];\dot{H}^{1-\epsilon,q})\big{)},$ (3.4)
-$y-G\in L^{p}\big{(}\Omega\times(0,T);\dot{H}^{1+2/p-\epsilon,q}\big{)},$ (3.5)
+Furthermore, for any $p,q\in(2,\infty)$ and $\epsilon>0$, the solution $y$ possesses the following regularity properties:
+$y\in L_{\mathbb{F}}^{p}\big{(}\Omega;C([0,T];L^{q})\big{)},$ (3.2)
+$y\in L_{\mathbb{F}}^{p}\big{(}\Omega\times(0,T);\dot{H}^{2/p-\epsilon,q}\big{)},$ (3.3)
+$y-G\in L_{\mathbb{F}}^{p}\big{(}\Omega;C([0,T];\dot{H}^{1-\epsilon,q})\big{)},$ (3.4)
+$y-G\in L_{\mathbb{F}}^{p}\big{(}\Omega\times(0,T);\dot{H}^{1+2/p-\epsilon,q}\big{)},$ (3.5)
 where $G$ is defined by Eq. 2.8.
 Theorem 3.2.
-Assume that $v\in L^{\infty}$.
+Assume that $y_{0}\in L^{\infty}$ is deterministic.
 Let $y$ be the mild solution of Eq.
 1.1, and let $y_{h}$ be the strong solution to the spatial semidiscretization Eq.
-2.11.
-Then, for any $\alpha\in(1,2)$ and $p\in(2,\infty)$, the following error estimate holds for all $t\in(0,T]$:
-$\lVert{(y-y_{h})(t)}\rVert_{L^{p}(\Omega;L^{2})}\leqslant c\left(h^{\alpha/p}+h^{2/p}t^{-1/p}\right).$ (3.6)
+2.9.
+Then, for any $p\in(2,\infty)$, the following error estimate holds for all $t\in(0,T]$:
+$\lVert{(y-y_{h})(t)}\rVert_{L^{p}(\Omega;L^{2})}\leqslant c\left(h^{2/p-\epsilon}+h^{2/p}t^{-1/p}\right),\quad\forall\epsilon>0.$ (3.6)
 Theorem 3.3.
-Assume that $v\in W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$.
+Assume that $y_{0}\in W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$ is deterministic.
 Let $y$ be the mild solution of Eq.
 1.1, and let $y_{h}$ be the strong solution of Eq.
-2.11.
-Then, for any $p\in(2,\infty)$, $q\in[2,\infty)$, and $\epsilon>0$, the following error estimates hold:
-$\lVert{y-y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2},$ (3.7)
-$\lVert{y-y_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2-\epsilon}.$ (3.8)
+2.9.
+Then the following error estimates hold:
+$\lVert{y-y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2},\quad\forall p,q\in(2,\infty),$ (3.7)
+$\lVert{y-y_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{\infty}))}\leqslant ch^{2-\epsilon},\quad\forall p\in(2,\infty),\,\forall\epsilon>0.$ (3.8)
 It is worth noting that the aforementioned theoretical results remain valid under the assumption that the operator $F$ possesses the growth and Lipschitz continuity properties outlined in Lemma 4.5.
 Remark 3.1.
-Liu [18] proposed a theoretical framework for a broad spectrum of nonlinear stochastic partial differential equations. Applying this framework to our model yields
+Liu [20] proposed a theoretical framework for a broad spectrum of nonlinear stochastic partial differential equations. Applying this framework to the model problem Eq. 1.1 yields
 $y\in L^{p}(\Omega;C([0,T];\dot{H}^{1,2})),\quad\forall p\in(2,\infty),$
-provided that $v\in\dot{H}^{1,2}$.
+provided that $y_{0}\in L^{p}(\Omega,\mathcal{F}_{0},\mathbb{P};\dot{H}^{1,2})$, where $p\in[2,\infty)$.
 Furthermore, Liu and Qiao [23] have established certain regularity estimates in the context of the spatial $L^{2}$-norm, and some regularity estimates with the spatial $L^{\infty}$-norm were derived by using Sobolev’s embedding theorem.
@@
 Remark 3.2.
-For Theorems 3.2 and 3.3, we have the following comments.
-  1. 1.
 In the numerical analysis of the three-dimensional stochastic Allen-Cahn equation with multiplicative noise, the existing literature often imposes strict regularity conditions on the diffusion coefficients.
@@
 1.1 under less restrictive conditions on $F$, especially the loosening of condition (2.1), has important theoretical and practical implications.
-  2. 2.
-    In Section 4.3, we show that the norm of $y_{h}$ in $L^{p}(\Omega\times(0,T);L^{q})$ is uniformly bounded with respect to the spatial mesh size $h$ for all $p,q\in[2,\infty)$, assuming $v\in W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$ (see Eq.
-4.53).
-It can be rigorously demonstrated that this result also holds when $v\in W_{0}^{\epsilon,\infty}(\mathcal{O})$, where $\epsilon>0$ can be any arbitrarily small positive number.
-This finding is critical for the numerical analysis of the fully discrete approximations of Eq.
-1.1.
-However, it remains an open question whether this conclusion applies when $v\in L^{\infty}$.
 4 Proofs
@@
 4.1 Preliminary Lemmas
-It is a well-established result that for any $q\in(1,\infty)$ and $s\in(0,2)$, the space $\dot{H}^{s,q}$ coincides with the complex interpolation space $[L^{q},\dot{H}^{2,q}]_{s/2}$, equipped with equivalent norms.
+It is a well-established result that for any $\alpha\in(0,2)$ and $q\in(1,\infty)$, the space $\dot{H}^{\alpha,q}$ coincides with the complex interpolation space $[L^{q},\dot{H}^{2,q}]_{\alpha/2}$, equipped with equivalent norms.
 This result is documented, for example, in Chapter 16 of [37].
-The combination of Proposition 1.4 from [22] and Theorem C.4.1 in [12] yields the following continuous embeddings.
+The combination of Propositions 1.3 and 1.4 and Theorem 4.17 from [24] and Theorem C.4.1 in [15] yields the following continuous embeddings.
 Lemma 4.1.
-Let $0\leqslant\beta<\gamma\leqslant 1$ and $q\in(1,\infty)$. Then the following continuous embeddings hold:
+Let $0\leqslant\alpha<\beta\leqslant 2$ and $p,q\in(1,\infty)$. Then the following continuous embeddings hold:
   • (i)
-    The space $\dot{H}^{\gamma,q}$ is continuously embedded in $(L^{q},\dot{H}^{1,q})_{\beta,q}$;
+    If $\gamma>(1-\theta)\alpha+\theta\beta$ for some $\theta\in(0,1)$, then $\dot{H}^{\gamma,q}$ is continuously embedded in the real interpolation space $(\dot{H}^{\alpha,q},\dot{H}^{\beta,q})_{\theta,p}$.
   • (ii)
-    The space $(L^{q},\dot{H}^{1,q})_{\gamma,q}$ is continuously embedded in $\dot{H}^{\beta,q}$.
+    If $\gamma<(1-\theta)\alpha+\theta\beta$ for some $\theta\in(0,1)$, then the real interpolation space $(\dot{H}^{\alpha,q},\dot{H}^{\beta,q})_{\theta,p}$ is continuously embedded in $\dot{H}^{\gamma,q}$.
 Theorem 2.12 and inequality (2.128) in [37] provide the following standard estimate for the growth of the analytic semigroup generated by the Laplace operator with homogeneous Dirichlet boundary conditions.
 Lemma 4.2.
-Let $0\leqslant\alpha\leqslant\beta<\infty$ and $q\in(1,\infty)$. For all $t\in(0,T]$, the following estimate holds:
-$\lVert{e^{t\Delta}}\rVert_{\mathcal{L}(\dot{H}^{\alpha,q},\dot{H}^{\beta,q})}\leqslant ct^{(\alpha-\beta)/2}.$
-For the deterministic convolution operator $S_{0}$, defined by Eq.
-2.7, we have the following celebrated deterministic maximal $L^{p}$-regularity estimate; see Theorem 6.1.11 in [30] and Corollary 1.14 from [22].
+Let $\alpha\in[0,\infty)$ and $q\in(1,\infty)$. For all $t>0$, the following estimate holds:
+$\|(-\Delta)^{\alpha}e^{t\Delta}\|_{\mathcal{L}(L^{q})}+t\Big{\|}(-\Delta)^{\alpha}\frac{\mathrm{d}}{\mathrm{d}t}e^{t\Delta}\Big{\|}_{\mathcal{L}(L^{q})}\leqslant ct^{-\alpha}.$
+For the convolution operator $S_{0}$, defined by Eq. 2.5, we have the following celebrated maximal $L^{p}$-regularity estimate; see Theorem 4.2 in [36] and Corollary 1.14 from [24].
 Lemma 4.3.
@@
 Lemma 4.5.
-Let $q\in(1,\infty)$. The following inequalities are valid:
+Let $q\in[2,\infty)$. The following inequalities are valid:
 $\lVert{F(v)}\rVert_{\gamma(H,L^{q})}\leqslant c\left(1+\lVert{v}\rVert_{L^{q}}\right),$ for all $\text{for all }v\in L^{q};$ (4.1)
@@
 where
-$A_{N}v:=\sum_{n=0}^{N}\gamma_{n}\sqrt{\lambda_{n}}f_{n}(v),\quad\forall v\in L^{q}.$
+$A_{N}v:=\sum_{n=0}^{N}\gamma_{n}f_{n}(v),\quad\forall v\in L^{q}.$
 For any $v\in\dot{H}^{1,q}$, we deduce that
-$\lVert{A_{N}v}\rVert_{L^{q}(\Omega_{\gamma};\dot{H}^{1,q})}$ $\leqslant c\bigg{[}\mathbb{E}_{\gamma}\int_{\mathcal{O}}\Big{\lvert}{\sum_{n=0}^{N}\gamma_{n}\sqrt{\lambda_{n}}\nabla f_{n}(v)}\Big{\rvert}^{q}\,\mathrm{d}\mu\bigg{]}^{1/q}$
-$\stackrel{{\scriptstyle\mathrm{(i)}}}{{\leqslant}}c\bigg{[}\int_{\mathcal{O}}\Big{(}\mathbb{E}_{\gamma}\Big{\lvert}{\sum_{n=0}^{N}\gamma_{n}\sqrt{\lambda_{n}}\nabla f_{n}(v)}\Big{\rvert}^{2}\Big{)}^{q/2}\,\mathrm{d}\mu\bigg{]}^{1/q}$
-$=c\bigg{[}\int_{\mathcal{O}}\Big{(}\sum_{n=0}^{N}\lambda_{n}\lvert{\nabla f_{n}(v)}\rvert^{2}\Big{)}^{q/2}\,\mathrm{d}\mu\bigg{]}^{1/q}$
-$\stackrel{{\scriptstyle\mathrm{(ii)}}}{{\leqslant}}c\bigg{[}\int_{\mathcal{O}}\Big{(}\sum_{n=0}^{N}\lambda_{n}(1+\lvert{v}\rvert+\lvert{\nabla v}\rvert)^{2}\Big{)}^{q/2}\,\mathrm{d}\mu\bigg{]}^{1/q},$
-where step $(\mathrm{i})$ uses the famous Kahane-Khintchine inequality (see Theorem 6.2.6 from [13]), and step $(\mathrm{ii})$ relies on Lipschitz conditions Eqs.
-2.3 and 2.4.
-Using the summability condition Eq.
-2.5 and the standard inequality
-$\left(\int_{\mathcal{O}}|v|^{q}+|\nabla v|^{q}\,\mathrm{d}\mu\right)^{1/q}\leqslant c\lVert{v}\rVert_{\dot{H}^{1,q}},$
+$\lVert{A_{N}v}\rVert_{L^{q}(\Omega_{\gamma};\dot{H}^{1,q})}$ $\leqslant c\bigg{[}\mathbb{E}_{\gamma}\int_{\mathcal{O}}\Big{\lvert}{\sum_{n=0}^{N}\gamma_{n}\nabla f_{n}(v)}\Big{\rvert}^{q}\,\mathrm{d}x\bigg{]}^{1/q}$
+$\stackrel{{\scriptstyle\mathrm{(i)}}}{{\leqslant}}c\bigg{[}\int_{\mathcal{O}}\Big{(}\mathbb{E}_{\gamma}\Big{\lvert}{\sum_{n=0}^{N}\gamma_{n}\nabla f_{n}(v)}\Big{\rvert}^{2}\Big{)}^{q/2}\,\mathrm{d}x\bigg{]}^{1/q}$
+$=c\bigg{[}\int_{\mathcal{O}}\Big{(}\sum_{n=0}^{N}\lvert{\nabla f_{n}(v)}\rvert^{2}\Big{)}^{q/2}\,\mathrm{d}x\bigg{]}^{1/q}$
+$\stackrel{{\scriptstyle\mathrm{(ii)}}}{{\leqslant}}c\bigg{[}\int_{\mathcal{O}}\Big{(}1+\lvert{v}\rvert^{2}+\lvert{\nabla v}\rvert^{2}\Big{)}^{q/2}\,\mathrm{d}x\bigg{]}^{1/q},$
+where step $(\mathrm{i})$ uses the famous Kahane-Khintchine inequality (see Theorem 6.2.6 from [16]), and step $(\mathrm{ii})$ relies on conditions Eqs. 2.3 and 2.2. Using the standard inequality
+$\left(\int_{\mathcal{O}}|v|^{q}+|\nabla v|^{q}\,\mathrm{d}x\right)^{1/q}\leqslant c\lVert{v}\rVert_{\dot{H}^{1,q}},$
 we further deduce that
 $\lVert{A_{N}v}\rVert_{L^{q}(\Omega_{\gamma};\dot{H}^{1,q})}\leqslant c(1+\lVert{v}\rVert_{\dot{H}^{1,q}}).$
-Analogously, employing the growth condition Eq. 2.2 yields
+Analogously, employing the condition Eq. 2.2 yields
 $\lVert{A_{N}v}\rVert_{L^{q}(\Omega_{\gamma};L^{q})}\leqslant c(1+\lVert{v}\rVert_{L^{q}}),\quad\forall v\in L^{q},$
-and utilizing the Lipschitz continuous condition Eq. 2.3 results in
+and utilizing the condition Eq. 2.3 results in
 $\lVert{A_{N}u-A_{N}v}\rVert_{L^{q}(\Omega_{\gamma};L^{q})}\leqslant c\lVert{u-v}\rVert_{L^{q}},\quad\forall u,v\in L^{q}.$
@@
 $\|F(v)\|_{\gamma(H,\dot{H}^{s,2})}\leqslant c(1+\|v\|_{\dot{H}^{s,2}}),\quad\text{for all $v\in\dot{H}^{s,2}$, where $s\in(0,1)$}.$
-4.2 Proof of Theorem 3.1
 It is well known that the space $L^{2}$ admits an orthonormal basis $(\phi_{n})_{n\in\mathbb{N}}$, consisting of eigenvectors of the negative Laplace operator $-\Delta$ with homogeneous Dirichlet boundary conditions.
@@
 Lemma 4.6.
-For each $n\in\mathbb{N}$ and $v_{n}\in\mathrm{span}\{\phi_{0},\ldots,\phi_{n}\}$, the stochastic Allen-Cahn equation
-$\mathrm{d}y_{n}(t)=\big{(}\Delta y_{n}+y_{n}-y_{n}^{3}\big{)}(t)\,\mathrm{d}t+F(y_{n}(t))\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T,$ (4.6a)
-$y_{n}(0)=v_{n}$ (4.6b)
+For each $n\in\mathbb{N}$ and $y_{0}^{(n)}\in\mathrm{span}\{\phi_{0},\ldots,\phi_{n}\}$, which is deterministic, the stochastic Allen-Cahn equation
+$\mathrm{d}y^{(n)}(t)=\left[\Delta y^{(n)}+y^{(n)}-(y^{(n)})^{3}\right](t)\,\mathrm{d}t+F(y^{(n)}(t))\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T,$ (4.6a)
+$y^{(n)}(0)=y_{0}^{(n)}$ (4.6b)
 has a unique mild solution. Moreover, for all $p,q\in(2,\infty)$, the following inequality holds:
-$\lVert{y_{n}}\rVert_{L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q})}\leqslant c(1+\lVert{v_{n}}\rVert_{L^{q}}).$ (4.7)
+$\lVert{y^{(n)}}\rVert_{L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q})}\leqslant c(1+\lVert{y_{0}^{(n)}}\rVert_{L^{q}}).$ (4.7)
 Proof.
 We split the proof into the following two steps.
-Step 1. Define $G_{n}(t):=e^{t\Delta}v_{n}$ for $t\in[0,T]$. Given that $v_{n}\in\dot{H}^{2,q}$ for every $q\in(2,\infty)$, it follows from Lemma 4.2 with $\alpha=\beta=2$ that
-$G_{n}\in L^{p}(0,T;\dot{H}^{2,q}),\quad\forall p,q\in(2,\infty).$ (4.8)
-In view of Eq. 4.2 and the fact that $v_{n}\in\dot{H}^{1,2}$, we use [18, Theorem 1.1] (see also, [21, Lemma 2.3 and Theorem 3.1]) to deduce that Eq. 4.6 has a unique mild solution $y_{n}$ and
-$y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{1,2}),\quad\forall p\in(2,\infty).$ (4.9)
+Step 1. Define $G^{(n)}(t):=e^{t\Delta}y_{0}^{(n)}$ for $t\in[0,T]$. Given that $y_{0}^{(n)}\in\dot{H}^{2,q}$ for every $q\in(2,\infty)$, it follows from Lemma 4.2 with $\alpha=\beta=2$ that
+$G^{(n)}\in L^{p}(0,T;\dot{H}^{2,q}),\quad\forall p,q\in(2,\infty).$ (4.8)
+In view of the estimates for $F$ in Eqs.
+4.2 and 4.4 and the fact that $y_{0}^{(n)}\in\dot{H}^{1,2}$ is deterministic, we use [20, Theorem 1.1] (see also, [23, Lemma 2.3 and Theorem 3.1]) to deduce that Eq.
+4.6 has a unique mild solution $y^{(n)}$ and
+$y^{(n)}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{1,2}),\quad\forall p\in(2,\infty).$ (4.9)
 The continuous embedding of $\dot{H}^{1,2}$ in $L^{6}$ then leads to
-$y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{6}),\quad\forall p\in(2,\infty),$
-which in turn implies that $y_{n}-y_{n}^{3}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{2})$ for all $p\in(2,\infty)$. Applying Lemma 4.3, we obtain
-$S_{0}(y_{n}-y_{n}^{3})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,2}),\quad\forall p\in(2,\infty).$ (4.10)
-Since (4.2) and (4.9) ensure that $F(y_{n})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,\dot{H}^{1,2}))$ for every $p\in(2,\infty)$, Lemma 4.4 allows us to conclude that
-$S_{1}F(y_{n})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,2}),\quad\forall p\in(2,\infty).$ (4.11)
+$y^{(n)}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{6}),\quad\forall p\in(2,\infty),$
+which in turn implies that $y^{(n)}-(y^{(n)})^{3}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{2})$ for all $p\in(2,\infty)$. Applying Lemma 4.3, we obtain
+$S_{0}(y^{(n)}-(y^{(n)})^{3})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,2}),\quad\forall p\in(2,\infty).$ (4.10)
+Since (4.2) and (4.9) ensure that $F(y^{(n)})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,\dot{H}^{1,2}))$ for every $p\in(2,\infty)$, Lemma 4.4 allows us to conclude that
+$S_{1}F(y^{(n)})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,2}),\quad\forall p\in(2,\infty).$ (4.11)
 Considering the mild solution representation
-$y_{n}=G_{n}+S_{0}(y_{n}-y_{n}^{3})+S_{1}F(y_{n}),$
+$y^{(n)}=G^{(n)}+S_{0}(y^{(n)}-(y^{(n)})^{3})+S_{1}F(y^{(n)}),$
 and combining (4.8), (4.10), and (4.11), we deduce that
-$y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,2}),\quad\forall p\in(2,\infty).$
+$y^{(n)}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,2}),\quad\forall p\in(2,\infty).$
 The continuous embeddings of $\dot{H}^{2,2}$ into $\dot{H}^{1,6}$ and $\dot{H}^{2,2}$ into $L^{q}$ for all $q\in(2,\infty)$ yield
-$y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}),\quad\forall p,q\in(2,\infty),$
-$y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{1,6}),\quad\forall p\in(2,\infty).$
+$y^{(n)}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}),\quad\forall p,q\in(2,\infty),$
+$y^{(n)}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{1,6}),\quad\forall p\in(2,\infty).$
 Hence, Eqs. 4.10 and 4.11 can be refined to
-$S_{0}(y_{n}-y_{n}^{3})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,q}),\quad\forall p,q\in(2,\infty),$ (4.12)
-$S_{1}F(y_{n})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,6}),\quad\forall p\in(2,\infty).$
+$S_{0}(y^{(n)}-(y^{(n)})^{3})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,q}),\quad\forall p,q\in(2,\infty),$ (4.12)
+$S_{1}F(y^{(n)})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,6}),\quad\forall p\in(2,\infty).$
 Combining these refined regularity results with (4.8), we arrive at
-$y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,6}),\quad\forall p\in(2,\infty),$
+$y^{(n)}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,6}),\quad\forall p\in(2,\infty),$
 which, through the continuous embedding of $\dot{H}^{2,6}$ into $\dot{H}^{1,q}$ for all $q\in(2,\infty)$, implies that
-$y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{1,q}),\quad\forall p,q\in(2,\infty).$
+$y^{(n)}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{1,q}),\quad\forall p,q\in(2,\infty).$
 This regularity result allows us to further improve (4.11) to
-$S_{1}F(y_{n})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,q}),\quad\forall p,q\in(2,\infty),$
+$S_{1}F(y^{(n)})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,q}),\quad\forall p,q\in(2,\infty),$
 which, in conjunction with (4.8) and (4.12), yields
-$y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,q}),\quad\forall p,q\in(2,\infty).$ (4.13)
-Step 2. Let $p,q\in(2,\infty)$ be arbitrary but fixed. Considering the regularity assertion Eq. 4.13 and the fact
-$y_{n}(t)=v_{n}+\int_{0}^{t}(\Delta y_{n}+y_{n}-y_{n}^{3})(s)\,\mathrm{d}s+\int_{0}^{t}F(y_{n}(s))\,\mathrm{d}W_{H}(s),\quad t\in[0,T],$
-we use Itô’s formula in UMD Banach spaces (see [7, Theorem 2.4]) to conclude that, for any $t\in(0,T]$,
-$\mathbb{E}\Big{[}\lVert{y_{n}(t)}\rVert_{L^{q}}^{p}-\lVert{v_{n}}\rVert_{L^{q}}^{p}\Big{]}$
+$y^{(n)}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,q}),\quad\forall p,q\in(2,\infty).$ (4.13)
+Step 2. Let $p,q\in(2,\infty)$ be arbitrary but fixed. Considering the regularity assertion Eq. 4.13 and the fact, $\mathbb{P}$-almost surely,
+$y^{(n)}(t)=y_{0}^{(n)}+\int_{0}^{t}\Delta y^{(n)}(s)+y^{(n)}(s)-(y^{(n)}(s))^{3}\,\mathrm{d}s+\int_{0}^{t}F(y^{(n)}(s))\,\mathrm{d}W_{H}(s),\quad t\in[0,T],$
+we use Itô’s formula in UMD Banach spaces (see [8, Theorem 2.4]) to conclude that, for any $t\in(0,T]$,
+$\mathbb{E}\Big{[}\lVert{y^{(n)}(t)}\rVert_{L^{q}}^{p}-\lVert{y_{0}^{(n)}}\rVert_{L^{q}}^{p}\Big{]}$
 […]
-$\quad{}+\frac{p(q-1)}{2}\,\mathbb{E}\Biggl{[}\int_{0}^{t}\lVert{y_{n}(s)}\rVert_{L^{q}}^{p-q}\,\int_{\mathcal{O}}\lvert{y_{n}(s)}\rvert^{q-2}\sum_{k=0}^{\infty}\lambda_{k}\lvert{f_{k}(y_{n}(s))}\rvert^{2}\,\mathrm{d}\mu\,\mathrm{d}s\Biggr{]}$
-$\quad{}+\frac{p(p-q)}{2}\,\mathbb{E}\Biggl{\{}\int_{0}^{t}\lVert{y_{n}(s)}\rVert_{L^{q}}^{p-2q}\sum_{k=0}^{\infty}\lambda_{k}\bigg{[}\int_{\mathcal{O}}\lvert{y_{n}(s)}\rvert^{q-2}y_{n}(s)f_{k}(y_{n}(s))\,\mathrm{d}\mu\bigg{]}^{2}\,\mathrm{d}s\Biggr{\}}$
+$\quad{}+\frac{p(q-1)}{2}\,\mathbb{E}\Biggl{[}\int_{0}^{t}\lVert{y^{(n)}(s)}\rVert_{L^{q}}^{p-q}\,\int_{\mathcal{O}}\lvert{y^{(n)}(s)}\rvert^{q-2}\sum_{k=0}^{\infty}\lvert{f_{k}(y^{(n)}(s))}\rvert^{2}\,\mathrm{d}x\,\mathrm{d}s\Biggr{]}$
+$\quad{}+\frac{p(p-q)}{2}\,\mathbb{E}\Biggl{\{}\int_{0}^{t}\lVert{y^{(n)}(s)}\rVert_{L^{q}}^{p-2q}\sum_{k=0}^{\infty}\bigg{[}\int_{\mathcal{O}}\lvert{y^{(n)}(s)}\rvert^{q-2}y^{(n)}(s)f_{k}(y^{(n)}(s))\,\mathrm{d}x\bigg{]}^{2}\,\mathrm{d}s\Biggr{\}}$
 […]
 Considering the inequality
-$\int_{\mathcal{O}}\lvert{y_{n}(s)}\rvert^{q-2}y_{n}(s)(\Delta y_{n}-y_{n}^{3})(s)\,\mathrm{d}\mu$
+$\int_{\mathcal{O}}\lvert{y^{(n)}(s)}\rvert^{q-2}y^{(n)}(s)[\Delta y^{(n)}(s)-(y^{(n)}(s))^{3}]\,\mathrm{d}x$
 […]
@@
 we infer the bound for $I_{1}$:
-$I_{1}\leqslant p\,\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{y_{n}(s)}\rVert_{L^{q}}^{p}\,\mathrm{d}s\bigg{]}.$
-For $I_{2}+I_{3}$, using Hölder’s inequality in conjunction with Young’s inequality, and employing the summability condition Eq.
-2.5 along with the growth condition Eq.
-2.2, we derive the following bound:
-$I_{2}+I_{3}\leqslant c\,\mathbb{E}\bigg{[}\int_{0}^{t}1+\lVert{y_{n}(s)}\rVert_{L^{q}}^{p}\,\mathrm{d}s\bigg{]}.$
+$I_{1}\leqslant p\,\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{y^{(n)}(s)}\rVert_{L^{q}}^{p}\,\mathrm{d}s\bigg{]}.$
+For $I_{2}+I_{3}$, using Hölder’s inequality in conjunction with Young’s inequality, and employing condition Eq. 2.2, we derive the following bound:
+$I_{2}+I_{3}\leqslant c\,\mathbb{E}\bigg{[}\int_{0}^{t}1+\lVert{y^{(n)}(s)}\rVert_{L^{q}}^{p}\,\mathrm{d}s\bigg{]}.$
 Combining the above bounds for $I_{1}$ and $I_{2}+I_{3}$, we arrive at
-$\mathbb{E}\lVert{y_{n}(t)}\rVert_{L^{q}}^{p}\leqslant\lVert{v_{n}}\rVert_{L^{q}}^{p}+c\,\mathbb{E}\bigg{[}\int_{0}^{t}1+\lVert{y_{n}(s)}\rVert_{L^{q}}^{p}\,\mathrm{d}s\bigg{]},\quad\forall t\in[0,T].$
-Applying Gronwall’s lemma to the above inequality yields
-$\lVert{y_{n}}\rVert_{C([0,T];L^{p}(\Omega;L^{q}))}\leqslant c\big{(}1+\lVert{v_{n}}\rVert_{L^{q}}\big{)},$
-which implies the desired inequality Eq. 4.7. This completes the proof. ∎
-Proof of Theorem 3.1. We split the proof into the following two steps.
+$\mathbb{E}\lVert{y^{(n)}(t)}\rVert_{L^{q}}^{p}\leqslant\lVert{y_{0}^{(n)}}\rVert_{L^{q}}^{p}+c\,\mathbb{E}\bigg{[}\int_{0}^{t}1+\lVert{y^{(n)}(s)}\rVert_{L^{q}}^{p}\,\mathrm{d}s\bigg{]},\quad\forall t\in[0,T].$
+An application of Gronwall’s inequality then yields
+$\lVert{y^{(n)}}\rVert_{C([0,T];L^{p}(\Omega;L^{q}))}\leqslant c\big{(}1+\lVert{y_{0}^{(n)}}\rVert_{L^{q}}\big{)},$
+which implies the desired estimate Eq. 4.7. This completes the proof. ∎
+4.2 Proof of Theorem 3.1
+We split the proof into the following three steps.
 Step 1.
@@
 3.1 is satisfied.
-Let $p,q\in(2,\infty)$ be fixed but arbitrary.
-For each $n\in\mathbb{N}$, denote by $y_{n}$ the mild solution of equation Eq.
-4.6, where $v_{n}$ is defined as
-$v_{n}:=\mathop{\mathrm{argmin}}_{w\in\mathrm{span}\{\phi_{0},\ldots,\phi_{n}\}}\lVert{v-w}\rVert_{L^{3q}}.$
-For any $m,n\in\mathbb{N}$, put $e_{m,n}:=y_{m}-y_{n}$. By definition, we infer that
-$\mathrm{d}e_{m,n}(t)=(\Delta e_{m,n}+e_{m,n}+y_{n}^{3}-y_{m}^{3})(t)\,\mathrm{d}t+\big{(}F(y_{m}(t))-F(y_{n}(t))\big{)}\,\mathrm{d}W_{H}(t),\quad t\in[0,T].$
-Observing that the integral over $\mathcal{O}$ of $\lvert{e_{m,n}(t)}\rvert^{q-2}e_{m,n}(t)(y_{n}^{3}-y_{m}^{3})(t)$ is non-positive for all $t\in[0,T]$, and following the proof of Eq.
-4.7, we can deduce that
-$\lVert{e_{m,n}}\rVert_{L_{\mathbb{F}}^{3p}(\Omega\times(0,T);L^{3q})}\leqslant c\lVert{v_{m}-v_{n}}\rVert_{L^{3q}}.$ (4.14)
-On the other hand, from the fact that $\dot{H}^{2,2}$ is dense in $L^{3q}$ and that $(\phi_{n})_{n\in\mathbb{N}}$ is a complete orthogonal system in $\dot{H}^{2,2}$, we conclude that $(v_{n})_{n\in\mathbb{N}}$ is a Cauchy sequence in $L^{3q}$, converging to $v$.
+Let $p\in(2,\infty)$ and $q\in[4,\infty)$ be arbitrarily chosen.
+For each $n\in\mathbb{N}$, denote by $y^{(n)}$ the mild solution of equation Eq.
+4.6, where $y_{0}^{(n)}$ is defined as
+$y_{0}^{(n)}:=\mathop{\mathrm{argmin}}_{w\in\mathrm{span}\{\phi_{0},\ldots,\phi_{n}\}}\lVert{y_{0}-w}\rVert_{L^{3q}}.$
+For any $m,n\in\mathbb{N}$, put $e_{m,n}:=y^{(m)}-y^{(n)}$. By definition, we infer that
+$\mathrm{d}e_{m,n}(t)=(\Delta e_{m,n}+e_{m,n}+(y^{(n)})^{3}-(y^{(m)})^{3})(t)\,\mathrm{d}t+\big{(}F(y^{(m)}(t))-F(y^{(n)}(t))\big{)}\,\mathrm{d}W_{H}(t),\quad t\in[0,T].$
+Observing that the integral over $\mathcal{O}$ of
+$\lvert{e_{m,n}(t)}\rvert^{3q-2}e_{m,n}(t)[\Delta e_{m,n}(t)+(y^{(n)}(t))^{3}-(y^{(m)}(t))^{3}]$
+is non-positive for all $t\in[0,T]$, and following the proof of Eq. 4.7, we can deduce that
+$\lVert{e_{m,n}}\rVert_{L_{\mathbb{F}}^{3p}(\Omega\times(0,T);L^{3q})}\leqslant c\lVert{y_{0}^{(m)}-y_{0}^{(n)}}\rVert_{L^{3q}}.$ (4.14)
+On the other hand, from the fact that $\dot{H}^{2,2}$ is dense in $L^{3q}$ and that $(\phi_{n})_{n\in\mathbb{N}}$ is a complete orthogonal system in $\dot{H}^{2,2}$, we conclude that $(y_{0}^{(n)})_{n\in\mathbb{N}}$ is a Cauchy sequence in $L^{3q}$ and converges to $y_{0}$ in $L^{3q}$, i.e.,
+$\lim_{n\to\infty}\lVert{y_{0}^{(n)}-y_{0}}\rVert_{L^{3q}}=0.$ (4.15)
 Consequently, there exists a unique process $y\in L_{\mathbb{F}}^{3p}(\Omega\times(0,T);L^{3q})$ such that
-$\lim_{n\to\infty}y_{n}=y\quad\text{ in }L_{\mathbb{F}}^{3p}(\Omega\times(0,T);L^{3q}).$ (4.15)
+$\lim_{n\to\infty}y^{(n)}=y\quad\text{ in }L_{\mathbb{F}}^{3p}(\Omega\times(0,T);L^{3q}).$ (4.16)
 For any $n\in\mathbb{N}$, applying Hölder’s inequality yields
-$\lVert{y_{n}^{3}-y^{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}$
+$\lVert{(y^{(n)})^{3}-y^{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}$
 […]
 Therefore, using the convergence established in Eq. 4.16, we deduce that
-$\lim_{n\to\infty}y_{n}^{3}=y^{3}\quad\text{ in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}),$
+$\lim_{n\to\infty}(y^{(n)})^{3}=y^{3}\quad\text{ in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}),$
 which consequently leads to
-$\lim_{n\to\infty}y_{n}-y_{n}^{3}=y-y^{3}\quad\text{ in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}).$
-Using Lemma 4.3, this directly implies the following convergence of $S_{0}(y_{n}-y_{n}^{3})$:
-$\lim_{n\to\infty}S_{0}(y_{n}-y_{n}^{3})=S_{0}(y-y^{3})\quad\text{ in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}).$
-Furthermore, invoking Eqs. 4.4 and 4.15 yields
-$\lim_{n\to\infty}F(y_{n})=F(y)\quad\text{ in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,L^{q})),$
-and subsequently by Lemma 4.4 we infer the following convergence for $S_{1}F(y_{n})$:
-$\lim_{n\to\infty}S_{1}F(y_{n})=S_{1}F(y)\quad\text{ in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}).$
-Additionally, applying Lemma 4.2 with $\alpha=\beta=0$ and noting the convergence $\lim_{n\in\infty}\lVert{v_{n}-v}\rVert_{L^{q}}=0$, we establish the following convergence for $G_{n}$:
-$\lim_{n\to\infty}G_{n}=G\quad\text{ in }L^{p}(0,T;L^{q}),$
+$\lim_{n\to\infty}y^{(n)}-(y^{(n)})^{3}=y-y^{3}\quad\text{ in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}).$
+Using Lemma 4.3, this directly implies the following convergence of $S_{0}(y^{(n)}-(y^{(n)})^{3})$:
+$\lim_{n\to\infty}S_{0}(y^{(n)}-(y^{(n)})^{3})=S_{0}(y-y^{3})\quad\text{ in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q})\cap L_{\mathbb{F}}^{p}(\Omega;C([0,T];L^{q})).$
+Furthermore, invoking Eqs. 4.4 and 4.16 yields
+$\lim_{n\to\infty}F(y^{(n)})=F(y)\quad\text{ in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,L^{q})),$
+and subsequently by Lemma 4.4 we infer the following convergence for $S_{1}F(y^{(n)})$:
+$\lim_{n\to\infty}S_{1}F(y^{(n)})=S_{1}F(y)\quad\text{ in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q})\cap L_{\mathbb{F}}^{p}(\Omega;C([0,T];L^{q})).$
+Additionally, applying Lemma 4.2 with $\alpha=\beta=0$ and using Eq. 4.15, we establish the following convergence for $G^{(n)}$:
+$\lim_{n\to\infty}G^{(n)}=G\quad\text{ in }C([0,T];L^{q}),$
 where $G$ is defined by Eq.
-2.10, and $G_{n}(t):=e^{t\Delta}v_{n}$ for each $t\in[0,T]$.
-Finally, combining the derived convergences of $y_{n}$, $G_{n}$, $S_{0}(y_{n}-y_{n}^{3})$, and $S_{1}F(y_{n})$, and recalling that $y_{n}=G_{n}+S_{0}(y_{n}-y_{n}^{3})+S_{1}F(y_{n})$, we obtain by passing to the limit $n\to\infty$ that
-$y=G+S_{0}(y-y^{3})+S_{1}F(y)\quad\text{in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}).$
-This confirms that $y$ is a mild solution of Eq.
-1.1, and it is evident that Eq.
-3.1 holds.
-To prove the uniqueness of the mild solution, assume $\widetilde{y}$ is another mild solution.
-Consider the difference $e=y-\widetilde{y}$.
-Then $e$ satisfies the equation
-$\begin{cases}\mathrm{d}e(t)=(\Delta e+e+\widetilde{y}^{3}-y^{3})(t)\,\mathrm{d}t+\big{(}F(y(t))-F(\widetilde{y}(t))\big{)}\,\mathrm{d}W_{H}(t),&t\in[0,T],\\ e(0)=0.\end{cases}$
-By an argument similar to that used in Eq. 4.14, we find that
-$\lVert{e}\rVert_{L^{2}(\Omega\times(0,T);L^{2})}\leqslant 0,$
-which implies that $e=0$ almost surely, confirming the uniqueness of the mild solution to Eq. 1.1.
+2.8, and $G^{(n)}(t):=e^{t\Delta}y_{0}^{(n)}$ for each $t\in[0,T]$.
+Finally, combining the derived convergences of $y^{(n)}$, $G^{(n)}$, $S_{0}(y^{(n)}-(y^{(n)})^{3})$, and $S_{1}F(y^{(n)})$, and recalling that $y^{(n)}=G^{(n)}+S_{0}(y^{(n)}-(y^{(n)})^{3})+S_{1}F(y^{(n)})$, we obtain by passing to the limit $n\to\infty$ that
+$y=G+S_{0}(y-y^{3})+S_{1}F(y)\quad\text{in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q})\cap L_{\mathbb{F}}^{p}(\Omega;C([0,T];L^{q})).$
+This confirms that $y$ is a mild solution of Eq. 1.1, and it is evident that Eq. 3.1 holds.
 Step 2.
-Fix any $p\in(2,\infty)$ and $q\in[2,\infty)$.
-By virtue of the regularity assertion in Eq.
+We now consider the uniqueness of the mild solution $y$, which can be proved by a standard localization argument.
+Suppose $\widetilde{y}\in L_{\mathbb{F}}^{0}(\Omega;C([0,T];L^{4}))$ is another mild solution of Eq.
+1.1, and define the difference $e:=y-\widetilde{y}$.
+Then $e$ satisfies the stochastic evolution equation
+$\begin{cases}\mathrm{d}e(t)=\left(\Delta e(t)+e(t)+\widetilde{y}^{3}(t)-y^{3}(t)\right)\mathrm{d}t\\ \qquad\qquad{}+\left(F(y(t))-F(\widetilde{y}(t))\right)\mathrm{d}W_{H}(t)\quad\text{in $\dot{H}^{-1,2}$,}\quad t\in[0,T],\\ e(0)=0.\end{cases}$
+For $m\geqslant 1$, define the stopping time
+$\sigma_{m}:=\inf\left\{t\in[0,T]\colon\max\left(\|y(t)\|_{L^{4}},\|\widetilde{y}(t)\|_{L^{4}}\right)\geqslant m\right\},$
+with $\sigma_{m}=T$ if the set is empty. Set $e_{m}(t):=e(t\wedge\sigma_{m})$ for $t\in[0,T]$. Then $e_{m}$ satisfies the stopped equation
+$\begin{cases}\mathrm{d}e_{m}(t)=\mathbbm{1}_{[0,\sigma_{m}]}(t)\left(\Delta e_{m}(t)+e_{m}(t)+\widetilde{y}^{3}(t\wedge\sigma_{m})-y^{3}(t\wedge\sigma_{m})\right)\mathrm{d}t\\ \qquad\qquad{}+\mathbbm{1}_{[0,\sigma_{m}]}(t)\left(F(y(t))-F(\widetilde{y}(t))\right)\mathrm{d}W_{H}(t)\quad\text{in $\dot{H}^{-1,2}$},\quad t\in[0,T],\\ e_{m}(0)=0.\end{cases}$
+Applying Itô’s formula to $\|e_{m}(t)\|_{L^{2}}^{2}$ and taking expectation yields, for all $t\in[0,T]$,
+$\mathbb{E}\|e_{m}(t)\|_{L^{2}}^{2}$ $=2\mathbb{E}\bigg{[}\mathbbm{1}_{[0,\sigma_{m}]}(s)\big{\langle}\Delta e_{m}(s)+e_{m}(s)+\widetilde{y}^{3}(s)-y^{3}(s),\,e_{m}(s)\big{\rangle}\,\mathrm{d}s\bigg{]}$
+$\quad+\mathbb{E}\bigg{[}\int_{0}^{t}\mathbbm{1}_{[0,\sigma_{m}]}(s)\|F(y(s))-F(\widetilde{y}(s))\|_{\gamma(H,L^{2})}^{2}\,\mathrm{d}s\bigg{]},$
+[…]
+$\mathbbm{1}_{[0,\sigma_{m}]}(s)\big{\langle}\Delta e_{m}(s)+e_{m}(s)+\widetilde{y}^{3}(s)-y^{3}(s),\,e_{m}(s)\big{\rangle}\leqslant\mathbbm{1}_{[0,\sigma_{m}]}(s)\|e_{m}(s)\|_{L^{2}}^{2},$
+it follows that
+$\mathbb{E}\|e_{m}(t)\|_{L^{2}}^{2}\leqslant\mathbb{E}\bigg{[}\int_{0}^{t}\mathbbm{1}_{[0,\sigma_{m}]}(s)\left(\|e_{m}(s)\|_{L^{2}}^{2}+\|F(y(s))-F(\widetilde{y}(s))\|_{\gamma(H,L^{2})}^{2}\right)\mathrm{d}s\bigg{]}.$
+Using the Lipschitz condition Eq. 4.4 of $F$, we deduce
+$\mathbb{E}\|e_{m}(t)\|_{L^{2}}^{2}\leqslant c\mathbb{E}\int_{0}^{t}\|e_{m}(s)\|_{L^{2}}^{2}\,\mathrm{d}s,\quad\forall t\in[0,T].$
+Gronwall’s inequality then yields
+$\sup_{t\in[0,T]}\mathbb{E}\|e_{m}(t)\|_{L^{2}}^{2}=0,\quad\forall m\geqslant 1.$
+Since $e_{m}\in L_{\mathbb{F}}^{0}(\Omega;C([0,T];L^{4}))$, this implies $\|e_{m}\|_{C([0,T];L^{4})}=0$ holds $\mathbb{P}$-almost surely. Consequently,
+$y=\widetilde{y}\quad\text{in }C([0,\sigma_{m}];L^{4})\text{ holds }\mathbb{P}\text{-almost surely, }\forall m\geqslant 1.$
+As $y,\widetilde{y}\in L_{\mathbb{F}}^{0}(\Omega;C([0,T];L^{4}))$, we have $\lim_{m\to\infty}\mathbb{P}(\sigma_{m}=T)=1$.
+Thus, $y=\widetilde{y}$ in $C([0,T];L^{4})$ holds $\mathbb{P}$-almost surely, proving uniqueness for the mild solution to Eq.
+1.1.
+Step 3.
+Fix any $p,q\in(2,\infty)$.
+By the regularity assertion in Eq.
 3.1 and the growth estimate Eq.
@@
 while Lemma 4.4 implies $S_{1}F(y)\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{1,q})$.
-Additionally, Lemma 4.2 ensures that $G\in L^{p}(0,T;\dot{H}^{2/p-\epsilon,q})$ for all $\epsilon>0$.
+Additionally, the initial condition $y_{0}\in L^{\infty}$ ensures by Lemma 4.2 that $G\in L^{p}(0,T;\dot{H}^{2/p-\epsilon,q})$ for all $\epsilon>0$.
 Consequently, by the mild solution representation Eq.
-2.9, we arrive at
+2.7, we arrive at
 $y\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2/p-\epsilon,q}),\quad\forall\epsilon>0,$
@@
 Finally, combining Eqs.
-2.9, 4.16 and 4.17 establishes the regularity assertions in Eqs.
+2.7, 4.17 and 4.18 establishes the regularity assertions in Eqs.
 3.4 and 3.5.
@@
 3.2 follows directly from Eq.
-3.4 and the observation that $G\in C([0,T];L^{q})$.
+3.4 and the observation that $G\in C([0,T];L^{q})$, which is justified by the initial condition $y_{0}\in L^{\infty}$ and Lemma 4.2.
 This completes the proof of Theorem 3.1.
@@
 4.3 Proof of Theorems 3.2 and 3.3
-We begin by summarizing several well-established results. As known from [8], it holds that
-$\|P_{h}\|_{\mathcal{L}(L^{q})}\text{ is uniformly bounded with respect to }h\text{ for all }q\in(1,\infty).$ (4.18)
-Consequently, the operator $P_{h}$ possesses the following approximation property:
+We begin by summarizing several well-known results. It is known from [9] that
+$\|P_{h}\|_{\mathcal{L}(L^{q})}\text{ is uniformly bounded in }h\text{ for all }q\in[1,\infty].$ (4.19)
+This uniform stability of $P_{h}$ implies the following approximation property:
 $\|I-P_{h}\|_{\mathcal{L}(\dot{H}^{\alpha,q},L^{q})}\leqslant ch^{\alpha}\quad\text{for all }\alpha\in[0,2]\text{ and }q\in(1,\infty).$ (4.20)
-Inequality (2.5) from [1], combined with the observation that
-$\lVert{\Delta_{h}^{-1}P_{h}\Delta}\rVert_{\mathcal{L}(\dot{H}^{1,2})}\leqslant 1,$
-yields, through interpolation, that for any $q\in[2,\infty)$, the operator norm
-$\lVert{\Delta_{h}^{-1}P_{h}\Delta}\rVert_{\mathcal{L}(\dot{H}^{1,q})}$
-is uniformly bounded with respect to $h$.
-Utilizing this fundamental stability estimate, the approximation property Eq.
-4.19, the interpolation theory as stated in [22, Theorem 2.6], and the fact that $\Delta^{-1}$ is a bounded linear operator mapping from $\dot{H}^{-s,q}$ to $\dot{H}^{2-s,q}$ for all $s\in[0,2]$, standard arguments then lead to the conclusion that
-$\|P_{h}\Delta^{-1}-\Delta_{h}^{-1}P_{h}\|_{\mathcal{L}(\dot{H}^{\beta-2,q},\dot{H}^{\alpha,q})}\leqslant ch^{\beta-\alpha}\quad\text{for all }0\leqslant\alpha\leqslant 1\leqslant\beta\leqslant 2.$ (4.20)
+Furthermore, it is a standard result (see, e.g., [13, Equations (12) and (13)]) that
+$\lVert{I-\Delta_{h}^{-1}P_{h}\Delta}\rVert_{\mathcal{L}(\dot{H}^{2,q},L^{q})}$ $\leqslant ch^{2},\quad\forall\,q\in(1,\infty),$ (4.21)
+$\lVert{I-\Delta_{h}^{-1}P_{h}\Delta}\rVert_{\mathcal{L}(\dot{H}^{1,q},L^{q})}$ $\leqslant ch,\quad\forall\,q\in[2,\infty).$ (4.22)
+By applying complex interpolation between (4.21) and (4.22), one obtains
+$\lVert{I-\Delta_{h}^{-1}P_{h}\Delta}\rVert_{\mathcal{L}(\dot{H}^{\alpha,q},L^{q})}\leqslant ch^{\alpha},\quad\forall\,\alpha\in[1,2),\quad\forall\,q\in[2,\infty).$ (4.23)
+Combining this with Eq. 4.20 via the triangle inequality yields
+$\lVert{P_{h}-\Delta_{h}^{-1}P_{h}\Delta}\rVert_{\mathcal{L}(\dot{H}^{\alpha,q},L^{q})}\leqslant ch^{\alpha},\quad\forall\alpha\in[1,2],\,\forall q\in[2,\infty),$ (4.24)
+which in turn implies the estimate
+$\lVert{P_{h}\Delta^{-1}-\Delta_{h}^{-1}P_{h}}\rVert_{\mathcal{L}(\dot{H}^{\alpha-2,q},L^{q})}\leqslant ch^{\alpha},\quad\forall\alpha\in[1,2],\,\forall q\in[2,\infty).$ (4.25)
+Remark 4.2.
+The projection operator $P_{h}$ can be extended to act as a linear operator from $\dot{H}^{-1,q}$ to $X_{h}$ for any $q\in(1,\infty)$. This extension is defined via the identity
+$\int_{\mathcal{O}}(P_{h}v)v_{h}\,\mathrm{d}x=\langle{v,v_{h}}\rangle,$
+[…]
+Next, we introduce the discrete Sobolev space and recall some standard results.
+Let $q\in(1,\infty)$.
+For any $\alpha\in\mathbb{R}$, define $\dot{H}_{h}^{\alpha,q}$ as the space $X_{h}$ equipped with the norm
+$\|v_{h}\|_{\dot{H}_{h}^{\alpha,q}}:=\|(-\Delta_{h})^{\alpha/2}v_{h}\|_{L^{q}},\quad v_{h}\in X_{h}.$
+According to Theorem 3.1 (ii) in [19], the purely imaginary powers of $-\Delta_{h}$ are bounded in $\mathcal{L}(\dot{H}_{h}^{0,q})$ uniformly in $h$.
+As a consequence, an application of Theorem 4.17 from [24] yields, for any $\alpha\in(0,2)$, the equivalence
+$c_{0}\|v_{h}\|_{\dot{H}_{h}^{\alpha,q}}\leqslant\|v_{h}\|_{[\dot{H}_{h}^{0,q},\dot{H}_{h}^{2,q}]_{\alpha/2}}\leqslant c_{1}\|v_{h}\|_{\dot{H}_{h}^{\alpha,q}},\quad v_{h}\in X_{h},$
+with constants $c_{0},c_{1}>0$ independent of $h$.
+Here, $[\dot{H}_{h}^{0,q},\dot{H}_{h}^{2,q}]_{\alpha/2}$ denotes the complex interpolation space between $\dot{H}_{h}^{0,q}$ and $\dot{H}_{h}^{2,q}$; see [24, Chapter 2] for details.
+Given the standard inverse estimate
+$\|v_{h}\|_{\dot{H}_{h}^{2,q}}\leqslant ch^{-2}\|v_{h}\|_{L^{q}},\quad v_{h}\in X_{h},$
+complex interpolation yields the following general inverse estimate:
+$\|v_{h}\|_{\dot{H}_{h}^{\beta,q}}\leqslant ch^{\alpha-\beta}\|v_{h}\|_{\dot{H}_{h}^{\alpha,q}},\quad v_{h}\in X_{h},\,\forall 0\leqslant\alpha<\beta<\infty.$ (4.26)
+Applying this inverse estimate to Eq. 4.25, we obtain the estimate
+$\|P_{h}\Delta^{-1}-\Delta_{h}^{-1}P_{h}\|_{\mathcal{L}(\dot{H}^{\alpha-2,q},\dot{H}_{h}^{\beta,q})}\leqslant ch^{\alpha-\beta}\quad\text{for all }\alpha\in[1,2],\,\beta\in[0,1],\,q\in[2,\infty).$ (4.27)
+In analogy with Lemma 4.1, we also have the following $h$-uniform embedding result.
+Lemma 4.7.
+Let $0\leqslant\beta<\gamma\leqslant 1$ and $p,q\in(1,\infty)$. Then
+$\|v_{h}\|_{\dot{H}_{h}^{\beta,q}}\leqslant c\|v_{h}\|_{(\dot{H}_{h}^{0,q},\dot{H}_{h}^{2,q})_{\gamma/2,p}},\quad v_{h}\in X_{h}.$
+Additionally, we establish the following $h$-uniform Sobolev-type embeddings into $L^{\infty}$, whose proofs are given in Appendix A.
+Lemma 4.8.
+The following continuous embeddings hold:
+  1. (i)
+    If $q\in(3/2,\infty)$, then $\|v_{h}\|_{L^{\infty}}\leqslant c\|v_{h}\|_{\dot{H}_{h}^{2,q}}$ for all $v_{h}\in X_{h}$.
+  2. (ii)
+    If $\alpha\in(0,2)$ and $q\in(3/\alpha,\infty)\cap[2,\infty)$, then $\|v_{h}\|_{L^{\infty}}\leqslant c\|v_{h}\|_{\dot{H}_{h}^{\alpha,q}}$ for all $v_{h}\in X_{h}$.
+Now, we introduce the discrete semigroup $e^{t\Delta_{h}}$ and the maximal $L^{p}$-regularity estimates associated with this semigroup.
 Let $e^{t\Delta_{h}}$, for $t\in[0,\infty)$, denote the analytic semigroup generated by the discrete Laplace operator $\Delta_{h}$.
-Using the established estimates for the resolvents of $\Delta_{h}$ from [1] and applying (2.128) from [38], we obtain the well-known growth estimate: for any $\alpha\in[0,1]$, $q\in(1,\infty)$, and $t\in(0,T]$,
+The following fundamental properties hold, with proofs provided in Appendix A.
+Lemma 4.9.
+The following estimates hold:
+  • (i)
+    For any $\alpha\in[0,1]$, $q\in(1,\infty)$, and $t>0$,
     $\lVert{(-\Delta_{h})^{\alpha}e^{t\Delta_{h}}}\rVert_{\mathcal{L}(L^{q})}\leqslant ct^{-\alpha}.$ (4.28)
-Furthermore, utilizing techniques analogous to those found in Theorem 6.10 and Lemma 20.7 of [35], one can readily derive the following well-known inequality: for any $\alpha\in[0,2]$, $q\in(1,\infty)$, and $t\in[0,T]$,
+  • (ii)
+    For any $\alpha\in[0,2]$, $q\in(1,\infty)$, and $t>0$,
     $\lVert{e^{t\Delta_{h}}P_{h}-P_{h}e^{t\Delta}}\rVert_{\mathcal{L}(\dot{H}^{\alpha,q},L^{q})}+\lVert{(I-P_{h})e^{t\Delta}}\rVert_{\mathcal{L}(\dot{H}^{\alpha,q},L^{q})}\leqslant c\min\Big{\{}h^{\alpha},\frac{h^{2}}{t^{1-\alpha/2}}\Big{\}}.$ (4.29)
-For the special case where $q=2$, the reader is referred to Theorem 3.2 of [35] for a detailed treatment of this result.
-Let $q\in(1,\infty)$. For any $\alpha\in\mathbb{R}$, denote by $\dot{H}_{h}^{\alpha,q}$ the space $X_{h}$ equipped with the norm
-$\|v_{h}\|_{\dot{H}_{h}^{\alpha,q}}:=\|(-\Delta_{h})^{\alpha/2}v_{h}\|_{L^{q}}\quad\text{for all }v_{h}\in X_{h}.$
-According to Theorem 3.1 (ii) in [17], the boundedness of the purely imaginary powers of $-\Delta_{h}$ in $\mathcal{L}(\dot{H}_{h}^{0,q})$ is independent of the spatial mesh size $h$.
-Consequently, by applying Theorem 4.17 from [22], we deduce that for any $\alpha\in(0,2)$,
-$c_{0}\|v_{h}\|_{\dot{H}_{h}^{\alpha,q}}\leqslant\|v_{h}\|_{[\dot{H}_{h}^{0,q},\dot{H}_{h}^{2,q}]_{\alpha/2}}\leqslant c_{1}\|v_{h}\|_{\dot{H}_{h}^{\alpha,q}}\quad\text{for all }v_{h}\in X_{h},$ (4.23)
-where $c_{0}$ and $c_{1}$ are positive constants independent of $h$.
-Here, $[\dot{H}_{h}^{0,q},\dot{H}_{h}^{2,q}]_{\alpha/2}$ denotes the complex interpolation space between $\dot{H}_{h}^{0,q}$ and $\dot{H}_{h}^{2,q}$ (see [22, Chapter 2]).
-This result is important as it enables the application of interpolation theory to the space $\dot{H}_{h}^{s,q}$ for all $s\in\mathbb{R}$.
-For example, given the standard inverse estimate
-$\|v_{h}\|_{\dot{H}_{h}^{2,q}}\leqslant ch^{-2}\|v_{h}\|_{L^{q}}\quad\text{for all }v_{h}\in X_{h},$
-we can invoke [22, Theorem 2.6] to derive the following inverse estimate: for any $0\leqslant\alpha<\beta<\infty$,
-$\|v_{h}\|_{\dot{H}_{h}^{\beta,q}}\leqslant ch^{\alpha-\beta}\|v_{h}\|_{\dot{H}_{h}^{\alpha,q}}\quad\text{for all }v_{h}\in X_{h}.$
-Furthermore, for any $\alpha\in[0,1]$, there exist positive constants $c_{0}$ and $c_{1}$, independent of $h$, such that
-$c_{0}\|v_{h}\|_{\dot{H}^{\alpha,q}}\leqslant\|v_{h}\|_{\dot{H}_{h}^{\alpha,q}}\leqslant c_{1}\|v_{h}\|_{\dot{H}^{\alpha,q}}\quad\text{for all }v_{h}\in X_{h}.$
-Henceforth, the aforementioned theoretical results will be utilized without further explicit reference.
-We now introduce a discrete deterministic maximal $L^{p}$-regularity estimate and its stochastic counterpart.
-For a function $g_{h}\in L^{1}(0,T;\dot{H}_{h}^{0,2})$, the discrete deterministic convolution operator $S_{0,h}$ is defined as
-$(S_{0,h}g_{h})(t):=\int_{0}^{t}e^{(t-s)\Delta_{h}}g_{h}(s)\,\mathrm{d}s,\quad\text{for }t\in[0,T].$ (4.24)
-For a process $g_{h}\in L_{\mathbb{F}}^{2}(\Omega\times(0,T);\gamma(H,\dot{H}_{h}^{0,2}))$, the discrete stochastic convolution operator $S_{1,h}$ is given by
-$(S_{1,h}g_{h})(t):=\int_{0}^{t}e^{(t-s)\Delta_{h}}g_{h}(s)\,\mathrm{d}W_{H}(s),\quad\text{for }t\in[0,T].$ (4.25)
-The proofs of the following two lemmas can be found in [16] for the first and in [17] for the second.
+For a function $g_{h}\in L^{1}(0,T;\dot{H}_{h}^{0,2})$, the discrete convolution operator $S_{0,h}$ is defined by
+$(S_{0,h}g_{h})(t):=\int_{0}^{t}e^{(t-s)\Delta_{h}}g_{h}(s)\,\mathrm{d}s,\quad t\in[0,T].$ (4.30)
+For a stochastic process $g_{h}\in L_{\mathbb{F}}^{2}(\Omega\times(0,T);\gamma(H,\dot{H}_{h}^{0,2}))$, the discrete stochastic convolution operator $S_{1,h}$ is given by
+$(S_{1,h}g_{h})(t):=\int_{0}^{t}e^{(t-s)\Delta_{h}}g_{h}(s)\,\mathrm{d}W_{H}(s),\quad t\in[0,T].$ (4.31)
+These operators satisfy the following fundamental discrete maximal $L^{p}$-regularity estimates.
 Lemma 4.10.
-Suppose $p,q\in(1,\infty)$. For any $g_{h}\in L^{p}(0,T;X_{h})$, the following discrete deterministic maximal $L^{p}$-regularity estimate is valid:
-$\Big{\lVert}{\frac{\mathrm{d}}{\mathrm{d}t}S_{0,h}g_{h}}\Big{\rVert}_{L^{p}(0,T;\dot{H}_{h}^{0,q})}+\lVert{S_{0,h}g_{h}}\rVert_{L^{p}(0,T;\dot{H}_{h}^{2,q})}\leqslant c\lVert{g_{h}}\rVert_{L^{p}(0,T;\dot{H}_{h}^{0,q})}.$
+Suppose $p,q\in(1,\infty)$. For any $g_{h}\in L^{p}(0,T;\dot{H}_{h}^{0,q})$, the following discrete maximal $L^{p}$-regularity estimate is valid:
+$\lVert{S_{0,h}g_{h}}\rVert_{C([0,T];(\dot{H}_{h}^{0,q},\dot{H}_{h}^{2,q})_{1-1/p,p}}+\lVert{S_{0,h}g_{h}}\rVert_{L^{p}(0,T;\dot{H}_{h}^{2,q})}\leqslant c\lVert{g_{h}}\rVert_{L^{p}(0,T;\dot{H}_{h}^{0,q})}.$
 Lemma 4.11.
@@
 […]
-Additionally, for the discrete deterministic convolution operator $S_{0,h}$, we establish the following property.
-Lemma 4.9.
-Let $p,q\in(1,\infty)$. For any $g_{h}\in L^{p}(0,T;\dot{H}_{h}^{0,q})$, the following inequality holds for all $\epsilon>0$:
-$\|S_{0,h}g_{h}(t)\|_{C([0,T];\dot{H}_{h}^{2-2/p-\epsilon,q})}\leqslant c\|g_{h}\|_{L^{p}(0,T;L^{q})}.$
-Proof.
-For any $t\in[0,T]$, from the identity $S_{0,h}g_{h}(t)=\int_{0}^{t}e^{(t-s)\Delta_{h}}g_{h}(s)\,\mathrm{d}s$, we deduce that
-$\|S_{0,h}g_{h}(t)\|_{\dot{H}_{h}^{2-2/p-\epsilon,q}}$ $\leqslant\int_{0}^{t}\left\|e^{(t-s)\Delta_{h}}g_{h}(s)\right\|_{\dot{H}_{h}^{2-2/p-\epsilon,q}}\,\mathrm{d}s$
-$\stackrel{{\scriptstyle\text{(i)}}}{{\leqslant}}c\int_{0}^{t}(t-s)^{-1+\frac{1}{p}+\frac{\epsilon}{2}}\|g_{h}(s)\|_{L^{q}}\,\mathrm{d}s$
-$\stackrel{{\scriptstyle\text{(ii)}}}{{\leqslant}}c\Big{(}\int_{0}^{t}(t-s)^{(-1+\frac{1}{p}+\frac{\epsilon}{2})\frac{p}{p-1}}\,\mathrm{d}s\Big{)}^{1-\frac{1}{p}}\|g_{h}\|_{L^{p}(0,T;L^{q})},$
-where step (i) uses the estimate Eq. 4.21 with $\alpha=2-2/p-\epsilon$, and step (ii) applies Hölder’s inequality. Since the integral
-$\int_{0}^{t}(t-s)^{\left(-1+\frac{1}{p}+\frac{\epsilon}{2}\right)\frac{p}{p-1}}\,\mathrm{d}s$
-is uniformly bounded with respect to $t\in[0,T]$, the desired inequality follows, thereby completing the proof of the lemma. ∎
+Remark 4.3.
+A proof of Lemma 4.10 can be found in [12, Theorem 3.2].
+Since the boundedness constant of the $H^{\infty}$-calculus for $-\Delta_{h}$ is independent of the spatial mesh size $h$, as shown in [19, Theorem 3.1], Lemma 4.11 follows directly from [32, Theorem 3.5]; see [19, Theorem 3.2] for further details.
 We proceed to present some auxiliary estimates.
 Lemma 4.12.
-Suppose that $v\in L^{\infty}$.
+Suppose that $y_{0}\in L^{\infty}$ is deterministic.
 Consider the mild solution $y$ to equation (1.1), the strong solution $y_{h}$ to equation (2.9), and the function $G$ defined by equation (2.8).
-Additionally, let $G_{h}(t):=e^{t\Delta_{h}}P_{h}v$ for $t\in[0,T]$.
+Additionally, let $G_{h}(t):=e^{t\Delta_{h}}P_{h}y_{0}$ for $t\in[0,T]$.
 For notational simplicity, let $z:=y-G$ and $\xi_{h}:=S_{0,h}(\Delta_{h}P_{h}-P_{h}\Delta)z$.
-For any $p\in(2,\infty)$, $q\in[2,\infty)$, and $\varepsilon>0$, the following inequalities hold:
+For any $p,q\in(2,\infty)$ and $\epsilon>0$, the following inequalities hold:
 $\lVert{G-G_{h}}\rVert_{L^{p}(0,T;L^{q})}+\lVert{P_{h}G-G_{h}}\rVert_{L^{p}(0,T;L^{q})}\leqslant ch^{\frac{2}{p}},$ (4.33)
-$\lVert{z-P_{h}z}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{1+\frac{2}{p}-\varepsilon},$ (4.28)
-$\lVert{\xi_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{1+\frac{2}{p}-\varepsilon},$ (4.29)
-$\lVert{\xi_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{1-\varepsilon},$ (4.30)
-$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{(y-P_{h}y)(t)}\rVert_{L^{4}}^{p}\lVert{y(t)}\rVert_{L^{8}}^{2p}\,\mathrm{d}t\bigg{]}\leqslant ch^{2-\varepsilon}.$ (4.31)
+$\lVert{z-P_{h}z}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{1+\frac{2}{p}-\epsilon},$ (4.34)
+$\lVert{\xi_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{1+\frac{2}{p}-\epsilon},$ (4.35)
+$\lVert{\xi_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{1-\epsilon},$ (4.36)
+$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{(y-P_{h}y)(t)}\rVert_{L^{4}}^{p}\lVert{y(t)}\rVert_{L^{8}}^{2p}\,\mathrm{d}t\bigg{]}\leqslant ch^{2-\epsilon}.$ (4.37)
 Furthermore, for any $\beta\in\Big{(}1,\frac{3p}{3p-4}\Big{)}$, the following inequalities hold:
 $\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{(P_{h}G-G_{h})(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\big{(}\lVert{y(t)}\rVert_{L^{2\beta}}^{p}+\lVert{y_{h}(t)}\rVert_{L^{2\beta}}^{p}\big{)}\,\mathrm{d}t\bigg{]}\leqslant ch^{2-\frac{3}{2}\big{(}1-\frac{1}{\beta}\big{)}p},$ (4.38)
-$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{\xi_{h}(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\big{(}\lVert{y(t)}\rVert_{L^{2\beta}}^{p}+\lVert{y_{h}(t)}\rVert_{L^{2\beta}}^{p}\big{)}\,\mathrm{d}t\bigg{]}\leqslant ch^{p+2-\frac{3}{2}\big{(}1-\frac{1}{\beta}\big{)}p-\varepsilon}.$ (4.33)
-Here, $\beta^{\prime}$ denotes the conjugate exponent of $\beta$.
+$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{\xi_{h}(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\big{(}\lVert{y(t)}\rVert_{L^{2\beta}}^{p}+\lVert{y_{h}(t)}\rVert_{L^{2\beta}}^{p}\big{)}\,\mathrm{d}t\bigg{]}\leqslant ch^{p+2-\frac{3}{2}\big{(}1-\frac{1}{\beta}\big{)}p-\epsilon}.$ (4.39)
+Here, $\beta^{\prime}$ denotes the Hölder conjugate of $\beta$.
 Proof.
@@
 […]
+$=c\lVert{(P_{h}\Delta^{-1}-\Delta_{h}^{-1}P_{h})\Delta z}\rVert_{L^{p}(\Omega\times(0,T);\dot{H}_{h}^{0,q})}$
 […]
-The inequality Eq. 4.30 is derived in a similarly manner:
+$=ch^{1+2/p-\epsilon}\lVert{z}\rVert_{L^{p}(\Omega\times(0,T);\dot{H}^{1+2/p-\epsilon,q})}$
+[…]
+The inequality Eq. 4.36 is derived in a similar manner:
 $\lVert{\xi_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}$
@@
 […]
-The inequality Eq.
-4.31 is straightforward to verify by applying Hölder’s inequality, in conjunction with the approximation property of $P_{h}$ from Eq.
-4.19 and the regularity assertions in Theorem 3.1.
-Now let us prove Eq. 4.32. By invoking Eq. 4.18 and replicating the reasoning behind [19, Equation (2.2)], we establish that
+[…]
+[…]
+[…]
+For Eq. 4.37, we apply Hölder’s inequality with $\alpha\in(1,2)$, yielding
+$\mathbb{E}\left[\int_{0}^{T}\lVert{(y-P_{h}y)(t)}\rVert_{L^{4}}^{p}\lVert{y(t)}\rVert_{L^{8}}^{2p}\,\mathrm{d}t\right]$
+[…]
+[…]
+[…]
+Choosing $\alpha>1$ sufficiently close to 1 so that $\frac{2}{\alpha}-\frac{\epsilon}{2}\geqslant 2-\epsilon$, and using the regularity estimate (3.3), we obtain (4.37).
+Now, let us prove Eq. 4.38. Fix any $\beta\in(1,\frac{3p}{3p-4})$. By using Eq. 4.19 and following the reasoning behind [21, Equation (2.2)], we establish that
 $\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{y_{h}(t)}\rVert_{L^{2}}^{\frac{4\beta}{3\beta-3}-2}\lVert{y_{h}(t)}\rVert_{\dot{H}^{1,2}}^{2}\,\mathrm{d}t\bigg{]}$
@@
 $\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{(P_{h}G-G_{h})(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\lVert{y_{h}(t)}\rVert_{2\beta}^{p}\,\mathrm{d}t\bigg{]}$
-$\leqslant\lVert{P_{h}G-G_{h}}\rVert_{L^{\frac{4p\beta}{4\beta-(3\beta-3)p}}(\Omega\times(0,T);L^{2\beta^{\prime}})}^{p}\lVert{y_{h}}\rVert_{L^{\frac{4\beta}{3\beta-3}}(\Omega\times(0,T);L^{2\beta})}^{p}$
-$\leqslant c\lVert{P_{h}G-G_{h}}\rVert_{L^{\frac{4p\beta}{4\beta-(3\beta-3)p}}(\Omega\times(0,T);L^{2\beta^{\prime}})}^{p}.$
-Invoking the inequality Eq. 4.27 with $p$ replaced by $4p\beta/(4\beta-(3\beta-3)p)$, we find
-$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{(P_{h}G-G_{h})(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\lVert{y_{h}(t)}\rVert_{2\beta}^{p}\,\mathrm{d}t\bigg{]}\leqslant ch^{2-\frac{3}{2}(1-\frac{1}{\beta})p}.$
-Given that Theorem 3.1 ensures $y$ belongs to $L^{\frac{4\beta}{3\beta-3}}(\Omega\times(0,T);L^{2\beta})$, we can replicate the previous argument to conclude that
-$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{(P_{h}G-G_{h})(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\lVert{y(t)}\rVert_{2\beta}^{p}\,\mathrm{d}t\bigg{]}\leqslant ch^{2-\frac{3}{2}(1-1/\beta)p}.$
-Hence, Eq. 4.32 follows by combining the aforementioned inequalities. To deduce Eq. 4.33, we rely on Eq. 4.29 and proceed analogously to the proof of Eq. 4.32. This completes the proof. ∎
-Proof of Theorem 3.2. Fix any $\alpha\in(1,2)$ and $p\in(2,\infty)$. Let
+[…]
+[…]
+Using Eq. 4.33 with $p$ replaced by $\frac{4p\beta}{4\beta-(3\beta-3)p}$ and $q$ by $2\beta^{\prime}$, we conclude that
+$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{(P_{h}G-G_{h})(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\lVert{y_{h}(t)}\rVert_{L^{2\beta}}^{p}\,\mathrm{d}t\bigg{]}\leqslant ch^{2-\frac{3p}{2}(1-\frac{1}{\beta})}.$
+Moreover, by Theorem 3.1, we have $y\in L^{\frac{4\beta}{3(\beta-1)}}(\Omega\times(0,T);L^{2\beta})$, so a similar argument yields
+$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{(P_{h}G-G_{h})(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\lVert{y(t)}\rVert_{L^{2\beta}}^{p}\,\mathrm{d}t\bigg{]}\leqslant ch^{2-\frac{3p}{2}(1-\frac{1}{\beta})}.$
+Combining both estimates gives Eq. 4.38. The estimate Eq. 4.39 follows by a similar argument using Eq. 4.35. This completes the proof.
+Finally, we are in a position to prove Theorems 3.2 and 3.3.
+Proof of Theorem 3.2. Fix any $p\in(2,\infty)$. Define
 $z(t)$ $:=y(t)-G(t),\quad 0\leqslant t\leqslant T,$ (4.40)
@@
 where $G$ is defined by Eq. 2.8, and $G_{h}$ is given by
-$G_{h}(t):=e^{t\Delta_{h}}P_{h}v,\quad t\in[0,T].$ (4.36)
-From Eq. 1.1 and the fact $G^{\prime}(t)=\Delta G(t)$ for $t\in[0,T]$, we deduce that
+$G_{h}(t):=e^{t\Delta_{h}}P_{h}y_{0},\quad t\in[0,T].$ (4.42)
+From Eq. 1.1 and the fact that $G^{\prime}(t)=\Delta G(t)$ for $t\in[0,T]$, we deduce that
 $\mathrm{d}z(t)=\big{(}\Delta z+z+G-y^{3}\big{)}(t)\,\mathrm{d}t+F(y(t))\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T.$
@@
 $\mathrm{d}P_{h}z(t)=P_{h}(\Delta z+z+G-y^{3})(t)\,\mathrm{d}t+P_{h}F(y(t))\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T.$
-Similarly, from Eq. 2.11 and the fact $G_{h}^{\prime}(t)=\Delta_{h}G_{h}(t)$ for $t\in[0,T]$, it follows that
+Similarly, from Eq. 2.9 and the fact that $G_{h}^{\prime}(t)=\Delta_{h}G_{h}(t)$ for $t\in[0,T]$, it follows that
 $\mathrm{d}z_{h}(t)=(\Delta_{h}z_{h}+z_{h}+G_{h}-P_{h}y_{h}^{3})(t)\,\mathrm{d}t+P_{h}F\big{(}y_{h}(t)\big{)}\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T.$
-Defining $e_{h}:=z_{h}-P_{h}z$, we then obtain
+Define $e_{h}:=z_{h}-P_{h}z$. Then we obtain
 $\mathrm{d}e_{h}(t)$ $=\Big{(}\Delta_{h}e_{h}+\Delta_{h}P_{h}z-P_{h}\Delta z+e_{h}+G_{h}-P_{h}G+P_{h}\big{(}y^{3}-y_{h}^{3}\big{)}\Big{)}(t)\,\mathrm{d}t+{}$ (4.43)
 $\quad P_{h}\Big{(}F\big{(}y_{h}(t)\big{)}-F\big{(}y(t)\big{)}\Big{)}\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T.$
-Defining $\eta_{h}:=e_{h}-\xi_{h}$, where $\xi_{h}:=S_{0,h}(\Delta_{h}P_{h}z-P_{h}\Delta z)$, and considering the differential equation
-$\Big{(}\frac{\mathrm{d}}{\mathrm{d}t}-\Delta_{h}\Big{)}\xi_{h}(t)=(\Delta_{h}P_{h}z-P_{h}\Delta z)(t),\quad t\geqslant 0,$
-we use Eq. 4.37 to conclude that
+Now define $\eta_{h}:=e_{h}-\xi_{h}$, where $\xi_{h}:=S_{0,h}(\Delta_{h}P_{h}z-P_{h}\Delta z)$, and consider the differential equation
+$\Big{(}\frac{\mathrm{d}}{\mathrm{d}t}-\Delta_{h}\Big{)}\xi_{h}(t)=(\Delta_{h}P_{h}z-P_{h}\Delta z)(t),\quad t\geqslant 0.$
+Using Eq. 4.43, we conclude that
 $\mathrm{d}\eta_{h}(t)$ $=\Big{(}\Delta_{h}\eta_{h}+\xi_{h}+\eta_{h}+G_{h}-P_{h}G+P_{h}\big{(}y^{3}-y_{h}^{3}\big{)}\Big{)}(t)\,\mathrm{d}t+{}$ (4.44)
-$\quad P_{h}\Big{(}F\big{(}y_{h}(t))-F\big{(}y(t)\big{)}\Big{)}\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T.$
-By applying Itô’s formula and considering the initial condition $\eta_{h}(0)=0$ and the definition of $F$ given by Eq. 2.6, we deduce that for any $0\leqslant t\leqslant T$,
+$\quad P_{h}\Big{(}F\big{(}y_{h}(t)\big{)}-F\big{(}y(t)\big{)}\Big{)}\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T.$
+By applying Itô’s formula to $\lVert{\eta_{h}(t)}\rVert_{L^{2}}^{p}$ and using the initial condition $\eta_{h}(0)=0$ together with the definition of $F$ given by Eq.
+2.4, we deduce that for any $0\leqslant t\leqslant T$,
 $\mathbb{E}\lVert{\eta_{h}(t)}\rVert_{L^{2}}^{p}$ $=p\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot\Big{(}\Delta_{h}\eta_{h}(s)+\xi_{h}(s)+\eta_{h}(s)$
-$\qquad\qquad\qquad\qquad\qquad{}+(G_{h}-P_{h}G)(s)+\big{(}y^{3}-y_{h}^{3}\big{)}(s)\Big{)}\,\mathrm{d}\mu\,\mathrm{d}s\biggr{]}$
-$\qquad{}+\frac{p}{2}\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\sum_{k=0}^{\infty}\lambda_{k}\lVert{P_{h}\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}}\rVert_{L^{2}}^{2}\,\mathrm{d}s\biggr{]}$
-$\qquad{}+\frac{p(p-2)}{2}\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-4}\sum_{k=0}^{\infty}\lambda_{k}\Big{(}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}\,\mathrm{d}\mu\Big{)}^{2}\,\mathrm{d}s\biggr{]}$
+$\qquad\qquad\qquad\qquad\qquad{}+(G_{h}-P_{h}G)(s)+\big{(}y^{3}-y_{h}^{3}\big{)}(s)\Big{)}\,\mathrm{d}x\,\mathrm{d}s\biggr{]}$
+$\qquad{}+\frac{p}{2}\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\sum_{k=0}^{\infty}\lVert{P_{h}\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}}\rVert_{L^{2}}^{2}\,\mathrm{d}s\biggr{]}$
+$\qquad{}+\frac{p(p-2)}{2}\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-4}\sum_{k=0}^{\infty}\Big{(}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}\,\mathrm{d}x\Big{)}^{2}\,\mathrm{d}s\biggr{]}$
 $=I_{1}+I_{2}+I_{3}+I_{4}+I_{5}+I_{6},$
 where
-$\allowdisplaybreaks I_{1}$ $:=p\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot\Delta_{h}\eta_{h}(s)\,\mathrm{d}\mu\,\mathrm{d}s\biggr{]},$
-$I_{2}$ $:=p\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot(\xi_{h}+\eta_{h}+G_{h}-P_{h}G)(s)\,\mathrm{d}\mu\,\mathrm{d}s\biggr{]},$
-$I_{3}$ $:=p\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{[}y^{3}-(P_{h}y)^{3}\big{]}(s)\,\mathrm{d}\mu\,\mathrm{d}s\biggr{]},$
-$I_{4}$ $:=p\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{[}(P_{h}y)^{3}-y_{h}^{3}\big{]}(s)\,\mathrm{d}\mu\,\mathrm{d}s\biggr{]},$
-$I_{5}$ $:=\frac{p}{2}\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\sum_{k=0}^{\infty}\lambda_{k}\lVert{P_{h}\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}}\rVert_{L^{2}}^{2}\,\mathrm{d}s\biggr{]},$
-$I_{6}$ $:=\frac{p(p-2)}{2}\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-4}\sum_{k=0}^{\infty}\lambda_{k}\Big{(}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}\,\mathrm{d}\mu\Big{)}^{2}\,\mathrm{d}s\biggr{]}.$
-Next, let us estimate $I_{1}$, $I_{2}$, $I_{3}$, and $I_{4}$. For the term $I_{1}$, by the definition of $\Delta_{h}$, we establish that
+$\allowdisplaybreaks I_{1}$ $:=p\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot\Delta_{h}\eta_{h}(s)\,\mathrm{d}x\,\mathrm{d}s\biggr{]},$
+$I_{2}$ $:=p\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot(\xi_{h}+\eta_{h}+G_{h}-P_{h}G)(s)\,\mathrm{d}x\,\mathrm{d}s\biggr{]},$
+$I_{3}$ $:=p\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{[}y^{3}-(P_{h}y)^{3}\big{]}(s)\,\mathrm{d}x\,\mathrm{d}s\biggr{]},$
+$I_{4}$ $:=p\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{[}(P_{h}y)^{3}-y_{h}^{3}\big{]}(s)\,\mathrm{d}x\,\mathrm{d}s\biggr{]},$
+$I_{5}$ $:=\frac{p}{2}\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\sum_{k=0}^{\infty}\lVert{P_{h}\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}}\rVert_{L^{2}}^{2}\,\mathrm{d}s\biggr{]},$
+$I_{6}$ $:=\frac{p(p-2)}{2}\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-4}\sum_{k=0}^{\infty}\Big{(}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}\,\mathrm{d}x\Big{)}^{2}\,\mathrm{d}s\biggr{]}.$
+Next, we estimate the terms $I_{1}$, $I_{2}$, $I_{3}$, and $I_{4}$. For the term $I_{1}$, by the definition of $\Delta_{h}$, we obtain
 $I_{1}=-p\mathbb{E}\biggl{[}\int_{0}^{t}\|{\eta_{h}(s)}\|_{L^{2}}^{p-2}\|{\eta_{h}(s)}\|_{\dot{H}^{1,2}}^{2}\,\mathrm{d}s\biggr{]}.$ (4.45)
-Regarding $I_{2}$, by Hölder’s inequality, Young’s inequality, and the inequalities Eq. 4.27 and Eq. 4.29, we can deduce that
-$I_{2}\leqslant ch^{2}+c\int_{0}^{t}\|{\eta_{h}(s)}\|_{L^{p}(\Omega;L^{2})}^{p}\,\mathrm{d}s,$ (4.40)
-For $I_{3}$, by employing Hölder’s inequality, Young’s inequality, and the stability property of $P_{h}$ as stated in Eq. 4.18, a straightforward calculation yields
+Regarding $I_{2}$, applying Hölder’s inequality, Young’s inequality, and the bounds given in Eq. 4.33 and Eq. 4.35, we deduce that
+$I_{2}\leqslant ch^{2}+c\int_{0}^{t}\|{\eta_{h}(s)}\|_{L^{p}(\Omega;L^{2})}^{p}\,\mathrm{d}s.$ (4.46)
+For $I_{3}$, using Hölder’s inequality, Young’s inequality, and the stability property of $P_{h}$ stated in Eq. 4.19, a straightforward computation gives
 $I_{3}$ $\leqslant c\,\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-1}\lVert{(y-P_{h}y)(s)}\rVert_{L^{4}}\Big{(}\lVert{y(s)}\rVert_{L^{8}}^{2}+\lVert{P_{h}y(s)}\rVert_{L^{8}}^{2}\Big{)}\,\mathrm{d}s\bigg{]}$
 $\leqslant c\,\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-1}\lVert{(y-P_{h}y)(s)}\rVert_{L^{4}}\lVert{y(s)}\rVert_{L^{8}}^{2}\,\mathrm{d}s\bigg{]}$
-$\leqslant c\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p}+\lVert{(y-P_{h}y)(s)}\rVert_{L^{4}}^{p}\lVert{y(s)}\rVert_{L^{8}}^{2p}\,\mathrm{d}s\bigg{]}.$
-By utilizing Eq. 4.31, we arrive at the following bound:
+$\leqslant c\mathbb{E}\bigg{[}\int_{0}^{t}\left(\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p}+\lVert{(y-P_{h}y)(s)}\rVert_{L^{4}}^{p}\lVert{y(s)}\rVert_{L^{8}}^{2p}\right)\,\mathrm{d}s\bigg{]}.$
+Using the estimate Eq. 4.37, we then arrive at the bound
 $I_{3}\leqslant ch^{2-\epsilon}+c\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{p}(\Omega;L^{2})}^{p}\,\mathrm{d}s,\quad\forall\epsilon>0.$ (4.47)
-To estimate $I_{4}$, we proceed as follows.
-Using the definitions in (4.34) and (4.35), along with the relation $z_{h}-P_{h}z=\xi_{h}+\eta_{h}$, we deduce that $P_{h}y-y_{h}=-\xi_{h}-\eta_{h}+P_{h}G-G_{h}$.
-For any $s\in[0,t]$, we can then write:
-$\int_{\mathcal{O}}\eta_{h}(s)\left[(P_{h}y)^{3}-y_{h}^{3}\right](s)\,\mathrm{d}\mu$
+To estimate $I_{4}$, we proceed as follows. Using the definitions in (4.40) and (4.41), together with the identity $z_{h}-P_{h}z=\xi_{h}+\eta_{h}$, we find that
+$P_{h}y-y_{h}=-\xi_{h}-\eta_{h}+P_{h}G-G_{h}.$
+Thus, for any $s\in[0,t]$, we compute:
+$\int_{\mathcal{O}}\eta_{h}(s)\left[(P_{h}y)^{3}-y_{h}^{3}\right](s)\,\mathrm{d}x$
 […]
@@
 […]
-Hence, applying Hölder’s inequality, Young’s inequality, and the stability property of $P_{h}$ from (4.18), we obtain the following bound for $I_{4}$:
+Hence, applying Hölder’s inequality, Young’s inequality, and the stability property of $P_{h}$ from (4.19), we derive the following estimate for $I_{4}$:
 $I_{4}$ $\leqslant c\mathbb{E}\int_{0}^{t}\bigg{[}\|\eta_{h}(s)\|_{L^{2}}^{p}+\left(\|\xi_{h}(s)\|^{p}_{L^{2\beta^{\prime}}}+\|(P_{h}G-G_{h})(s)\|^{p}_{L^{2\beta^{\prime}}}\right)$
-×$\qquad\qquad\qquad{}\times\Big{(}\|y(s)\|^{p}_{L^{2\beta}}+\|y_{h}(s)\|^{p}_{L^{2\beta}}\Big{)}\bigg{]}\,\mathrm{d}s,\quad\forall\beta\in\big{(}1,\frac{3p}{3p-4}\big{)}.$
-Using (4.32) and (4.33), we further simplifies this bound to:
+×$\qquad\qquad\qquad{}\times\Big{(}\|y(s)\|^{p}_{L^{2\beta}}+\|y_{h}(s)\|^{p}_{L^{2\beta}}\Big{)}\bigg{]}\,\mathrm{d}s,\quad\forall\beta\in\left(1,\frac{3p}{3p-4}\right),$
+where $\beta^{\prime}$ denotes the Hölder conjugate of $\beta$. Using the estimates (4.38) and (4.39), this simplifies further to
 $I_{4}\leqslant c\int_{0}^{t}\|\eta_{h}(s)\|^{p}_{L^{p}(\Omega;L^{2})}\,\mathrm{d}s+ch^{2-\frac{3}{2}p\left(1-\frac{1}{\beta}\right)},\quad\forall\beta\in\left(1,\frac{3p}{3p-4}\right).$
-By selecting $\beta$ sufficiently small within $\big{(}1,\frac{3p}{3p-4}\big{)}$ for any $\epsilon>0$, we arrive at the final bound for $I_{4}$:
-$I_{4}\leqslant ch^{2-\epsilon}+c\int_{0}^{t}\|\eta_{h}(s)\|^{p}_{L^{p}(\Omega;L^{2})}\,\mathrm{d}s.$ (4.42)
-Now, let us proceed to estimate $I_{5}+I_{6}$. Using Hölder’s inequality and the stability property of $P_{h}$ from Eq. 4.18, we obtain the following bound:
-$I_{5}+I_{6}\leqslant c\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\sum_{k=0}^{\infty}\lambda_{k}\lVert{f_{k}(y_{h}(s))-f_{k}(y(s))}\rVert_{L^{2}}^{2}\,\mathrm{d}s\bigg{]}.$
-Invoking the Lipschitz continuity condition Eq. 2.3 and the summability condition Eq. 2.5, we further deduce that
+By choosing $\beta$ sufficiently close to 1 within the interval $\left(1,\frac{3p}{3p-4}\right)$, for any $\epsilon>0$, we finally obtain
+$I_{4}\leqslant ch^{2-\epsilon}+c\int_{0}^{t}\|\eta_{h}(s)\|^{p}_{L^{p}(\Omega;L^{2})}\,\mathrm{d}s,\quad\forall\epsilon>0.$ (4.48)
+Now, we proceed to estimate $I_{5}+I_{6}$. Using Hölder’s inequality and the stability property of $P_{h}$ given in Eq. 4.19, we obtain the following bound:
+$I_{5}+I_{6}\leqslant c\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\sum_{k=0}^{\infty}\lVert{f_{k}(y_{h}(s))-f_{k}(y(s))}\rVert_{L^{2}}^{2}\,\mathrm{d}s\bigg{]}.$
+Invoking the condition Eq. 2.3, we further deduce that
 $I_{5}+I_{6}\leqslant c\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\lVert{(y_{h}-y)(s)}\rVert_{L^{2}}^{2}\,\mathrm{d}s\bigg{]}.$
-Subsequently, by Young’s inequality, we obtain
-$I_{5}+I_{6}\leqslant c\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p}+\lVert{(y_{h}-y)(s)}\rVert_{L^{2}}^{p}\,\mathrm{d}s\bigg{]}.$
-Recalling the definitions in Eqs. 4.34 and 4.35 and considering the relation $z_{h}-P_{h}z=\xi_{h}+\eta_{h}$, we infer that
-$(y_{h}-y)(s)=\xi_{h}(s)+\eta_{h}(s)-(z-P_{h}z)(s)-(G-G_{h})(s),\quad\forall s\in[0,T],$ (4.43)
-which, in conjunction with Eqs. 4.27, 4.28 and 4.29, yields
-$\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{(y_{h}-y)(s)}\rVert_{L^{2}}^{p}\,\mathrm{d}s\bigg{]}\leqslant ch^{2}+\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{p}(\Omega;L^{2})}^{p}\,\mathrm{d}s.$
-Consequently, we obtain the following bound for $I_{5}+I_{6}$:
+Applying Young’s inequality, we then have
+$I_{5}+I_{6}\leqslant c\mathbb{E}\bigg{[}\int_{0}^{t}\left(\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p}+\lVert{(y_{h}-y)(s)}\rVert_{L^{2}}^{p}\right)\,\mathrm{d}s\bigg{]}.$
+Recalling the definitions in Eqs. 4.40 and 4.41 and using the identity $z_{h}-P_{h}z=\xi_{h}+\eta_{h}$, we infer that
+$(y_{h}-y)(s)=\xi_{h}(s)+\eta_{h}(s)-(z-P_{h}z)(s)-(G-G_{h})(s),\quad\forall s\in[0,T].$ (4.49)
+Combining this with the estimates Eq. 4.33, Eq. 4.34, and Eq. 4.35, we find that
+$\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{(y_{h}-y)(s)}\rVert_{L^{2}}^{p}\,\mathrm{d}s\bigg{]}\leqslant ch^{2}+c\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{p}(\Omega;L^{2})}^{p}\,\mathrm{d}s.$
+As a result, we arrive at the following bound for $I_{5}+I_{6}$:
 $I_{5}+I_{6}\leqslant ch^{2}+c\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{p}(\Omega;L^{2})}^{p}\,\mathrm{d}s.$ (4.50)
-Finally, by integrating the estimates for $I_{1}$, $I_{2}$, $I_{3}$, $I_{4}$, and $I_{5}+I_{6}$, as given by Eqs. 4.39, 4.40, 4.41, 4.42 and 4.44, we obtain the following inequality:
+Finally, by combining the estimates for $I_{1}$, $I_{2}$, $I_{3}$, $I_{4}$, and $I_{5}+I_{6}$, as established in Eqs. 4.45, 4.46, 4.47, 4.48 and 4.50, we obtain
 $\|\eta_{h}(t)\|_{L^{p}(\Omega;L^{2})}^{p}+\mathbb{E}\bigg{[}\int_{0}^{t}\|\eta_{h}(s)\|_{L^{2}}^{p-2}\|\eta_{h}(s)\|_{\dot{H}_{h}^{1,2}}^{2}\,\mathrm{d}s\bigg{]}$
 […]
-Applying Gronwall’s lemma to the above inequality, we infer that
-$\max_{t\in[0,T]}\|\eta_{h}(t)\|_{L^{p}(\Omega;L^{2})}+\left[\mathbb{E}\int_{0}^{T}\|\eta_{h}(t)\|_{L^{2}}^{p-2}\|\eta_{h}(t)\|_{\dot{H}_{h}^{1,2}}^{2}\,\mathrm{d}t\right]^{1/p}\leqslant ch^{\alpha/p}.$
-Moreover, for any $\epsilon>0$ and $t\in(0,T]$, we establish the following bounds:
-$\lVert{(z-P_{h}z)(t)}\rVert_{L^{p}(\Omega;L^{2})}\leqslant ch^{1-\epsilon}$ (by Eqs. 4.19 and 3.4),
+An application of Gronwall’s inequality then yields
+$\max_{t\in[0,T]}\|\eta_{h}(t)\|_{L^{p}(\Omega;L^{2})}\leqslant ch^{(2-\epsilon)/p},\quad\forall\epsilon>0.$
+It follows that
+$\max_{t\in[0,T]}\|\eta_{h}(t)\|_{L^{p}(\Omega;L^{2})}\leqslant ch^{2/p-\epsilon},\quad\forall\epsilon>0.$ (4.51)
+Moreover, for any $\epsilon>0$ and $t\in(0,T]$, the following bounds hold:
+$\lVert{(z-P_{h}z)(t)}\rVert_{L^{p}(\Omega;L^{2})}\leqslant ch^{1-\epsilon},$ (see Eqs. 4.20 and 3.4),
 […]
 […]
-Since the equality Eq. 4.43 implies
+Now, from the decomposition Eq. 4.49, we deduce that
 $\|(y-y_{h})(t)\|_{L^{p}(\Omega;L^{2})}$ $\leqslant\|(z-P_{h}z)(t)\|_{L^{p}(\Omega;L^{2})}+\|(G-G_{h})(t)\|_{L^{p}(\Omega;L^{2})}$
-$\qquad{}+\|\xi_{h}(t)\|_{L^{p}(\Omega;L^{2})}+\|\eta_{h}(t)\|_{L^{p}(\Omega;L^{2})},\quad\forall t\in[0,T],$
-the combination of these bounds yields the desired error estimate Eq. 3.6. This completes the proof of Theorem 3.2.
+$\qquad{}+\|\xi_{h}(t)\|_{L^{p}(\Omega;L^{2})}+\|\eta_{h}(t)\|_{L^{p}(\Omega;L^{2})},\quad\forall t\in[0,T].$
+Combining the above estimates, we conclude that for all $t\in(0,T]$,
+$\lVert{(y-y_{h})(t)}\rVert_{L^{p}(\Omega;L^{2})}\leqslant ch^{2/p-\epsilon}+ch^{1-\epsilon}+ch^{2/p}t^{-1/p},\quad\forall\epsilon>0.$
+Since $h^{1-\epsilon}$ is absorbed into $h^{2/p-\epsilon}$ for $p>2$, the desired error estimate Eq. 3.6 follows immediately. This completes the proof of Theorem 3.2.
 $\blacksquare$
 Proof of Theorem 3.3.
-Let $G$ be defined by Eq.
-2.10, and let $z$ be defined by Eq.
-4.34.
-Extending the reasoning from the proof of Theorem 3.1, we establish the following enhanced regularity properties for the processes $y$ and $z$:
-$y\in L^{p}(\Omega;C([0,T];\dot{H}^{2-\epsilon,q})),$ $\forall p\in(2,\infty),\,\forall q\in[2,\infty),\,\forall\epsilon>0;$ (4.45)
-$z\in L^{p}(\Omega\times(0,T);\dot{H}^{2,q}),$ $\forall p\in(2,\infty),\,\forall q\in[2,\infty);$ (4.46)
-$z\in L^{p}(\Omega;C([0,T];\dot{H}^{2-\epsilon,q})),$ $\forall p\in(2,\infty),\,\forall q\in[2,\infty),\,\forall\epsilon>0.$ (4.47)
-Let $G_{h}$ be defined by Eq. 4.36. We adopt the same notation as introduced in the proof of Theorem 3.2:
+Let $G$ and $z$ be defined by Eqs.
+2.8 and 4.40, respectively.
+Since the initial condition $y_{0}\in W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$ implies that $y_{0}\in\dot{H}^{2,q}$ for all $q\in(2,\infty)$, it follows from Lemma 4.2 that $G\in C([0,T];\dot{H}^{2,q})$ for all $q\in(2,\infty)$.
+Consequently, by adapting the argument in Step 3 of the proof of Theorem 3.1, we obtain the following improved regularity properties for the processes $y$ and $z$:
+$y\in L_{\mathbb{F}}^{p}(\Omega;C([0,T];\dot{H}^{2-\epsilon,q})),$ $\forall p,q\in(2,\infty),\,\forall\epsilon\in(0,1);$ (4.52)
+$z\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,q}),$ $\forall p,q\in(2,\infty);$ (4.53)
+$z\in L_{\mathbb{F}}^{p}(\Omega;C([0,T];\dot{H}^{2-\epsilon,q})),$ $\forall p,q\in(2,\infty),\,\forall\epsilon\in(0,1).$ (4.54)
+Let $G_{h}$ be defined as in Eq. 4.42. We retain the notation introduced in the proof of Theorem 3.2:
 $z_{h}$ $:=y_{h}-G_{h},$ $e_{h}:=z_{h}-P_{h}z,$
 $\xi_{h}$ $:=S_{0,h}(\Delta_{h}P_{h}-P_{h}\Delta)z,$ $\eta_{h}:=e_{h}-\xi_{h}.$
-The remainder of the proof is structured into the following five steps.
+The remainder of the proof is structured into the following three steps.
 Step 1. Let us present some key estimates that form the foundation of our subsequent analysis. We begin by establishing the following estimates:
-$\left\|z-P_{h}z\right\|_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2},\quad\forall p\in(2,\infty),\,\forall q\in[2,\infty),$ (4.48)
-$\left\|G-G_{h}\right\|_{C([0,T];L^{q})}+\left\|P_{h}G-G_{h}\right\|_{C([0,T];L^{q})}\leqslant ch^{2},\quad\forall q\in[2,\infty).$ (4.49)
+$\left\|z-P_{h}z\right\|_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2},\quad\forall p,q\in(2,\infty),$ (4.55)
+$\left\|G-G_{h}\right\|_{C([0,T];L^{q})}+\left\|P_{h}G-G_{h}\right\|_{C([0,T];L^{q})}\leqslant ch^{2},\quad\forall q\in(2,\infty).$ (4.56)
 Estimate (4.55) is a direct consequence of (4.20) combined with the regularity property (4.53).
-Estimate (4.49) follows from (4.22) and the condition $v\in W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$.
-In analogy to the estimate (4.29), we utilize Lemma 4.7, along with (4.20) and the regularity assertion (4.46), to derive the following bound for $\xi_{h}$:
-$\left\|\xi_{h}\right\|_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2},\quad\forall p\in(2,\infty),\,\forall q\in[2,\infty).$ (4.50)
-Following the reasoning from Theorem 3.2, we obtain the following estimate for $\eta_{h}$:
-$\max_{t\in[0,T]}\left\|\eta_{h}(t)\right\|_{L^{p}(\Omega;L^{2})}\leqslant ch^{2},\quad\forall p\in(2,\infty).$ (4.51)
-In light of the decomposition Eq. 4.43, combining Eqs. 4.50, 4.51, 4.49 and 4.48 leads to the error estimate
+Estimate (4.56) follows from (4.29) and the initial condition $y_{0}\in W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$.
+Given the regularity Eq.
+4.53, estimate Eq.
+4.35 improves to
+$\left\|\xi_{h}\right\|_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2},\quad\forall p,q\in(2,\infty),$ (4.57)
+and estimate Eq. 4.36 improves to
+$\lVert{\xi_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2-2/p-\epsilon},\quad\forall p,q\in(2,\infty),\,\forall\epsilon\in(0,1).$ (4.58)
+Following the derivation of estimate Eq. 4.51 for $\eta_{h}$, we obtain the improved bound
+$\max_{0\leqslant t\leqslant T}\left\|\eta_{h}(t)\right\|_{L^{p}(\Omega;L^{2})}\leqslant ch^{2},\quad\forall p\in(2,\infty).$ (4.59)
+Therefore, combining estimates Eqs. 4.57, 4.59, 4.56 and 4.55 via the decomposition Eq. 4.49 yields the error estimate
 $\left\|y-y_{h}\right\|_{L^{p}(\Omega\times(0,T);L^{2})}\leqslant ch^{2},\quad\forall p\in(2,\infty).$ (4.60)
-Furthermore, for any $p\in(2,\infty)$ and $q\in[2,\infty)$, we deduce the following chain of inequalities:
-$\lVert{y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}$
+Furthermore, for any $p\in(2,\infty)$, we deduce the following chain of inequalities:
+$\lVert{y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{\infty})}$
 […]
 […]
-$\stackrel{{\scriptstyle\mathrm{(i)}}}{{\leqslant}}{}$ $\lVert{P_{h}y}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}+ch^{3(1/q-1/2)}\lVert{P_{h}(y-y_{h})}\rVert_{L^{p}(\Omega\times(0,T);L^{2})}$
-$\stackrel{{\scriptstyle\mathrm{(ii)}}}{{\leqslant}}{}$ $c\lVert{y}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}+ch^{3(1/q-1/2)}\lVert{y-y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{2})}$
-$\stackrel{{\scriptstyle\mathrm{(iii)}}}{{\leqslant}}{}$ $c\lVert{y}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}+ch^{2+3(1/q-1/2)},$
-where in (i) we apply the inverse estimate $\lVert{v_{h}}\rVert_{L^{q}}\leqslant ch^{3(1/q-1/2)}\lVert{v_{h}}\rVert_{L^{2}}$ for all $v_{h}\in X_{h}$; in (ii) we utilize the stability property of $P_{h}$ as stated in Eq.
-4.18; and in (iii) we employ the error estimate Eq.
-4.52.
-By invoking the regularity assertion Eq.
-4.45, we subsequently conclude a uniform boundedness with respect to the spatial mesh size $h$:
-$\lVert{y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\text{ is uniformly bounded with respect to $h$ for all $p,q\in[2,\infty)$}.$ (4.53)
-Step 2. Let us prove
-$\lVert{\eta_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{6})}\leqslant ch^{2},\quad\forall p\in(2,\infty).$ (4.54)
-Using the equality Eq. 4.38, along with the definitions of the discrete convolution operators $S_{0,h}$ and $S_{1,h}$ given in Eqs. 4.24 and 4.25, we deduce the following decomposition:
-$\eta_{h}=II_{1}+II_{2}+II_{3},$ (4.55)
+$\stackrel{{\scriptstyle\mathrm{(i)}}}{{\leqslant}}{}$ $\lVert{P_{h}y}\rVert_{L^{p}(\Omega\times(0,T);L^{\infty})}+ch^{-3/2}\lVert{P_{h}(y-y_{h})}\rVert_{L^{p}(\Omega\times(0,T);L^{2})}$
+$\stackrel{{\scriptstyle\mathrm{(ii)}}}{{\leqslant}}{}$ $c\lVert{y}\rVert_{L^{p}(\Omega\times(0,T);L^{\infty})}+ch^{-3/2}\lVert{y-y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{2})}$
+$\stackrel{{\scriptstyle\mathrm{(iii)}}}{{\leqslant}}{}$ $c\lVert{y}\rVert_{L^{p}(\Omega\times(0,T);L^{\infty})}+ch^{1/2},$
+where in (i) we apply the inverse estimate $\lVert{v_{h}}\rVert_{L^{\infty}}\leqslant ch^{-3/2}\lVert{v_{h}}\rVert_{L^{2}}$ for all $v_{h}\in X_{h}$ (see, e.g., [7, Theorem 4.5.11]); in (ii) we utilize the stability property of $P_{h}$ as stated in Eq.
+4.19; and in (iii) we employ the error estimate Eq.
+4.60.
+By Eq.
+4.52 and the continuous embedding $\dot{H}^{2-\epsilon,q}\hookrightarrow L^{\infty}$ for sufficiently small $\epsilon>0$, we conclude $h$-uniform boundedness:
+$\lVert{y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{\infty})}\text{ is uniformly bounded with respect to $h$ for all $p\in(2,\infty)$}.$ (4.61)
+Step 2. We now prove the error estimate Eq. 3.7. The proof comprises two parts, (a) and (b).
+Part (a). We establish the estimate
+$\|\eta_{h}\|_{L^{p}(\Omega\times(0,T);L^{6})}\leqslant ch^{2},\quad\forall p\in(2,\infty).$ (4.62)
+Utilizing (4.44) and the definitions of the discrete convolution operators $S_{0,h}$ and $S_{1,h}$ given in Eqs. 4.30 and 4.31, we derive the decomposition
+$\eta_{h}=II_{1}+II_{2},$ (4.63)
 where
-$II_{1}$ $:=S_{0,h}\big{(}\xi_{h}+\eta_{h}+G_{h}-P_{h}G\big{)},$
-$II_{2}$ $:=S_{0,h}P_{h}(y^{3}-y_{h}^{3}),$
-$II_{3}$ $:=S_{1,h}P_{h}\big{(}F(y_{h})-F(y)\big{)}.$
-Fix an arbitrary $p\in(2,\infty)$. For the term $II_{1}$, we establish the following bound for any $q\in[2,\infty)$:
-$\lVert{II_{1}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}$ $\leqslant c\lVert{S_{0,h}\big{(}\xi_{h}+\eta_{h}+G_{h}-P_{h}G\big{)}}\rVert_{L^{p}(\Omega\times(0,T);\dot{H}_{h}^{2,2})}$
-$\leqslant c\lVert{\xi_{h}+\eta_{h}+G_{h}-P_{h}G}\rVert_{L^{p}(\Omega\times(0,T);L^{2})}$
-$\leqslant ch^{2},$
-where the first inequality uses the discrete Sobolev’s inequality $\lVert{v_{h}}\rVert_{L^{\infty}}\leqslant c\lVert{v_{h}}\rVert_{\dot{H}_{h}^{2,2}}$ for all $v_{h}\in X_{h}$; the subsequent inequality relies on Lemma 4.7; and the final inequality is justified by Eqs.
-4.51, 4.49 and 4.50.
-Turning to $II_{2}$, we establish the bound: for any $q\in[2,\infty)$,
-$\lVert{II_{2}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}$ $\leqslant c\lVert{S_{0,h}P_{h}(y^{3}-y_{h}^{3})}\rVert_{L^{p}(\Omega\times(0,T);\dot{H}_{h}^{2,5/3})}$
-$\leqslant c\lVert{y^{3}-y_{h}^{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{5/3})}$
-$\leqslant c\lVert{y-y_{h}}\rVert_{L^{2p}(\Omega\times(0,T);L^{2})}\Big{(}\lVert{y}\rVert_{L^{4p}(\Omega\times(0,T);L^{20})}^{2}+\lVert{y_{h}}\rVert_{L^{4p}(\Omega\times(0,T);L^{20})}^{2}\Big{)}$
-$\leqslant ch^{2},$
-where the first inequality uses the discrete Sobolev’s inequality $\lVert{v_{h}}\rVert_{L^{\infty}}\leqslant c\lVert{v_{h}}\rVert_{\dot{H}_{h}^{2,5/3}}$ for all $v_{h}\in X_{h}$; the second inequality uses Lemma 4.7 and Eq.
-4.18; the third inequality uses Hölder’s inequality; and the last inequality uses Eqs.
-4.45, 4.52 and 4.53.
-For $II_{3}$ we establish the following bound:
-$\lVert{II_{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{6})}$ $\leqslant c\lVert{S_{1,h}P_{h}\big{(}F(y_{h})-F(y)\big{)}}\rVert_{L^{p}(\Omega\times(0,T);\dot{H}_{h}^{1,2})}$
-$\leqslant c\lVert{y-y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{2})}$
-[…]
-where the first inequality employs the discrete Sobolev’s inequality $\lVert{v_{h}}\rVert_{L^{6}}\leqslant c\lVert{v_{h}}\rVert_{\dot{H}_{h}^{1,2}}$ for all $v_{h}\in X_{h}$, and the second inequality invokes Lemma 4.8, along with Eqs.
-4.18 and 4.4.
-In view of the decomposition Eq.
-4.55, combining the above bounds for $II_{1}$, $II_{2}$, and $II_{3}$ yields Eq.
-4.54.
-Step 3.
-We proceed to establish the error estimate (3.7).
-Considering the decomposition of $y_{h}-y$ given in (4.43), we integrate the inequalities (4.48), (4.49), (4.50), and (4.54) to validate (3.7) for the specific case where $p\in(2,\infty)$ and $q=6$.
-Building upon this outcome and employing a similar bounding technique as utilized for $II_{3}$ in Step 2, we obtain an enhanced bound for $II_{3}$:
-$\lVert{II_{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2},\quad\forall p\in(2,\infty),\,\forall q\in[2,\infty).$
-Integrating this refined bound with the previously established bounds for $II_{1}$ and $II_{2}$ from Step 2, and incorporating the decomposition (4.55), we arrive at the following conclusion:
-$\lVert{\eta_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2},\quad\forall p\in(2,\infty),\,\forall q\in[2,\infty).$ (4.56)
-Ultimately, revisiting the decomposition (4.43) and synthesizing the inequalities (4.48), (4.49), (4.50), and (4.56), we affirm the validity of the error estimate (3.7) for all $p\in(2,\infty)$ and $q\in[2,\infty)$.
-Step 4. Let us demonstrate the following inequality:
-$\lVert{\eta_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2},\quad\forall p\in(2,\infty),\,\forall q\in[2,\infty).$ (4.57)
-To this end, let $p\in(2,\infty)$ and $q\in[2,\infty)$ be arbitrarily chosen.
-Invoking the definitions of $II_{1}$ and $II_{2}$ from (4.55), in conjunction with Lemma 4.9 and the continuous embedding of $\dot{H}_{h}^{1,q}$ into $L^{q}$, we deduce
-$\lVert{II_{1}+II_{2}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant c\lVert{\xi_{h}+\eta_{h}+G_{h}-P_{h}G+P_{h}(y^{3}-y_{h}^{3})}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}.$
-Relying on (4.49), (4.50), and (4.56), it follows that
-$\lVert{II_{1}+II_{2}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2}+c\lVert{P_{h}(y^{3}-y_{h}^{3})}\rVert_{L^{p}(\Omega\times(0,T);L^{q})},$
-which, combined with the stability property of $P_{h}$ as stated in (4.18), leads to
-$\lVert{II_{1}+II_{2}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2}+c\lVert{y^{3}-y_{h}^{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}.$
-Moreover, using Hölder’s inequality yields
-$\lVert{y^{3}-y_{h}^{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}$
-[…]
-which, in light of the regularity result (4.45) and the stability assertion (4.53), simplifies to
-$\lVert{y^{3}-y_{h}^{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant c\lVert{y-y_{h}}\rVert_{L^{2p}(\Omega\times(0,T);L^{2q})}.$
-Consequently, we obtain the following bound for $II_{1}+II_{2}$:
-$\lVert{II_{1}+II_{2}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2}+c\lVert{y-y_{h}}\rVert_{L^{2p}(\Omega\times(0,T);L^{2q})}.$
-We proceed to bound the term $II_{3}$ from Eq. 4.55 as follows:
-$\lVert{II_{3}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}$ $\leqslant c\lVert{P_{h}(F(y_{h})-F(y))}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}$
-$\leqslant c\lVert{F(y_{h})-F(y)}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}$
-$\leqslant c\lVert{y-y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})},$
-where the first inequality is justified by Lemma 4.8 and the continuous embedding of $(\dot{H}_{h}^{0,q},\dot{H}_{h}^{2,q})_{1/2-1/p,p}$ in $L^{q}$; the second inequality is a consequence of the stability of the projection $P_{h}$ from Eq.
-4.18; and the final inequality relies on the Lipschitz continuity of $F$ as stated in Eq.
-4.4.
-Combining the above bounds for $II_{1}+II_{2}$ and $II_{3}$, we arrive at
-$\lVert{II_{1}+II_{2}+II_{3}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}$
-[…]
-Given the decomposition Eq. 4.55 and the derived error estimate Eq. 3.7 for all $p\in(2,\infty)$ and $q\in[2,\infty)$, the desired inequality Eq. 4.57 is thereby established.
-Step 5.
-We proceed to prove the error estimate Eq.
-3.8.
-Fix arbitrarily $p\in(2,\infty)$, $q\in[2,\infty)$, and $\epsilon>0$.
-Using the regularity assertion Eq.
-4.46 and following the proof technique used for Eq.
-4.30, we deduce the following bound for $\xi_{h}$:
-$\lVert{\xi_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2-\epsilon}.$
-Employing Eq. 4.19 and the regularity assertion Eq. 4.47, we establish the following bound for $z-P_{h}z$:
-$\|z-P_{h}z\|_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2-\epsilon}.$
-Furthermore, by combining Eqs. 4.49 and 4.57, we obtain the following bound:
-$\|G-G_{h}+\eta_{h}\|_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2}.$
-Since the decomposition Eq. 4.43 implies
+$II_{1}$ $:=S_{0,h}\big{(}\xi_{h}+\eta_{h}+G_{h}-P_{h}G+P_{h}(y^{3}-y_{h}^{3})\big{)},$
+$II_{2}$ $:=S_{1,h}P_{h}\big{(}F(y_{h})-F(y)\big{)}.$
+Fix $p\in(2,\infty)$. For $II_{1}$, applying Lemma 4.8(i) and Lemma 4.10 yields
+$\|II_{1}\|_{L^{p}(\Omega\times(0,T);L^{\infty})}$ $\leqslant c\|II_{1}\|_{L^{p}(\Omega\times(0,T);\dot{H}_{h}^{2,5/3})}$
+$\leqslant c\|\xi_{h}+\eta_{h}+G_{h}-P_{h}G+P_{h}(y^{3}-y_{h}^{3})\|_{L^{p}(\Omega\times(0,T);L^{5/3})}$
+$\leqslant c\|\xi_{h}\|_{L^{p}(\Omega\times(0,T);L^{5/3})}+c\|\eta_{h}\|_{L^{p}(\Omega\times(0,T);L^{5/3})}$
+$\quad{}+c\|G_{h}-P_{h}G\|_{L^{p}(\Omega\times(0,T);L^{5/3})}+c\|P_{h}(y^{3}-y_{h}^{3})\|_{L^{p}(\Omega\times(0,T);L^{5/3})}.$
+Employing the estimates Eqs. 4.59, 4.56 and 4.57 and the embedding $L^{q}\hookrightarrow L^{5/3}$ for $q>5/3$, we obtain
+$\|II_{1}\|_{L^{p}(\Omega\times(0,T);L^{\infty})}\leqslant ch^{2}+\|P_{h}(y^{3}-y_{h}^{3})\|_{L^{p}(\Omega\times(0,T);L^{5/3})}.$
+Since $P_{h}$ is bounded on $L^{5/3}$ (see (4.19)), it follows that
+$\|II_{1}\|_{L^{p}(\Omega\times(0,T);L^{\infty})}\leqslant ch^{2}+\|y^{3}-y_{h}^{3}\|_{L^{p}(\Omega\times(0,T);L^{5/3})}.$
+To estimate $\|y^{3}-y_{h}^{3}\|_{L^{p}(\Omega\times(0,T);L^{5/3})}$, we use the identity
+$y^{3}-y_{h}^{3}=(y-y_{h})(y^{2}+yy_{h}+y_{h}^{2})$
+and apply Hölder’s inequality, leading to
+$\|y^{3}-y_{h}^{3}\|_{L^{p}(\Omega\times(0,T);L^{5/3})}\leqslant c\|y-y_{h}\|_{L^{2p}(\Omega\times(0,T);L^{2})}\Big{(}\|y\|_{L^{4p}(\Omega\times(0,T);L^{20})}^{2}+\|y_{h}\|_{L^{4p}(\Omega\times(0,T);L^{20})}^{2}\Big{)}.$
+By the regularity of $y$ (see (4.52)), the stability of $y_{h}$ (see (4.61)), and the error estimate (4.60), we deduce
+$\|y^{3}-y_{h}^{3}\|_{L^{p}(\Omega\times(0,T);L^{5/3})}\leqslant ch^{2}.$ (4.64)
+Therefore,
+$\|II_{1}\|_{L^{p}(\Omega\times(0,T);L^{\infty})}\leqslant ch^{2}.$ (4.65)
+For $II_{2}$, the continuous embedding $\dot{H}^{1,2}\hookrightarrow L^{6}$ and the isometric embedding of $\dot{H}_{h}^{1,2}$ into $\dot{H}^{1,2}$ imply
+$\|II_{2}\|_{L^{p}(\Omega\times(0,T);L^{6})}\leqslant c\|II_{2}\|_{L^{p}(\Omega\times(0,T);\dot{H}_{h}^{1,2})}.$
+Hence, applying Lemma 4.11, the ideal property of $\gamma$-radonifying operators [16, Theorem 9.1.10], the boundedness of $P_{h}$ on $L^{2}$, the Lipschitz condition (4.4), and the estimate (4.60), we obtain
+$\|II_{2}\|_{L^{p}(\Omega\times(0,T);L^{6})}$ $\leqslant c\|P_{h}\big{(}F(y_{h})-F(y)\big{)}\|_{L^{p}(\Omega\times(0,T);\gamma(H,\dot{H}_{h}^{0,2}))}$
+$\leqslant c\|F(y_{h})-F(y)\|_{L^{p}(\Omega\times(0,T);\gamma(H,L^{2}))}$
+$\leqslant c\|y-y_{h}\|_{L^{p}(\Omega\times(0,T);L^{2})}$
+$\leqslant ch^{2}.$
+Combining (4.65) with this estimate via the decomposition (4.63) establishes (4.62).
+Part (b).
+We consider the decomposition of $y_{h}-y$ given in Eq.
+4.49, and combine it with the estimates Eq.
+4.55, Eq.
+4.56, Eq.
+4.57, and Eq.
+4.62 to establish Eq.
+3.7 for the particular case $p\in(2,\infty)$, $q=6$.
+Based on this, we now apply Lemma 4.8(ii) and adapt the same bounding strategy used for $II_{2}$ in part (a).
+This yields an improved estimate:
+$\lVert{II_{2}}\rVert_{L^{p}(\Omega\times(0,T);L^{\infty})}$ $\leqslant c\lVert{II_{2}}\rVert_{L^{p}(\Omega\times(0,T);\dot{H}_{h}^{1,6})}$
+$\leqslant c\lVert{P_{h}(F(y_{h})-F(y))}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,\dot{H}_{h}^{0,6}))}$
+$\leqslant c\lVert{F(y_{h})-F(y)}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{6}))}$
+$\leqslant c\lVert{y_{h}-y}\rVert_{L^{p}(\Omega\times(0,T);L^{6})}$
+$\leqslant ch^{2},\quad\forall p\in(2,\infty).$
+Combining this refined bound with the estimate for $II_{1}$ given in Eq. 4.65, and recalling the decomposition Eq. 4.63, we deduce that
+$\lVert{\eta_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{\infty})}\leqslant ch^{2},\quad\forall p\in(2,\infty).$ (4.66)
+Finally, revisiting the decomposition Eq. 4.49 and combining the estimates Eq. 4.55, Eq. 4.56, Eq. 4.57, and Eq. 4.66, we conclude the error estimate Eq. 3.7 for all $p,q\in(2,\infty)$.
+Step 3. We now establish the error estimate Eq. 3.8, proceeding in three parts, (a)–(c).
+Part (a). We aim to prove the inequality
+$\|\eta_{h}\|_{L^{p}(\Omega;C([0,T];L^{\infty}))}\leqslant ch^{2},\quad\forall p\in(4,\infty).$ (4.67)
+To this end, fix $p\in(4,\infty)$ and let $q\in(6,\infty)$ be arbitrary. Recall the decomposition Eq. 4.63 for $\eta_{h}$. By Lemma 4.8(ii) and Lemma 4.7, we have
+$\|II_{1}\|_{L^{p}(\Omega;C([0,T];L^{\infty}))}$ $\leqslant\|II_{1}\|_{L^{p}(\Omega;C([0,T];\dot{H}_{h}^{2-3/p,q}))}$
+$\leqslant\|II_{1}\|_{L^{p}(\Omega;C([0,T];(\dot{H}_{h}^{0,q},\dot{H}_{h}^{2,q})_{1/2-1/p,p}))}.$
+An application of Lemma 4.10 then gives
+$\|II_{1}\|_{L^{p}(\Omega;C([0,T];L^{\infty}))}$ $\leqslant c\|\xi_{h}+\eta_{h}+G_{h}-P_{h}G+P_{h}(y^{3}-y_{h}^{3})\|_{L^{p}(\Omega\times(0,T);L^{q})}.$
+Using the estimates Eq. 4.56, Eq. 4.57, and Eq. 4.66, we deduce
+$\|II_{1}\|_{L^{p}(\Omega;C([0,T];L^{\infty}))}$ $\leqslant ch^{2}+c\|P_{h}(y^{3}-y_{h}^{3})\|_{L^{p}(\Omega\times(0,T);L^{q})}.$
+Since the projection operator $P_{h}$ is bounded on $L^{q}$ (cf. Eq. 4.19), it follows that
+$\|II_{1}\|_{L^{p}(\Omega;C([0,T];L^{\infty}))}$ $\leqslant ch^{2}+c\|y^{3}-y_{h}^{3}\|_{L^{p}(\Omega\times(0,T);L^{q})}.$
+As in Eq. 4.64, we use the identity
+$y^{3}-y_{h}^{3}=(y-y_{h})(y^{2}+yy_{h}+y_{h}^{2})$
+and apply Hölder’s inequality to obtain
+$\|y^{3}-y_{h}^{3}\|_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant c\|y-y_{h}\|_{L^{2p}(\Omega\times(0,T);L^{2q})}.$
+Therefore,
+$\|II_{1}\|_{L^{p}(\Omega;C([0,T];L^{\infty}))}$ $\leqslant ch^{2}+c\|y-y_{h}\|_{L^{2p}(\Omega\times(0,T);L^{2q})}.$
+Next, we estimate $II_{2}$ from Eq. 4.63. Applying Lemma 4.7 and Lemma 4.8(ii), we get
+$\|II_{2}\|_{L^{p}(\Omega;C([0,T];L^{\infty}))}$ $\leqslant\|II_{2}\|_{L^{p}(\Omega;C([0,T];(\dot{H}_{h}^{0,q},\dot{H}_{h}^{2,q})_{1/2-1/p,p}))}$
+$\leqslant c\|P_{h}(F(y_{h})-F(y))\|_{L^{p}(\Omega\times(0,T);\gamma(H,\dot{H}_{h}^{0,q}))},$
+where the second inequality follows from Lemma 4.11. As in the estimate of $II_{2}$ in Step 2, Part (a), we then obtain
+$\|II_{2}\|_{L^{p}(\Omega;C([0,T];L^{\infty}))}\leqslant c\|y-y_{h}\|_{L^{p}(\Omega\times(0,T);L^{q})}.$
+Combining the bounds for $II_{1}$ and $II_{2}$, and recalling the decomposition Eq. 4.63, we arrive at
+$\|\eta_{h}\|_{L^{p}(\Omega;C([0,T];L^{\infty}))}$ $\leqslant ch^{2}+c\|y-y_{h}\|_{L^{p}(\Omega\times(0,T);L^{q})}+c\|y-y_{h}\|_{L^{2p}(\Omega\times(0,T);L^{2q})}.$
+Now, since the error estimate Eq. 3.7 has already been established for all $p,q\in(2,\infty)$, we conclude that the right-hand side is of order $h^{2}$. This completes the proof of Eq. 4.67.
+Part (b). We now prove the estimate
+$\|y-y_{h}\|_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2-2/p-\varepsilon},\quad\forall p,q\in(2,\infty),\ \forall\varepsilon\in(0,1).$ (4.68)
+By the approximation property Eq. 4.20 and the regularity result Eq. 4.54, we have
+$\|z-P_{h}z\|_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2-\varepsilon},\quad\forall p,q\in(2,\infty),\ \forall\varepsilon\in(0,1).$ (4.69)
+Moreover, combining the approximation result Eq. 4.56 with the bound Eq. 4.67, we deduce
+$\|G-G_{h}+\eta_{h}\|_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2},\quad\forall p,q\in(2,\infty).$ (4.70)
+Using the decomposition Eq. 4.49, we write
 $\|y-y_{h}\|_{L^{p}(\Omega;C([0,T];L^{q}))}$ $\leqslant\|\xi_{h}\|_{L^{p}(\Omega;C([0,T];L^{q}))}+\|z-P_{h}z\|_{L^{p}(\Omega;C([0,T];L^{q}))}$
-$\qquad{}+\|G-G_{h}+\eta_{h}\|_{L^{p}(\Omega;C([0,T];L^{q}))},$
-the desired error estimate Eq. 3.8 follows by combining the above bounds. This concludes the proof of Theorem 3.3.
+$\quad+\|G-G_{h}+\eta_{h}\|_{L^{p}(\Omega;C([0,T];L^{q}))}.$
+Substituting the bounds Eqs. 4.69 and 4.70 into this inequality, we obtain
+$\|y-y_{h}\|_{L^{p}(\Omega;C([0,T];L^{q}))}$ $\leqslant ch^{2-\varepsilon}+\|\xi_{h}\|_{L^{p}(\Omega;C([0,T];L^{q}))},\quad\forall p,q\in(2,\infty),\ \forall\varepsilon\in(0,1).$
+Finally, invoking the estimate Eq. 4.58, we conclude the desired result Eq. 4.68.
+Part (c). Let $\Pi_{h}$ denote the standard Lagrange interpolation operator onto the finite element space $X_{h}$. Combining the regularity result Eq. 4.52 with standard interpolation theory yields
+$\lVert{y-\Pi_{h}y}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}$ $\leqslant ch^{2-\epsilon},\quad\forall p,q\in(2,\infty),\,\forall\epsilon\in(0,1).$
+The triangle inequality applied to this estimate and Eq. 4.68 then gives
+$\lVert{y_{h}-\Pi_{h}y}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}$ $\leqslant ch^{2-2/p-\epsilon},\quad\forall p,q\in(2,\infty),\,\forall\epsilon\in(0,1).$
+Applying the standard finite element inverse estimate (cf. [7, Theorem 4.5.11])
+$\lVert{v_{h}}\rVert_{L^{\infty}}\leqslant ch^{-3/q}\lVert{v_{h}}\rVert_{L^{q}},\quad\forall v_{h}\in X_{h},\,q\in(2,\infty),$
+to the preceding bound implies
+$\lVert{y_{h}-\Pi_{h}y}\rVert_{L^{p}(\Omega;C([0,T];L^{\infty}))}$ $\leqslant ch^{2-2/p-3/q-\epsilon},\quad\forall p,q\in(2,\infty),\,\forall\epsilon\in(0,1).$
+Furthermore, the regularity Eq. 4.52 combined with standard interpolation estimates yields
+$\lVert{y-\Pi_{h}y}\rVert_{L^{p}(\Omega;C([0,T];L^{\infty}))}$ $\leqslant ch^{2-3/q-\epsilon},\quad\forall p\in(2,\infty),\,\forall q\in(3,\infty),\,\forall\epsilon\in(0,1).$
+Applying the triangle inequality to these last two estimates now provides
+$\lVert{y-y_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{\infty}))}$ $\leqslant ch^{2-2/p-3/q-\epsilon},\quad\forall p,q\in(2,\infty),\,\forall\epsilon\in(0,1).$
+Since $p,q\in(2,\infty)$ can be chosen arbitrarily large, the term $2/p+3/q$ can be made arbitrarily small and is therefore absorbed into $\epsilon>0$.
+This yields the desired error estimate Eq.
+3.8, completing the proof of Theorem 3.3.
 $\blacksquare$
 5 Conclusions
-In this work, we have rigorously investigated the regularity properties of the three-dimensional stochastic Allen-Cahn equation with multiplicative noise, particularly in the presence of non-smooth initial data.
-For the spatial semidiscretization using the finite element method, we have established a convergence rate of $O(h^{\alpha/p}+h^{2/p}t^{-1/p})$ with non-smooth initial data.
+In this work, we have rigorously investigated the regularity properties of the three-dimensional stochastic Allen-Cahn equation with multiplicative noise, in the presence of non-smooth initial data.
+For the spatial semidiscretization using the finite element method, we have established a convergence rate of $O(h^{2/p-\epsilon}+h^{2/p}t^{-1/p})$ with non-smooth initial data.
 Meanwhile, under the condition of smooth initial data, employing the discrete stochastic maximal $L^{p}$-regularity theory, we have achieved a convergence rate of $O(h^{2})$ in the $L^{p}(\Omega\times(0,T);L^{q})$ norm, along with a pathwise uniform convergence rate of $O(h^{2-\epsilon})$.
-Significantly, these findings cover a variety of spatial $L^{q}$-norms, enhancing their adaptability and universality.
-The theoretical results derived here lay a foundation for the numerical analysis of fully discrete approximations based on the finite element method for the three-dimensional stochastic Allen-Cahn equation with multiplicative noise, especially within the context of general spatial $L^{q}$-norms.
+The theoretical results derived here lay a foundation for the numerical analysis of fully discrete approximations based on the finite element method for the three-dimensional stochastic Allen-Cahn equation with multiplicative noise, within the context of general spatial $L^{q}$-norms.
+Appendix A Proofs of Lemmas 4.8 and 4.9
+The discrete Laplace operator $\Delta_{h}$ has the following fundamental property; see [1].
+Lemma A.1.
+For any $q\in(1,\infty)$ and $\theta\in(\pi/2,\pi)$, there exists a constant $c>0$ independent of $h$, depending only on $\theta$, $q$, and the regularity constants of $\mathcal{K}_{h}$, such that
+$\left\|(re^{i\omega}-\Delta_{h})^{-1}\right\|_{\mathcal{L}(\dot{H}_{h}^{0,q})}\leqslant\frac{c}{1+r},\quad\forall r\geqslant 0,\quad\forall\omega\in[-\theta,\theta].$
+Lemma A.2.
+For any $\alpha\in[0,2]$ and $q\in(1,\infty)$, the operator norm $\|P_{h}\|_{\mathcal{L}(\dot{H}^{\alpha,q},\dot{H}_{h}^{\alpha,q})}$ is uniformly bounded with respect to $h$.
+Proof.
+Fix $q\in(1,\infty)$. By the estimates Eqs. 4.20 and 4.21, and applying the triangle inequality, we obtain
+$\lVert{P_{h}-\Delta_{h}^{-1}P_{h}\Delta}\rVert_{\mathcal{L}(\dot{H}^{2,q},L^{q})}\leqslant ch^{2}.$
+Using the inverse estimate (4.26), it follows that
+$\lVert{P_{h}-\Delta_{h}^{-1}P_{h}\Delta}\rVert_{\mathcal{L}(\dot{H}^{2,q},\dot{H}_{h}^{2,q})}\leqslant c.$
+On the other hand, the stability estimate (4.19) gives
+$\lVert{\Delta_{h}^{-1}P_{h}\Delta}\rVert_{\mathcal{L}(\dot{H}^{2,q},\dot{H}_{h}^{2,q})}=\lVert{P_{h}\Delta}\rVert_{\mathcal{L}(\dot{H}^{2,q},L^{q})}\leqslant c\lVert{\Delta}\rVert_{\mathcal{L}(\dot{H}^{2,q},L^{q})}=c,$
+where we have used the fact that $\Delta:\dot{H}^{2,q}\to L^{q}$ is an isometric isomorphism. Combining the above bounds and applying the triangle inequality, we conclude that
+$\lVert{P_{h}}\rVert_{\mathcal{L}(\dot{H}^{2,q},\dot{H}_{h}^{2,q})}\leqslant c.$
+This proves the uniform boundedness for $\alpha=2$.
+The case $\alpha=0$ is already established by (4.19).
+The general case $\alpha\in[0,2]$ follows from complex interpolation between the endpoint cases $\alpha=0$ and $\alpha=2$, which completes the proof.
+∎
+By the approximation property of $P_{h}$ in Eq. 4.20 and the smoothing estimate for $e^{t\Delta}$ in Lemma 4.2, a direct calculation yields the following lemma.
+Lemma A.3.
+For any $\alpha\in[0,2]$, $q\in(1,\infty)$, and $t>0$,
+$\left\|(I-P_{h})e^{t\Delta}\right\|_{\mathcal{L}(\dot{H}^{\alpha,q},L^{q})}\leqslant c\min\left\{h^{\alpha},\,\frac{h^{2}}{t^{1-\alpha/2}}\right\}.$ (A.1)
+A.1 Proof of Lemma 4.8
+We only prove (ii), as (i) follows by similar arguments using Eq.
+4.21 instead of Eq.
+4.23.
+Fix $\alpha\in(0,2)$ and $q\in(3/\alpha,\infty)\cap[2,\infty)$, and let $v_{h}\in X_{h}$ be arbitrary.
+By Eq.
+4.23, we have
+$\|v_{h}-\Delta^{-1}\Delta_{h}v_{h}\|_{L^{q}}=\|-(I-\Delta_{h}^{-1}P_{h}\Delta)\Delta^{-1}\Delta_{h}v_{h}\|_{L^{q}}\leqslant ch^{\alpha}\|\Delta^{-1}\Delta_{h}v_{h}\|_{\dot{H}^{\alpha,q}}.$
+Moreover, Eq. 4.20 gives
+$\|(I-P_{h})\Delta^{-1}\Delta_{h}v_{h}\|_{L^{q}}\leqslant ch^{\alpha}\|\Delta^{-1}\Delta_{h}v_{h}\|_{\dot{H}^{\alpha,q}}.$
+Combining these two estimates via the triangle inequality yields
+$\|v_{h}-P_{h}\Delta^{-1}\Delta_{h}v_{h}\|_{L^{q}}\leqslant ch^{\alpha}\|\Delta^{-1}\Delta_{h}v_{h}\|_{\dot{H}^{\alpha,q}}.$
+Applying a standard inverse estimate [7, Theorem 4.5.11], we deduce
+$\|v_{h}-P_{h}\Delta^{-1}\Delta_{h}v_{h}\|_{L^{\infty}}\leqslant ch^{\alpha-3/q}\|\Delta^{-1}\Delta_{h}v_{h}\|_{\dot{H}^{\alpha,q}}.$
+Since $\alpha>3/q$, it follows that
+$\|v_{h}-P_{h}\Delta^{-1}\Delta_{h}v_{h}\|_{L^{\infty}}\leqslant c\|\Delta^{-1}\Delta_{h}v_{h}\|_{\dot{H}^{\alpha,q}}.$
+The Sobolev embedding $\dot{H}^{\alpha,q}\hookrightarrow L^{\infty}$, which holds due to $\alpha q>3$, together with the $L^{\infty}$-stability of $P_{h}$ (see Eq. 4.19), implies
+$\|P_{h}\Delta^{-1}\Delta_{h}v_{h}\|_{L^{\infty}}\leqslant c\|\Delta^{-1}\Delta_{h}v_{h}\|_{\dot{H}^{\alpha,q}}.$
+Applying the triangle inequality once more, we obtain
+$\|v_{h}\|_{L^{\infty}}\leqslant c\|\Delta^{-1}\Delta_{h}v_{h}\|_{\dot{H}^{\alpha,q}}=c\|\Delta_{h}v_{h}\|_{\dot{H}^{\alpha-2,q}}.$
+To estimate $\|\Delta_{h}v_{h}\|_{\dot{H}^{\alpha-2,q}}$, we use the dual characterization
+$\|\Delta_{h}v_{h}\|_{\dot{H}^{\alpha-2,q}}=\sup_{\begin{subarray}{c}\varphi\in\dot{H}^{2-\alpha,q^{\prime}}\\ \varphi\neq 0\end{subarray}}\frac{\langle\Delta_{h}v_{h},\varphi\rangle}{\|\varphi\|_{\dot{H}^{2-\alpha,q^{\prime}}}}.$
+Since $\Delta_{h}v_{h}\in X_{h}$, we may replace $\varphi$ with $P_{h}\varphi$, yielding
+$\|\Delta_{h}v_{h}\|_{\dot{H}^{\alpha-2,q}}=\sup_{\begin{subarray}{c}\varphi\in\dot{H}^{2-\alpha,q^{\prime}}\\ \varphi\neq 0\end{subarray}}\frac{\int_{\mathcal{O}}(-\Delta_{h})^{\alpha/2}v_{h}\cdot(-\Delta_{h})^{1-\alpha/2}P_{h}\varphi\,\mathrm{d}x}{\|\varphi\|_{\dot{H}^{2-\alpha,q^{\prime}}}}.$
+By Hölder’s inequality and Lemma A.2, we deduce
+$\|\Delta_{h}v_{h}\|_{\dot{H}^{\alpha-2,q}}\leqslant\sup_{\begin{subarray}{c}\varphi\in\dot{H}^{2-\alpha,q^{\prime}}\\ \varphi\neq 0\end{subarray}}\frac{\|v_{h}\|_{\dot{H}_{h}^{\alpha,q}}\|P_{h}\varphi\|_{\dot{H}_{h}^{2-\alpha,q^{\prime}}}}{\|\varphi\|_{\dot{H}^{2-\alpha,q^{\prime}}}}\leqslant c\|v_{h}\|_{\dot{H}_{h}^{\alpha,q}}.$
+Combining the above estimates, we conclude
+$\|v_{h}\|_{L^{\infty}}\leqslant c\|v_{h}\|_{\dot{H}_{h}^{\alpha,q}},$
+which completes the proof of (ii), and hence the lemma.
+$\blacksquare$
+A.2 Proof of Lemma 4.9
+Property (i) follows directly from Proposition 2.5 and Equation (2.128) in [37], in light of Lemma A.1.
+To prove property (ii), let $\alpha\in[0,2]$, $q\in(1,\infty)$, and $t>0$ be arbitrary.
+By Lemma A.3, the desired estimate for the second term on the left-hand side of Eq.
+4.29 is already established.
+It therefore remains to show that
+$\lVert{e^{t\Delta_{h}}P_{h}-P_{h}e^{t\Delta}}\rVert_{\mathcal{L}(\dot{H}^{\alpha,q},L^{q})}\leqslant c\min\left\{h^{\alpha},\frac{h^{2}}{t^{1-\alpha/2}}\right\}.$ (A.2)
+This estimate is well known and can be derived using arguments found in [31, Theorem 6.10, Lemma 20.7]. For the sake of completeness, we provide a concise proof of (A.2) in two steps.
+Step 1. We establish the estimate
+$\left\|e^{t\Delta_{h}}P_{h}v-P_{h}e^{t\Delta}v\right\|_{L^{q}}\leqslant ch^{2}\|v\|_{\dot{H}^{2,q}},\quad\forall v\in\dot{H}^{2,q}.$ (A.3)
+Fix $v\in\dot{H}^{2,q}$. Utilizing the Laplace inversion formula, we represent the semigroups as:
+$e^{t\Delta}v$ $=\frac{1}{2\pi i}\int_{\Gamma}e^{t\lambda}\lambda^{-1}(\lambda-\Delta)^{-1}\Delta v\,\mathrm{d}\lambda+v,$
+$e^{t\Delta_{h}}P_{h}v$ $=\frac{1}{2\pi i}\int_{\Gamma}e^{t\lambda}\lambda^{-1}(\lambda-\Delta_{h})^{-1}\Delta_{h}P_{h}v\,\mathrm{d}\lambda+P_{h}v,$
+where the contour $\Gamma$, oriented such that the negative real axis lies to its left, is defined by
+$\Gamma=\left\{re^{\pm i\theta}\mid\,r\in[t^{-1},\infty)\right\}\cup\left\{t^{-1}e^{i\omega}|\,\omega\in[-\theta,\theta]\right\},$
+with $\theta\in(\pi/2,\pi)$ and $i$ denoting the imaginary unit. Subtracting these expressions yields:
+$e^{t\Delta_{h}}P_{h}v-P_{h}e^{t\Delta}v=\frac{1}{2\pi i}\int_{\Gamma}e^{t\lambda}\lambda^{-1}\left[(\lambda-\Delta_{h})^{-1}\Delta_{h}P_{h}-P_{h}(\lambda-\Delta)^{-1}\Delta\right]v\,\mathrm{d}\lambda.$ (A.4)
+A direct computation establishes the identity
+$\left[(\lambda-\Delta_{h})^{-1}\Delta_{h}P_{h}-P_{h}(\lambda-\Delta)^{-1}\Delta\right]v=(\lambda-\Delta_{h})^{-1}\Delta_{h}\left(P_{h}\Delta^{-1}-\Delta_{h}^{-1}P_{h}\right)\lambda(\lambda-\Delta)^{-1}\Delta v.$
+This implies
+$\left\|\left((\lambda-\Delta_{h})^{-1}\Delta_{h}P_{h}-P_{h}(\lambda-\Delta)^{-1}\Delta\right)v\right\|_{L^{q}}$
+[…]
+[…]
+[…]
+where the second inequality follows from Lemma A.1 and Eq.
+4.25 with $\alpha=2$, and the third inequality uses the uniform boundedness of $\|\lambda(\lambda-\Delta)^{-1}\|_{\mathcal{L}(L^{q})}$ for $\lambda\in\Gamma$.
+Applying this bound to the integral representation in (A.4), we deduce
+$\lVert{e^{t\Delta_{h}}P_{h}v-P_{h}e^{t\Delta}v}\rVert_{L^{q}}\leqslant ch^{2}\lVert{v}\rVert_{\dot{H}^{2,q}}\int_{\Gamma}\big{|}e^{t\lambda}\big{|}\,\frac{|\mathrm{d}\lambda|}{|\lambda|}\leqslant ch^{2}\lVert{v}\rVert_{\dot{H}^{2,q}},$
+which establishes (A.3).
+Step 2. For arbitrary $v\in L^{q}$, the Laplace inversion formula gives
+$e^{t\Delta}v$ $=\frac{1}{2\pi i}\int_{\Gamma}e^{t\lambda}(\lambda-\Delta)^{-1}v\,\mathrm{d}\lambda,$
+$e^{t\Delta_{h}}P_{h}v$ $=\frac{1}{2\pi i}\int_{\Gamma}e^{t\lambda}(\lambda-\Delta_{h})^{-1}P_{h}v\,\mathrm{d}\lambda,$
+where $\Gamma$ is the contour defined in Step 1. Therefore, the difference is
+$e^{t\Delta_{h}}P_{h}v-P_{h}e^{t\Delta}v$ $=\frac{1}{2\pi i}\int_{\Gamma}e^{t\lambda}\bigl{[}(\lambda-\Delta_{h})^{-1}P_{h}-P_{h}(\lambda-\Delta)^{-1}\bigr{]}v\,\mathrm{d}\lambda.$
+Mimicking the derivation of (A.3) leads to the low-regularity estimate
+$\lVert{e^{t\Delta_{h}}P_{h}v-P_{h}e^{t\Delta}v}\rVert_{L^{q}}\leqslant ch^{2}t^{-1}\lVert{v}\rVert_{L^{q}},\quad\forall v\in L^{q}.$ (A.5)
+Applying complex interpolation between (A.3) and (A.5) yields
+$\lVert{e^{t\Delta_{h}}P_{h}v-P_{h}e^{t\Delta}v}\rVert_{L^{q}}\leqslant ch^{2}t^{\alpha/2-1}\lVert{v}\rVert_{\dot{H}^{\alpha,q}},\quad\forall v\in\dot{H}^{\alpha,q}.$ (A.6)
+Simultaneously, the following stability estimates hold for all $v\in L^{q}$:
+$\lVert{e^{t\Delta_{h}}P_{h}v}\rVert_{L^{q}}$ $\leqslant c\lVert{P_{h}v}\rVert_{L^{q}}\leqslant c\lVert{v}\rVert_{L^{q}},$
+$\lVert{P_{h}e^{t\Delta}v}\rVert_{L^{q}}$ $\leqslant c\lVert{e^{t\Delta}v}\rVert_{L^{q}}\leqslant c\lVert{v}\rVert_{L^{q}}.$
+These rely on the boundedness of $P_{h}$ on $L^{q}$ (see Eq.
+4.19), the boundedness of $e^{t\Delta_{h}}$ on $L^{q}$ (see Eq.
+4.28), and the boundedness of $e^{t\Delta}$ on $L^{q}$ (see Lemma 4.2).
+Consequently,
+$\lVert{e^{t\Delta_{h}}P_{h}v-P_{h}e^{t\Delta}v}\rVert_{L^{q}}\leqslant c\lVert{v}\rVert_{L^{q}},\quad\forall v\in L^{q}.$ (A.7)
+Complex interpolation between (A.3) and (A.7) then gives
+$\lVert{e^{t\Delta_{h}}P_{h}v-P_{h}e^{t\Delta}v}\rVert_{L^{q}}\leqslant ch^{\alpha}\lVert{v}\rVert_{\dot{H}^{\alpha,q}},\quad\forall v\in\dot{H}^{\alpha,q}.$ (A.8)
+Combining estimates (A.6) and (A.8), we conclude that
+$\lVert{e^{t\Delta_{h}}P_{h}v-P_{h}e^{t\Delta}v}\rVert_{L^{q}}\leqslant c\min\left\{h^{\alpha},\,h^{2}t^{\alpha/2-1}\right\}\lVert{v}\rVert_{\dot{H}^{\alpha,q}},\quad\forall v\in\dot{H}^{\alpha,q},$
+which establishes (A.2). This completes the proof of property (ii).
+$\blacksquare$